### v2

# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

Abstract.
This work is intended to present the basic properties of $KO$-theory for real $C^{*}$-algebras and to explain its relationship with complex $K$-theory and with $KR$- theory. Whenever possible we will rely upon proofs in printed literature, particularly the work of Karoubi [52], Wood [89], Schröder [76], and more recent work of Boersema [15] and J. M. Rosenberg [70]. In addition, we shall explain how $KO$-theory is related to the Ten-Fold Way in physics and point out how some deeper features of $KO$-theory for operator algebras may provide powerful new tools there. Commutative real $C^{*}$-algebras not of the form $\real(X)$ will play a special role.
Unfortunately, there is no single reference for $KO$-theory for operator algebras that begins to compare with Blackadar’s wonderful exposition of complex $K$-theory [14]. This work is intended to provide a platform upon which mathematicians and mathematical physicists can rely in order to use these new tools in their research. As we are writing for a diverse audience of functional analysts, topologists, and physicists, we often present material well-known to one group of people and unfamiliar to another.
Part I is a general introduction to real $C^{*}$-algebras, presenting basic examples and constructions, homotopy properties, and developing the general structure of the category of real $C^{*}$-algebras. After describing general homology theories, we define $KO_{*}$, display its basic properties and lead up to Bott Periodicity.
Part II is devoted to $KO^{*}$ for topological spaces. For this we review basic topological constructions, fibrations, fibre bundles, and vector bundles. We define $KO^{*}(X)$ and present a novel proof of Swan’s theorem relating $KO^{*}(X)$ to $KO_{*}(C^{\mathbb{R}}(X))$ for compact spaces.
Part III introduces Clifford algebras and their relationship with Bott periodicity and moves on to symmetric spaces and quantum symmetry. There is a section entitled ”Topology and Physics” written by our colleague Nadav Orion, which aims to make the path to physics clearer. We summarize this path via the Homotopy Road Map.
Part IV consists of a deeper look at some of the more advanced tools of $KO$-theory. We discuss various relationships between $KO$ and $K$-theory, display some concrete calculations of various $KO$-geoups, introduce $KO$-theory with coefficients, and demonstrate how elements of each $KO$-group may be represented in terms of unitaries.
1. Introduction
“Real $K$-theory behaves differently from complex $K$-theory.”
M.F. Atiyah [8], p. 4.
A little history
First, a little history to put $KO$-theory into historical context. The group $K_{0}(R)$ was introduced by Alexander Grothendieck in 1957 as part of his revolution in algebraic geometry, as a way to formulate what is now known as the Grothendieck-Riemann-Roch theorem. It was based upon coherent sheaves on an algebraic variety. It was extended using modules over arbitrary rings and is difficult to compute in general. [^1] For everything you would want to know about the early history of algebraic $K$-theory, see Weibel [87]. The group $K_{1}(R)$ was defined in 1962, $K_{2}(R)$ in 1967, and $K_{n}(R)$ in 1970.
At about the same time, Raoul Bott was studying the homotopy theory of the classical groups and discovered Bott periodicity - that homotopy groups of the classical unitary groups $\pi_{n}(\cup_{k}U_{k})$ were periodic of period 2 and the homotopy groups of the classical orthogonal groups $\pi_{n}(\cup_{k}O_{k})$ were periodic of period 8. This fit in, somehow, with Elie Cartan’s earlier classification of symmetric spaces, though certainly that was not clear at the time, and the connection to the Ten-Fold Way in physics was not understood.222other than the fact that $2+8=10.$
Michael Atiyah and Friedrich Hirzebruch made the fundamental breakthrough in 1959 (cf. [10]), introducing $K$-theory for complex vector bundles and $KO$-theory for real vector bundles, [^3] Atiyah used the notation $K(X)$ for complex bundles, $KO(X)$ for real bundles, and $KR(X)$ for Real bundles. Others used $KU(X)$ for the complex case. using Bott’s periodicity result to show that the associated cohomology theories on compact spaces were periodic (of period 2 and 8 respectively), and solving several major outstanding open questions in classical algebraic topology. The Atiyah-Singer Index Theorem, perhaps the most important mathematics result in the 1960’s, grew out of this milieu, and that is just one of many important results in algebraic topology established in the 1960’s using $K$-theory for compact spaces.
Atiyah, Bott, and Arnold Shapiro (1964) [9] (ABS) classified Clifford algebras, related them to $K$-theory and $KO$-theory, and predicted the existence of a proof of the Periodicity theorem based upon Clifford algebras.
As their paper was about to appear, Atiyah’s student R. Wood produced (1966)[89] a proof of the Periodicity Theorem along the lines suggested by ABS, where Clifford algebras play an essential role and where Bott Periodicity is a consequence of eight homotopy equivalences. He took a fundamental step forward, in that he stated the theorem in terms of the general linear groups of real and complex Banach algebras, and not just compact spaces, though his proof was rather sketchy. (See also Milnor [62]). This was followed by several detailed papers by Max Karoubi (cf. [45]). Karoubi generalized the Periodicity Theorem to Banach categories and Banach rings, proving periodicity in detail. Karoubi and Orlando Villamayor ([50], [51] ) explored its relationship to algebraic $K$-theory. A few years later Karoubi introduced Hermitian $K$-theory ([47], [49]), where he proved what he calls “A Fundamental Theorem” which is another road to Bott Periodicity (see the introduction to [47]). This theorem has important consequences nowadays in algebraic geometry and in quadratic forms over any field, not just the field of real numbers. Karoubi anticipated many of the results that we use in this work , In particular, the idea of desuspension was used to produce the negative algebraic $K$-theory groups developed by Bass.
The first major application of complex $K$-theory for $C^{*}$-algebras was the work of Brown-Douglas-Fillmore ( [28](1973), [29](1977) ), answering a question of Paul Halmos. Suppose that $T$ is an essentially normal operator.444meaning that $T$ is a bounded operator on complex Hilbert space and $T^{*}T-TT^{*}$ is a compact operator. When is $T$ the sum of a normal operator and a compact operator? The answer turns out to depend upon the complex $K$- homology theory of the essential spectrum of the operator and hence was best described in $C^{*}$-algebra terms. (Initially this involved their functor $Ext(X)$ for $X$ a compact subset of the plane555the essential spectrum of the operator, which was isomorphic to $K_{1}(X)$, but then their work was tied to the groups $KK_{*}(A,B)$ for $C^{*}$-algebras $A$ and $B$ of Gennadi Kasparov, which definitely required complex $K$-theory for $C^{*}$-algebras. [^6] The $KK$-groups to which we refer are a two-variable generalization of complex $K$-theory for $C^{*}$-algebras in several senses. For example, $KK_{*}({\mathbb{C}},B)\cong K_{*}(B)$ and $KK_{1}(C^{{\mathbb{C}}}(X),{\mathbb{C}})\cong Ext(X)$. The Kasparov groups are also defined for real Banach algebras and there are equivariant groups as well.
This success led to a spectacular flowering of work in functional analysis in the 1970’s that continues, using complex $K$-theory and Kasparov theory as primary tools. For example, the on-going classification of simple, separable, nuclear complex $C^{*}$-algebras relies heavily upon the Elliott invariant, which is built on complex $K$-theory.
Real $K$-theory was not as popular. For example, there are several different modern texts on complex $K$-theory for $C^{*}$-algebras, such as [14], [40], [44], [52], [69], but none for real $K$-theory. However, there has been a steady stream of results for $KO$ - both for spaces and for real $C^{*}$-algebras during this period, and in the last ten years its usefulness in various areas of mathematics has been established. For example, in Loring and Sorensen [57] (2016), $KO$-theory for real $C^{*}$-algebras is used to prove that almost commuting pairs of real $n\times n$ matrices can be approximated in a uniform way, independent of $n$, by exactly commuting pairs of matrices. This result extends to the real case what had been proven in the complex case much earlier by Lin [56].
There have been serious applications in algebraic topology as well. $KO_{*}$ plays a fundamental role in providing obstructions for the existence of a metric of positive scalar curvature on a compact manifold. It interacts with various versions of the Baum-Connes conjecture, and helps bring light to orbifold string theories in physics. For all of these matters and much more, we refer to Rosenberg’s deep and valuable survey article [70].
There has been a recent resurgence of interest in various problems in physics related to the Cartan classification of symmetric spaces, though the physicists would better describe it as studying symmetries in the atomic structure of matter. This generally involves $KO$-theory (since 8 of the 10 classes are related to $KO$, as we shall explain). Topological insulators and tilings are especially fertile areas for $K$ and for $KO$, as witness, for example, the work of Stone et al [81], Hastings and Loring [39], and Akkermans, Don, Rosenberg, and Schochet [2].
Section Summaries
This work is designed to provide a basic source for those in mathematics and physics who wish to use $KO$-theory in their work. There is no such source at present: we mean to remedy that problem.
We now briefly summarize our work.
Part 1: $KO_{*}$- Theory for Real $C^{*}$-algebras
Section 2 deals with the basics of real $C^{*}$-algebras, giving many examples. There is a fundamental difference between the real and complex commutative settings. Every complex commutative $C^{*}$-algebra is isomorphic to an algebra of continuous complex-valued functions on a locally compact space, by the Gelfand theory. The analogous statement is not true for real $C^{*}$-algebras and will lead to the possibility of “desuspending” a real $C^{*}$-algebra.
[…]
Section 4 is all about homotopy constructions for real $C^{*}$-algebras.. We recall the definitions of the cylinder, the cone, and the suspension and we introduce the desuspension of a real $C^{*}$-algebra as well as cofibrations.
Section 5 Here we introduce axioms for homology theories on real $C^{*}$-algebras and show how to use them. Our interest is in showing how much of the general theory of $KO_{*}$ goes through without invoking the depth of the Bott Periodicity theorem, as well as illustrating many of the similarities between real and complex $K$-theory. We also introduce cofibre homology theories.
Section 6 In Section 6 we introduce the homology theory $KO_{*}$ for real $C^{*}$-algebras, the main focus of this work . There are several distinct approaches to $KO$. We use what we believe to be the simplest approach, not relating much to algebraic $K$-theory, Hermitian $K$-theory, or other homotopy constructions. We show that $KO_{*}$ is a homology theory and introduce some of its relationships to $K_{*}$. Still no mention of Bott periodicity.
Section 7 Ordinary cohomology of a topological space is a graded ring, and this is also true for the $K$ and $KO$-theories when focused on spaces, but when we pass to $C^{*}$-algebras the ring structure disappears. Nevertheless, we show that some partial structure is retained. This structure allows us to state Künneth theorems that facilitate computation in many useful cases.
Section 8 In Section 8 we (finally!) introduce Bott Periodicity, both in the complex and the real settings, and we explain how this allows us to extend $KO$-theory and $K$-theory to periodic homology theories defined for all integers.
Part 2: $KO^{*}$-theory for Spaces
Section 9 Section 9 introduces basic constructions in topological spaces such as the product, the join, the smash product and the suspension, and establishes the fundamentally important identification $F(SX,Y)\cong F(X,\Omega Y)$. We will assume throughout the work that all topological spaces are compactly generated. We introduce $CW$-complexes briefly.
Section 10 This is a classical homotopy section. We introduce fibrations, fibre bundles, principal bundles and briefly explain their important properties. We conclude by discussing the universal principal bundle $\mathbf{O}\longrightarrow{\bf{EO}}\longrightarrow{\bf{BO}}$.
Section 11 In this short but very important section we revisit the Bott Periodicity theorem, state it in terms of the classifying space $\bf{BO}$} and present the first version of the Homotopy Road Map which connects Bott periodicity via classical homotopy to Clifford algebras to symmetric spaces to physics.
Section 12 Bott Periodicity tells us to concentrate on the homotopy groups of spaces like $\mathbf{U}/\mathbf{O}$, and it is true that
[…]
and so when $n$ is large then
$\pi_{k}(\mathbf{U}/\mathbf{O})\cong\pi_{k}(U_{n}/O_{n})$
but sometimes (and specifically, in some recent developments in physics) we need to know $\pi_{k}(U_{n}/O_{n})$ when $n$ is small. So we pause to spend a section computing such groups for $k=0,1,2$.
Section 13 Section 13 introduces real and complex vector bundles, mostly over compact spaces, and states the classification theorem with which isomorphism classes of vector bundles over $X$ are related to homotopy classes of maps from $X$ into the appropriate Grassmann manifold.
Section 14 Here we introduce $K$ and $KO$-theory for compact spaces, defined via complex and real vector bundles respectively. This material is very well-treated in various books and so we simply state the highlights. We mention the Chern character and the Pontrjagin carrier which relate $K$ and $KO$-theory for compact spaces to the cohomology of those spaces.
Section 15 Section 15 is devoted to establishing the relationship between $KO^{*}$-theory for spaces and $KO_{*}$-theory for real $C^{*}$-algebras and similarly for $K$-theory and complex $C^{*}$-algebras. We present a novel proof of Swan’s classical result relating the $KO_{*}$-theory of a real $C^{*}$-algebra $C^{\mathbb{R}}(X)$ and the topological $KO^{*}$-theory of the compact space $X$.
Section 16 Section 16 is special. Atiyah introduced $KR^{*}(X,\tau)$ (where $X$ is a compact space and $\tau$ is an involution of $X$) for specific applications in functional analysis, as we shall briefly explain. We show that there is a natural isomorphism of the form
$KR^{*}(X,\tau)\cong KO_{*}(A),$
where $A$ is a real commutative $C^{*}$-algebra constructed from $X$ and from the involution $\tau$.
Part 3: Clifford Algebras and Symmetric Spaces
Section 17 This section is a quick exposé of the basic information about real Clifford algebras that we need going forward. We concentrate upon the Lie groups of each Clifford algebra in the standard sequence and the homogeneous spaces constructed from them. We emphasize the relationship between this sequence of homogenous spaces and the sequence of homogeneous spaces obtained from the Bott Periodicity theorem, referring to the Homotopy Road Map for reference. We contrast the Clifford perspective and the Lie algebra perspective for this list of spaces and explain their relationships.
Section 18 gives a brief discussion of the Wood-Karoubi proof of Bott periodicity and, specifically, its relationship to Clifford algebras.
Section 19 In this short section we mention complex Clifford algebras and indicate how they too relate to the Homotopy Road Map via Bott Periodicity.
Section 20 Section 20 gives a basic introduction to symmetric spaces. Every symmetric space arises as a homogeneous space of Lie groups and so the discussion turns at once to associated Lie groups and Lie algebras. Compact symmetric spaces have been classified and there are ten different types. These ten types correspond to the ten rows on the Homotopy Road Map and we present Step 2 of the Road Map adding on this correspondence. We then present a ”Timely Example”, namely the symmetric space $SU_{4}/O_{4}$. We present a basis for the Lie algebras that reflect their geometry and define a symmetry $T$ that acts on the Lie algebra with very specific characteristics. Appendix 20.1 is by Nadav Orion. In it he explains the basic connection between topology and physics in our context.
Section 21 This section connects us directly with the Ten-Fold Way in physics. We discuss the actions of Hamiltonians and the symmetries that they may or may not respect. We introduce the $T,C,S$ symmetries and link them to the ten types of compact symmetric spaces, adding them in order to complete the Homotopy Road Map.
Part 4: A Deeper Look at the Structure of $KO$ and $K$
Section 22 Here we present several inter-related deeper properties of real and complex $K$-theory, namely certain natural transformations, general KKOK-KO-relationships, product structures, and a spectral sequence greatly generalizing the Atiyah-Hirzebruch spectral sequence.
Section 23 The section is entitled ”Examples and Calculations” which describes exactly what is contained in it.
Section 24 This section presents a novel approach to the groups $KO_{*}(A)$, showing that each of the eight groups may be described concretely in terms of unitary elements of A and associated algebras such as $M_{k}(A)$.
Section 25 This section contains our treatment of $KO$-theory with coefficients. It is well-known that $H^{*}(X;G)$, the cohomology of a space with coefficients in a groups $G$ is sometimes much more useful and certainly more tractable then when you take $G={\mathbb{Z}}$. We suspect that the same will be true for $KO$ and hence we present how to define it.
What’s not in the work ? We have not discussed the Künneth theorem in full generality, nor united $K$-theory, which combines $K$, KOKO and $KSc$ theories. We have not discussed some of the applications of $KO$-theory in differential geometry, as this requires quite a bit of background to appreciate. Nor have we discussed $KO$-homology theory for compact spaces. We have not discussed the Hopkins-Hovey isomorphism [43]
$MSpin_{*}(X)\otimes_{Mspin(pt)}KO_{*}(pt)\cong KO_{*}(X)$
relating $KO_{*}(X)$ and Spin bordism. We have not discussed the generalization of $KO$-homology for spaces to $KO^{*}(A)$ to a cohomology theory for real $C^{*}$-algebras, defined most easily using the real version of Kasparov’s $KK$-theory. There is some literature on each topic, but these are substantially more technical and, to this point, less central than the parts of the theory that we have covered.
This work arose in connection with the second author’s collaboration with Erik Akkermanns. We are most grateful to him and to his students, particularly to Nadav Orion, who have provided invaluable feedback and true insight in the development of this work. We are also very grateful to Jean Bellissard, Joachim Cuntz, Yigal Kamel, John Klein, Chris Phillips, Emil Prodan, Jonathan Rosenberg and Chuck Weibel, for providing us with various deep insights as this work developed. We especially want to acknowledge with great pleasure the extensive philosophical and detailed help that we have received from Max Karoubi, who, more than anyone, is responsible for the very existence of $KO$-theory for spaces and algebras as we know it.
Part I $KO_{*}$-Theory for Real $C^{*}$-Algebras
2. Real $C^{*}$-algebras
Two points on notation. First, we have been confused over the years by various papers using $C(X)$, $K_{0}(A)$, and the like, without specifying whether they had real or complex versions in mind.777and we each plead guilty to this sin. Frequently authors deal with this by stating at the beginning that everything is complex or that everything is real. In this work both possibilities will arise. We will be more precise.
A second point. Usually the real numbers occur in this work as a real vector space or with their real $C^{*}$-algebra structure, and when that happens we shall denote them as $\mathbb{R}$}. Sometimes, though, we want to think of the real numbers simply as a locally compact topological space, In that case we shall use the notation $\bf{R}$}. We will label the points of $\bf{R}$} as usual, so that we can write continuous functions, such as $x\to x^{2}$ or the involution $x\to-x$.
For any compact [^8] We assume that locally compact spaces and compact spaces are always Hausdorff. space $X$ and normed algebra $A$, let $C(X,A)$ denote the algebra of continuous functions from $X$ to $A$ with pointwise operations and the sup norm. Similarly, for a locally compact space $X$, let $C_{o}(X,A)$ denote the algebra of continuous functions from $X$ to $A$ which vanish at infinity. We then abbreviate as follows:
$\real(X)=C(X,{\mathbb{R}})\qquad\qquad C_{o}^{{\mathbb{R}}}(X)=C_{o}(X,{\mathbb{R}})$
$C^{{\mathbb{C}}}(X)=C(X,{\mathbb{C}})\qquad\qquad C_{o}^{{\mathbb{C}}}(X)=C_{o}(X,{\mathbb{C}})$
When developing $C^{*}$-algebras and its $K$-theory in the past, people generally used $C(X)$ for complex-valued functions on $X$ and $K_{*}$ for complex $K$-theory, because almost everything involving $C^{*}$-algebras was over the complex numbers. That is no longer the case, but, following tradition, we shall continue to use $K_{*}(A)$ for complex $K$-theory. We shall use $KO_{*}$ for real $K$-theory throughout.
Definition 2.1.
A real $C^{*}$-algebra is a real Banach *-algebra AA
  1. (1)
    which is isometrically *-isomorphic to a norm-closed *-algebra of bounded operators on real Hilbert space.
  2. (2)
    or, equivalently, such that for all $a\in A$,
[…]
    and in addition, for all $a$, the element $a^{*}a$ has spectum contained in $0,\infty)$.
See T. Palmer [67] for a helpful discussion of these definitions. We let $\mathscr{C}^{*}_{real}$ denote the category of real $C^{*}$-algebras and *-homomorphisms (always assumed to be continuous; frequently we simply say ”maps”).
We focus on examples.
  1. (1)
    If $A$ is a finite-dimensional real $C^{*}$-algebra, then it is a direct sum of matrix algebras $M_{n}({\mathbb{F}})$, where $\mathbb{F}$$\mathbb{R}$${\mathbb{C}}$,\,\, or $\mathbb{H}$}, the quaternions.
  2. (2)
    If $X$ is a locally compact space then $C_{o}^{{\mathbb{R}}}(X)$ with the sup norm is a commutative real $C^{*}$-algebra. If $X$ is compact then $\real(X)$ is a unital commutative real $C^{*}$-algebra.
  3. (3)
    If we take $\mathcal{H}$ to be a real Hilbert space, then $\mathcal{B}(\mathcal{H})$, the bounded linear operators on $\mathcal{H}$, is itself a real $C^{*}$-algebra. If $\mathcal{S}$ is any finite or infinite subset of $\mathcal{B}(\mathcal{H})$, then we let $C_{\mathbb{R}}^{*}(\mathcal{S})$ denote the smallest closed *-subalgebra which contains the set $\mathcal{S}$. So we may, for example, take two bounded operators $T_{1},\,T_{2}$ on $\mathcal{H}$, form the *-subalgebra of $\mathcal{B}(\mathcal{H})$ that they generate, close it in norm, and this gives us the real $C^{*}$-algebra $C_{\mathbb{R}}^{*}(T_{1},\,T_{2})$. It will be commutative if and only if $T_{1}T_{2}=T_{2}T_{1}$.
  4. (4)
    Take $\mathcal{H}_{\mathbb{C}}$ to be a complex Hilbert space and choose an orthonormal basis. The closed real span of that basis forms a real Hilbert space $\mathcal{H}_{\mathbb{R}}$, and then $\mathcal{B}(\mathcal{H}_{\mathbb{R}})$ is a real $C^{*}$-algebra.
  5. (5)
    Any *-closed ideal of a real $C^{*}$-algebra is again a real $C^{*}$-algebra - for example, the real compact operators ${\mathcal{K}}_{\mathbb{R}}\subset\mathcal{B}(\mathcal{H}_{\mathbb{R}})$.
  6. (6)
    Suppose that $A$ is a real $C^{*}$-algebra and $X$ is a locally compact space. Then, as mentioned above, we may form $C_{o}^{{\mathbb{R}}}(X,A)$, all continuous functions from $X$ to $A$ vanishing at infinity. With pointwise operations and sup norm, this forms a real $C^{*}$-algebra. Similarly, if $X$ is compact and $A$ is unital then $\real(X,A)$ is a unital real $C^{*}$-algebra.
  7. (7)
    Suppose that $X$ is a locally compact space and $\tau:X\to X$ is an involution of $X$; that is, a homeomorphism of $X$ with $\tau^{2}=1$. Define
    $C_{o}^{{\mathbb{R}}}(X,\tau)=\{f\in C_{o}^{{\mathbb{C}}}(X):f(\tau(x))=\overline{f(x)}\,\,\,\forall x\in X\}.$
    Then $C_{o}^{{\mathbb{R}}}(X,\tau)$ is a real commutative $C^{*}$-algebra. It is a nice exercise to show that
    $C_{o}^{{\mathbb{R}}}(X,\tau)\cong C_{o}^{{\mathbb{R}}}(X)$
    if and only if $\tau=id$ is the identity homeomorphism.
  8. (8)
[…]
    $C_{o}^{{\mathbb{R}}}(i{\bf{R}})=\{f\in C_{o}^{{\mathbb{C}}}({\bf{R}}):f(-x)=\overline{f(x)}\,\,\,\forall x\in{\mathbb{R}}\}.$
  9. (9)
    Similarly, the circle $S^{1}$ has three obvious involutions. Regarding $S^{1}$ as the unit circle in the complex plane, they are
    $z\overset{id}{\longrightarrow}z\qquad\qquad z\overset{\hat{\tau}}{\longrightarrow}\bar{z}\qquad\qquad z\overset{\sigma}{\longrightarrow}-z$
[…]
    $C^{{\mathbb{R}}}(iS^{1})=\real(S^{1},\hat{\tau})=\{f\in C^{{\mathbb{C}}}(S^{1}):f(\bar{z})=\overline{f(z)}\}.$
  10. (10)
    The real Toeplitz algebra $\mathcal{T}$$_{\mathbb{R}}$ is defined as the real $C^{*}$-algebra [^9] To be more precise, take $\mathcal{H}=L^{2}_{\mathbb{R}}$ to be real $L^{2}$-functions on the circle and $\mathcal{H}_{\mathbb{R}}^{2}\subset L^{2}_{\mathbb{R}}$ to be the real Hardy space, with canonical projection $P:L^{2}_{\mathbb{R}}\longrightarrow\mathcal{H}_{\mathbb{R}}^{2}$ If $\phi$ is a complex-valued continuous function on the circle then define $T_{\phi}f=P(\phi f)$.
    $\mathcal{T}_{\mathbb{R}}=\{T_{\phi}+K:\phi\in C^{{\mathbb{R}}}(iS^{1}),K\in{\mathcal{K}}_{\mathbb{R}}\}.$
    with associated short exact sequence
    $0\longrightarrow{\mathcal{K}}_{\mathbb{R}}\longrightarrow\mathcal{T}_{\mathbb{R}}\overset{\sigma}{\longrightarrow}C^{{\mathbb{R}}}(iS^{1})\to 0.$
  11. (11)
    Let $X=\{a,b\}$, the space with two discrete points. It has two obvious involutions, namely the identity and the involution $\tau$ which interchanges the two points. In the first case, we have
    $\real(X,id)\cong\real(X)\cong{\mathbb{R}}\oplus{\mathbb{R}}$
    with pointwise operations. In the second case, we have
    $\real(X,\tau)\cong{\mathbb{C}}$
    the complex numbers, regarded as a real $C^{*}$-algebra. This is yet another example of a real, commutative unital $C^{*}$-algebra which is not of the form $\real(X)$.
Recall that every complex unital commutative $C^{*}$-algebra $A$ is isometrically isomorphic to $C^{{\mathbb{C}}}(X)$ for some compact space $X$, and a non-unital commutative $C^{*}$-algebra is isometrically isomorphic to $C_{o}^{{\mathbb{C}}}(X)$ for some locally compact space $X$. The situation is profoundly different for real $C^{*}$-algebras, as witness the previous example. Here is the full story.
Theorem 2.2.
(Arens-Kaplansky [4], Thm. 9.1) Suppose that $A$ is a commutative real $C^{*}$-algebra. Then:
  1. (1)
    If $A$ is unital, then there exists a compact space $X$ and an involution h $\tau:X\to X$ such that $A\cong\real(X,\tau)$.
  2. (2)
    If $A$ is not unital, then there exists a locally compact space $X$ and an involution $\tau:X\to X$ such that $A\cong C_{o}^{{\mathbb{R}}}(X,\tau)$.
  3. (3)
    There is an equivalence of categories between the categoy of commutative real unital $C^{*}$-algebras and the category of compact spaces with involution and involution-respecting maps.
  4. (4)
    Similarly, there is an equivalence of categories between the categoy of commutative real $C^{*}$-algebras and the category of locally compact spaces with involution and involution-respecting maps.
Remark 2.3.
The real commutative $C^{*}$-algebras $C_{o}^{{\mathbb{R}}}(i{\mathbb{R}})$ and $\real(iS^{1})$ should be regarded as the poster children for commutative real $C^{*}$-algebras that have no complex analog. They will both be very important in our story.
3. New $C^{*}$-algebras from Old Ones
In this section we review several methods of constructing new $C^{*}$-algebras from old ones.
Definition 3.1.
Let $H$ be a real Hilbert space. Given a real $C^{*}$-algebra $A$, its complexification $A_{\mathbb{C}}=A+iA$ is a complex $C^{*}$-algebra:
  1. (1)
    $A_{\mathbb{C}}$ is a self-adjoint complex subalgebra.
  2. (2)
    $A_{\mathbb{C}}$ is complete in the induced norm and hence is a complex $C^{*}$-algebra. [^10] We refer the reader to Goodearl [36] pp. 66-67.
  3. (3)
    If $f:A\to B$ is a morphism of real $C^{*}$-algebras then it extends in an obvious way to $f_{\mathbb{C}}:A_{\mathbb{C}}\to B_{\mathbb{C}}$ by
    $f_{\mathbb{C}}(s+it)=f(s)+if(t).$
  4. (4)
    Complexification is a covariant functor from the category of real $C^{*}$-algebras to the category of complex $C^{*}$-algebras, denoted
    $c:\mathscr{C}^{*}_{real}\longrightarrow\mathscr{C}^{*}_{complex}.$
The complexification functor is not an equivalence of categories. Not every complex $C^{*}$-algebra arises as the complexification of a real $C^{*}$-algebra. Further, two non-isomorphic real $C^{*}$ algebras can have isomorphic complexifications. For instance, $M_{2}({\mathbb{R}})$ is not isomorphic to $\mathbb{H}$} as real $C^{*}$-algebras, and yet
$M_{2}({\mathbb{R}})_{\mathbb{C}}\cong{\mathbb{H}}_{\mathbb{C}}\cong M_{2}({\mathbb{C}}).$
Similarly, the three different examples of involutions on $S^{1}$ all lead to isomorphic complexifications:
$\real(S^{1})_{\mathbb{C}}\cong\real(iS^{1})_{\mathbb{C}}\cong\real(S^{1},\sigma)_{\mathbb{C}}\cong C^{{\mathbb{C}}}(S^{1}).$
Proposition 3.2.
Let $A$ be any real $C^{*}$-algebra. Then there is a antimultiplicative involutive complex-linear *-automorphism $\sigma:A_{\mathbb{C}}\to A_{\mathbb{C}}$ given by
$\sigma(s+it)=s^{*}+it^{*}\;$
which satisfies
$A=\{a\in A_{\mathbb{C}}\mid a^{*}=\sigma(a)\}\;.$
Both of the statements of this Proposition are easy to verify directly. The fact that $A$ can be recovered from $\sigma$ means that there is an equivalence between the category of real $C^{*}$-algebras and the category of pairs $(B,\sigma)$ where $B$ is a complex $C^{*}$-algebra and $\sigma$ is a antimultiplicative involutive complex-linear *-automorphism of $B$.
Example 3.3.
Here are some examples of how this cateogrical equivalence works for particular real $C^{*}$-algebras.
  • The real $C^{*}$-algebra $M_{n}({\mathbb{R}})$ corresponds to the pair $(M_{n}({\mathbb{C}}),\text{Tr})$ where Tr is the matrix transpose operator.
  • The $C^{*}$-algebra of quaternions $\mathbb{H}$} corresponds to the pair $(M_{2}({\mathbb{C}}),\sharp)$ where
    $\sharp\colon\begin{pmatrix}a&b\\ c&d\end{pmatrix}\mapsto\begin{pmatrix}d&-b\\ -c&a\end{pmatrix}\;$.
  • Let $\mathcal{H}$ be a real Hilbert space and let $\mathcal{B}(\mathcal{H})$ be the real $C^{*}$-algebra of bounded operators. This corresponds to the pair $(\mathcal{B}(\mathcal{H}_{\mathbb{C}}),\text{Tr})$ where Tr is the transpose operator defined by
    $\text{Tr}(T)=(\overline{T^{*}(\overline{x}))}\;.$
  • For a topological space $X$, the real $C^{*}$-algebra $\real(X)$ corresponds to the pair $(C^{{\mathbb{C}}}(X),\mathrm{id})$.
  • For a topological space $X$ with continuous involution, the real $C^{*}$-algebra $\real(X,\tau)$ corresponds to the pair $(C^{{\mathbb{C}}}(X),\tau^{*})$ where $\tau^{*}$ is the involution on $C^{{\mathbb{C}}}(X)$ defined by $\tau^{*}(f)(x)=f(\tau(x))$.
We note that for a real $C^{*}$-algebras, we could alternatively focus on the involution
$s+it\mapsto s-it\;$
on $A_{\mathbb{C}}$ and recover $A$ as the set of fixed points. Under this perspective there is an equivalence between the category of real $C^{*}$-algebras and the category of pairs $(B,\sigma)$ where $B$ is a complex $C^{*}$-algebra and $\sigma$ is a conjugative-linear involutive *-automorphism of $B$.
Definition 3.4.
Suppose that $A$ is a real $C^{*}$-algebra, unital or not. Define its unitization $A^{+}$ by $A^{+}=A\oplus{\mathbb{R}}$ additively, with multiplication given by
$(a,r)\times(b,s)=(ab+rb+as,rs)$
and norm given by
(a,r)=a+r.||(a,r)||=||a||+||r||.
Then $A^{+}$ is a real unital $C^{*}$-algebra, $A$ is a closed ideal in $A^{+}$ and $A^{+}/A\cong{\mathbb{R}}$, so there is a natural short exact sequence
$0\longrightarrow A\longrightarrow A^{+}\overset{\pi}{\longrightarrow}{\mathbb{R}}\longrightarrow 0.$
If $A$ already were unital, then $A^{+}\cong A\oplus{\mathbb{R}}$ as real $C^{*}$-algebras.
Here are some examples:
  1. (1)
    If $X$ is a locally compact space, then
    $C_{o}^{{\mathbb{R}}}(X)^{+}\cong\real(X^{+})$
    where $X^{+}$ denotes the one-point compactification of $X$. For example, $C_{o}^{{\mathbb{R}}}({\bf{R}})^{+}\cong\real(S^{1})$ since the one-point compactification of the real numbers is the circle.
  2. (2)
    Similarly,
    $C_{o}^{{\mathbb{R}}}(i{\bf{R}})^{+}\cong C^{{\mathbb{R}}}(iS^{1}).$
Definition 3.5.
Given real $C^{*}$-algebras $A_{1},A_{2},\dots A_{n}$ we may define their sum $A_{1}\oplus\dots\oplus A_{n}$ just as in the complex setting, by taking their algebraic direct sum with the evident * and close it in the evident norm. Similarly, we may take arbitrary sums $\oplus_{j}A_{j}$ by taking the algebraic direct sum and giving it the appropriate norm, and closing it just as in the complex setting. It is not hard to show that the direct sum operation respects complexification. That is,
$(A\oplus B)_{\mathbb{C}}\cong A_{\mathbb{C}}\oplus B_{\mathbb{C}}$
and similarly for arbitrary direct sums.
Definition 3.6.
More generally, given a sequence
$A_{1}\longrightarrow A_{2}\longrightarrow\dots\longrightarrow A_{n}\longrightarrow\dots$
[…]
The direct limit operation also respects complexification. That is, there is a natural isomorphism
[…]
Definition 3.7.
Suppose that $A$ and $B$ are real $C^{*}$-algebras. Then we may form their tensor product $A\otimes B$ by taking their algebraic tensor product over $\mathbb{R}$} and then completing that algebra in a suitable $C^{*}$-norm. The complication here is that in general, just as in the complex case, there is more than one “suitable” norm on the algebraic tensor product - there is a minimal norm, a maximal norm, and possibly many in between. The minimal norm has the very nice property of being concrete. The maximal norm has the very nice property of preserving exact sequences. We want both properties and hence we make a blanket assumption: whenever we write $A\otimes B$ then at least one of the two $C^{*}$-algebras has the property that its complexification is nuclear and exact. [^12] For example, any commutative complex $C^{*}$-algebra is nuclear and exact. Please see Blackadar [14] for definitions of nuclear and exact $C^{*}$-algebras as well as for other operator algebra terms that may seem mysterious. This does the job, because, using the minimal tensor product norm,
  1. (1)
    The complexification of a tensor product of two real $C^{*}$-algebras is isomorphic to the the tensor product of the complexification of the two real $C^{*}$-algebras, and
  2. (2)
    A short exact sequence of real $C^{*}$-algebras is exact if and only if the complexification is exact.
We are particularly indebted to J. Cuntz for clarifying this matter for us.
Proposition 3.8.
  1. (1)
    Let $A_{n}$ be a directed system of complex $C^{*}$-algebras and let $B$ be a complex $C^{*}$-algebra. Then there is an isomorphism
[…]
  2. (2)
    The same result holds in the category of real $C^{*}$-algebras.
The tensor product is the maximal tensor product. If we assume, say, that $B$ is nuclear then there is a unique tensor product in each case.
Proof. From the statement for complex $C^{*}$-algebras, we prove the statement for real $C^{*}$-algebras by working in the complexification. So consider first the complex case.
Assume first that $A_{n}$ is unital for all $n$ and that all of the connecting homomorphisms are unital. Assume also that $B$ is unital. The general case can be easily deduced from this. Let $A=\lim_{n\to\infty}A_{n}$ and let
$\alpha_{n}\colon A_{n}\rightarrow A$
be the natural homomorphism into the limit, which is also unital.
First of all, note that the sequence of maps
$\alpha_{n}\otimes 1\colon A_{n}\otimes B\rightarrow A\otimes B$
are coherent with respect to the inductive sequence $\{A_{n}\otimes B\}_{n}$, and so there is a homomorphism in the limit
[…]
For the other direction, there are maps
$\beta_{n}\colon A_{n}\rightarrow A_{n}\otimes B$
given by $\beta_{n}(a)=a\otimes 1$. If we compose $\beta_{n}$ with the limit map
[…]
then we have a system of coherent homomorphisms
[…]
which yields the map
[…]
in the limit. There is also a map
[…]
which restrict to $f$ and $g$ on $A\otimes 1$ and $1\otimes B$ respectively.
Finally, check that $\phi$ and $\psi$ are inverses to each other.
Here are some more examples:
  1. (1)
    $C_{o}^{{\mathbb{R}}}({\bf{R}})^{\otimes n}\,\cong\,C_{o}^{{\mathbb{R}}}({\bf{R}}^{n})$
  2. (2)
    $C^{{\mathbb{R}}}(iS^{1})^{\otimes n}\,\cong\,\real(S^{1}\times\dots\times S^{1},\hat{\tau}^{n})$
    where $\hat{\tau}^{n}:(S^{1}\times\dots\times S^{1})\longrightarrow(S^{1}\times\dots\times S^{1})$ is given by
    $\hat{\tau}^{n}(x_{1},\dots,x_{n})=(\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{n}).$
  3. (3)
    If $X$ is a locally compact space with involution $\tau$ then
    $C_{o}^{{\mathbb{R}}}(i{\bf{R}})\otimes C_{o}^{{\mathbb{R}}}(X,\tau)\cong C_{o}^{{\mathbb{R}}}({\bf{R}}\times X,\hat{\tau}\times\tau).$
  4. (4)
    More generally, if $A$ is a real $C^{*}$-algebra and the associated conjugate linear involution on $A_{\mathbb{C}}$ is $\sigma$, then
    $C_{o}^{{\mathbb{R}}}(i{\bf{R}})\otimes A\cong\{f\in C_{o}^{{\mathbb{C}}}({\bf{R}},A_{\mathbb{C}}):f(-x)=\sigma(f(x))\}.$
  5. (5)
    Finally,
    $C_{o}^{{\mathbb{R}}}(X_{1},\tau_{1})\oplus C_{o}^{{\mathbb{R}}}(X_{2},\tau_{2})\cong C_{o}^{{\mathbb{R}}}(X_{1}\times X_{2},\tau_{1}\times\tau_{2}).$
The section is titled “New $C^{*}$-algebras from Old Ones” but our final example should be titled “Old $C^{*}$-algebras from Old Ones.” This fact will lead eventually, as we shall see, to a strong connection between real and quaternionic $K$-theory.
Proposition 3.9.
(cf. [9] 4.1) There is a natural isomorphism
${\mathbb{H}}\otimes_{\mathbb{R}}{\mathbb{H}}\cong M_{4}({\mathbb{R}}).$
Proof.
We start by consider $\mathbb{H}$} in terms of the complex $C^{*}$-algebra with involution $(M_{n}({\mathbb{C}}),\sharp)$, as in Example 3.3. Then $\mathbb{H}$$\otimes_{{\mathbb{R}}}{\mathbb{H}}$ corresponds to
$(M_{2}({\mathbb{C}})\otimes_{{\mathbb{C}}}M_{2}({\mathbb{C}}),\sharp\otimes\sharp)=(M_{4}({\mathbb{C}}),\sharp\otimes\sharp)$
where $\sharp\otimes\sharp$ is the involution on $M_{4}({\mathbb{C}})$ defined by
$\sharp\colon\begin{bmatrix}A&B\\ C&D\end{bmatrix}\mapsto\begin{bmatrix}D^{\sharp}&-B^{\sharp}\\ -C^{\sharp}&A^{\sharp}\end{bmatrix}\;$.
We claim that there is an isomorphism $(M_{4}({\mathbb{C}}),\rm{Tr})\cong(M_{4}({\mathbb{C}}),\sharp\otimes\sharp)$ as complex $C^{*}$-algebras with involution. It will follow from this that there is a real $C^{*}$-algebra isomorphism $M_{4}({\mathbb{R}})\cong{\mathbb{H}}\otimes_{{\mathbb{R}}}{\mathbb{H}}$.
To produce this isomorphism, let
$Q=\tfrac{1}{\sqrt{2}}\begin{bmatrix}1&0&0&-i\\ 0&1&i&0\\ 0&i&1&0\\ -i&0&0&1\end{bmatrix}\;$.
Check that $Q$ is a unitary and define an isomorphism $\phi\colon M_{4}({\mathbb{C}})\rightarrow M_{4}({\mathbb{C}})$ by
$\phi(A)=QAQ^{*}\;.$
Then check that
$\phi(A^{\rm{Tr}})=\phi(A)^{\sharp\otimes\sharp}$
which therefore gives us the desired isomorphism $(M_{4}({\mathbb{C}}),\rm{Tr})\cong(M_{4}({\mathbb{C}}),\sharp\otimes\sharp)$. ∎
4. Homotopy Constructions
In this section we present some of the basic homotopy constructions that we shall need in subsequent sections. This is important philosophically, since it illustrates how much the homology theories $K_{*}$ and $KO_{*}$ have in common and opens the door for other possibilities in the future. [^13] For example, the homology theories $K_{*}(A;G)$ (complex $K$-theory with coefficients in an abelian group $G$ such as ${\mathbb{Z}}_{2}$ or $\mathbb{Q}$} ) have been found to be quite useful in certain contexts. It may turn out that similar constructions starting with $KO_{*}$ may also be of service. All of these theories are homology theories.
To set up a one needs first of all to specify a category which will be the domain of the theory. For $K_{*}$ we sometimes use, for example,
  1. (1)
    All complex $C^{*}$-algebras and their homomorphisms
  2. (2)
    All separable complex $C^{*}$-algebras and their homomorphisms
  3. (3)
    The bootstrap category [^14] First defined by Schochet in [73] (a certain subcategory of the previous example)
  4. (4)
    All commutative complex $C^{*}$-algebras and their homomorphisms151515which, by the Gelfand transform, corresponds to all compact spaces and continuous functions.
For $KO_{*}$ we will use a similar list:
  1. (1)
    All real $C^{*}$-algebras and their homomorphisms
  2. (2)
    All separable real $C^{*}$-algebras and their homomorphisms
  3. (3)
    The real bootstrap category (a certain subcategory of the previous example as yet unexplored as far as we know)
  4. (4)
    All commutative real $C^{*}$-algebras and their homomorphisms161616which, as we have noted, does not correspond to all compact spaces and continuous functions but is larger!.
The second author’s paper [74] lays out almost all of the constructions that we need, but restricted to the complex categories specified above, as the concern then was $K_{*}$. These constructions carry over immediately to the real category. There is one very striking addition that comes up - you cannot “desuspend” in the complex world, but in the real world you can. [^17] This is slightly exaggerated. If $A$ is a real $C^{*}$-algebra then we have real $C^{*}$-algebras $SA$ and $VA$ defined below, and they behave very differently. In the complex world the analogous $C^{*}$-algebras are isomorphic as complex $C^{*}$-algebras. Stay tuned!
We begin by defining some elementary constructions. Let $A$ be a real $C^{*}$-algebra.
Definition 4.1.
The cylinder of $A$, denoted $IA$, is $\real([0,1],A)$ with canonical maps $p_{t}:IA\to A$ given by $p_{t}(\xi)=\xi(t)$.
Definition 4.2.
The cone of $A$, denoted by $CA$, is defined by
$CA=\{\xi\in IA\,|\,\xi(0)=0\}=Ker(p_{0}:IA\to A)$
with associated short exact sequence
$0\longrightarrow CA\longrightarrow IA\overset{p_{0}}{\longrightarrow}A\longrightarrow 0.$
Definition 4.3.
The suspension of $A$, denoted $SA$, is defined by
$SA=\{\xi\in IA\,|\,\xi(0)=\xi(1)=0\}\cong C_{o}^{{\mathbb{R}}}({\bf{R}},A)\cong C_{o}^{{\mathbb{R}}}({\bf{R}})\otimes A$
with associated short exact sequence
$0\longrightarrow SA\longrightarrow CA\overset{p_{1}}{\longrightarrow}A\longrightarrow 0.$
For each $n=0,1,2,$... we define the n-th suspension of $A$, as
$S^{n}A=C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},A)\cong C_{o}^{{\mathbb{R}}}({\bf{R}}^{n})\otimes A,$
and we note at once that
$S(S^{n}A)\cong S^{n+1}A.$
For example, if $A=C^{\mathbb{R}}(X)$, with $X$ compact, then
$S^{n}A=C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},C(X,{\mathbb{R}}))\cong C_{o}^{{\mathbb{R}}}({\bf{R}}^{n}\times X,{\mathbb{R}})$
and
$(S^{n}A)^{+}\cong C(S^{n}\times X,{\mathbb{R}}).$
Taking $X=x_{0}$ a point, we have $C_{o}^{{\mathbb{R}}}(x_{0},{\mathbb{R}})={\mathbb{R}}$, and so
$S^{n}{\mathbb{R}}=C_{o}({\mathbb{R}}^{n},{\mathbb{R}})$
and
$(S^{n}{\mathbb{R}})^{+}=\real(S^{n},{\mathbb{R}})$
Next we come finally to desuspension. A construction similar to ours was first considered by Karoubi ([48], [49]) in the framework of Banach rings that is related to the development of negative algebraic $K$-theory of Bass (cf. Weibel [87] ).
Definition 4.4.
Let $A$ be a real $C^{*}$-algebra. The desuspension of $A$, denoted $VA$, is defined by
$VA=C_{o}^{{\mathbb{R}}}(i{\bf{R}})\otimes A$
and the n’th desuspension of $A$, denoted $V^{n}A$, is defined by
$V^{n}A=V(V^{n-1}A)\,\cong\,C_{o}^{{\mathbb{R}}}(i{\bf{R}})^{\otimes n}\otimes A\,\cong\,C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},\widetilde{\tau}^{\otimes n})\otimes A.$
We will need a little machinery in order to justify the use of the word “desuspension”. It will soon become clear.
Definition 4.5.
Suppose that $f_{0},f_{1}:A\to B$ are homomorphisms of real $C^{*}$-algebras. We say that $f_{0}$ is homotopic to $f_{1}$ (written $f_{0}\simeq f_{1}$) if there exists a homomorphism
$H:A\longrightarrow IB$
such that $p_{0}H=f_{0}$ and $p_{1}H=f_{1}$. We frequently write $f_{t}(a)=H(a)(t)$ and then think of $\{f_{t}\}$ as a continuous family of homomorphisms from $f_{0}$ to $f_{1}$. The first definition encodes what we mean by “continuous family.”
This definition of homotopy generalizes the definition of homotopy between compact spaces. If $f,g:X\to Y$ are based-point preserving continuous maps, then $f\simeq g$ if and only if the induced maps $f^{*},\,g^{*}:\real(Y)\to\real(X)$ are homotopic as real base-point preserving [^18] Suppose that $x_{o}$ is the base-point of $X$ and $\phi^{o}\in\real(X)$ is the function $\phi^{o}(x)=x_{o}$ for all $x$. Similarly define $\psi^{o}\in\real(Y)$. Then a function $\alpha:\real(Y)\to\real(X)$ is said to be base-point preserving if $\alpha(\psi^{o})=\phi^{o}$. This implies that $\alpha^{*}:X\to Y$ preserves base-points. When we say that $f^{*}\simeq g^{*}:\real(Y)\to\real(X)$ are base-point preserving homotopic, we mean that at each stage the homotopy must preserve base-points. $C^{*}$-algebra maps.
We note that homotopy is an equivalence relation, and we denote the set of homotopy classes of maps from $A$ to $B$ by $[A,B]$.
Recall that a topological space $X$ is contractible if the identity map $i:X\to X$ is homotopic to a constant map of the form $f(x)=x_{o}\,\,\,\forall x\in X$.
Definition 4.6.
A real $C^{*}$-algebra $A$ is contractible if the identity map $i:A\to A$ is homotopic to the trivial map $A\to A$ sending every element of $A$ to 0.
The simplest examples of contractible real $C^{*}$-algebras are those that come from contractible spaces. Here is a more interesting example.
Proposition 4.7.
If $A$ is a real $C^{*}$-algebra then $CA$, the cone of $A$, is contractible.
Proof.
Define $f_{s}:CA\to CA$ by
$f_{s}(\xi)(t)=\xi(st).$
This is well-defined since $f_{s}(\xi)(0)=\xi(0)=0$ for all $s$. Then
$f_{1}(\xi)(t)=\xi(t)\qquad{\text{so $f_{1}$ is the identity map $CA\to CA$ }}$
and
$f_{0}(\xi)(t)=\xi(0)=0\qquad{\text{so $f_{0}$ is the zero map.}}$
The basepoint of $A$ is the function $\phi:[0,1]\to A$ with $\phi(t)=0$ for all $t$. Then
$f_{s}(\phi)(t)=\phi(st)=0\,\,\forall t$
and hence the homotopy preserves basepoints.
Next we introduce pullbacks.
Definition 4.8.
Suppose given real $C^{*}$-algebras $A,B,C$ and maps
[…]
The pullback $D$
[…]
is defined by
$D=\{(a,b)\in A\oplus B:f(a)=g(b)\}$
with $\hat{f}$ and $\hat{g}$ the evident projections.
The notion of “pullback” occurs in many different settings.191919and different names. Sometimes it is called the fibre product or the Cartesian square, for instance. Indeed, later in this work we will need to take the pullback of a vector bundle! It is a categorical definition and enjoys the following universal property: in the notation of the diagram above, suppose that one is given a real $C^{*}$-algebra $L$ and maps
$r:L\longrightarrow A\qquad and\qquad s:L\longrightarrow B$
such that $fr=gs:L\to C$. Then there exists a unique map $k:L\to D$ such that
$r=\hat{g}k:L\longrightarrow D\longrightarrow A\qquad and\qquad s=\hat{f}k:L\longrightarrow D\longrightarrow B.$
Definition 4.9.
The mapping cone $Cf$ of a map $f:A\to B$ is defined to be the pullback
[…]
with associated short exact sequence
$0\longrightarrow SB\longrightarrow Cf\overset{\pi(f)}{\longrightarrow}A\longrightarrow 0.$
Definition 4.10.
The mapping cylinder $Mf$ of a map $f:A\to B$ is defined to be the pullback
[…]
with associated short exact sequence
$0\longrightarrow Cf\longrightarrow Mf\longrightarrow B\longrightarrow 0.$
Proposition 4.11.
Suppose given the short exact sequence and map of real $C^{*}$-algebras
[…]
Then the pullback
[…]
extends, up to isomorphism, to a commuting diagram of short exact sequences
[…]
Here is a very important example.
Definition 4.12.
The anonymous [^20] What else do you say when one’s identity is taken away? real Toeplitz algebra $\mathcal{T}_{{\mathbb{R}}0}$ is defined by the pullback diagram
[…]
and fits into the commuting diagram of short exact sequences
[…]
Definition 4.13.
A homotopy equivalence or equivalence $f:A\to B$ of real $C^{*}$-algebras is a map $f$ for which there exists a map $g:B\to A$ (called a homotopy inverse) such that $fg\simeq 1_{B}$ and $gf\simeq 1_{A}$. If there is a homotopy equivalence from $A$ to $B$ then we say that the two real $C^{*}$-algebras are homotopy equivalent and we write $A\simeq B$. This is an equivalence relation on real $C^{*}$-algebras.
This again is a straight-forward generalization of the notion for spaces. For example, it is easy to see that $D^{2}-(0,0)$, the unit disk minus the origin, is homotopically equivalent to a circle. At a much deeper level, we have the following result, which Schröder ([76] §1.5). attributes to J. Cuntz:
Theorem 4.14.
(Cuntz) Let $e_{1}:C^{{\mathbb{R}}}(iS^{1})\longrightarrow{\mathbb{R}}$ be evaluation at $1$ and let
$\pi:\mathcal{T}_{\mathbb{R}}\to{\mathbb{R}}$
be defined as the composition
$\mathcal{T}_{\mathbb{R}}\longrightarrow C^{{\mathbb{R}}}(iS^{1})\overset{e_{1}}{\longrightarrow}{\mathbb{R}}.$
Then $\pi:\mathcal{T}_{\mathbb{R}}\to{\mathbb{R}}$ is a homotopy equivalence with homotopy inverse $\iota:{\mathbb{R}}\to\mathcal{T}_{\mathbb{R}}$ the inclusion of the identity. Thus there is a split short exact sequence [^21] A short exact sequece $0\to A\to B\overset{p}{\to}C\to 0$ is said to split if there is a map $i:C\to B$ in the category such that $pi=id$. This implies that there is an isomorphism $B\cong A\oplus C$. Sequences need not split: for example, the short exact sequence of abelian groups $0\to{\mathbb{Z}}\overset{\times 2}{\longrightarrow}{\mathbb{Z}}\to{\mathbb{Z}}_{2}\to 0$ doesn’t split: $\mathbb{Z}$} is NOT isomorphic to ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$.
$0\to\mathcal{T}_{{\mathbb{R}}0}\longrightarrow\mathcal{T}_{\mathbb{R}}\overset{\pi}{\longrightarrow}{\mathbb{R}}\to 0$
with $\pi$ a homotopy equivalence. Further, for any real $C^{*}$-algebra $A$, there is a split short exact sequence
$0\to\mathcal{T}_{{\mathbb{R}}0}\otimes A\longrightarrow\mathcal{T}_{\mathbb{R}}\otimes A\overset{\pi\otimes 1}{\longrightarrow}A\to 0$
with $\pi\otimes 1$ a homotopy equivalence, so that there is a non-natural isomophism
$\mathcal{T}_{\mathbb{R}}\otimes A\cong\big(\mathcal{T}_{{\mathbb{R}}0}\otimes A\big)\oplus A.$
Next we move to cofibrations. For everything to do with this topic, please see the second author’s paper [74].
Definition 4.15.
. A map $p:A\to B$ of real $C^{*}$-algebras is a cofibration if whenever given a diagram
[…]
with real $C^{*}$-algebra $D$, a homotopy of maps
$h_{t}:D\longrightarrow B$
and a map
$H_{0};D\longrightarrow A$
with $pH_{0}=h_{0}$, then there exists a homotopy of maps
$H_{t}:D\to A$
such that the diagram
[…]
commutes for $0\leq t\leq 1$.
Remark 4.16.
The word ”cofibration” appears in classical algebraic topology. It has to do with a subspace $Y\subset X$ sitting nicely inside $X$. For example, the circle $S^{n}$ sits nicely inside the $n+1$ disk $D^{n+1}$ but the set of numbers $\{1,1/2,1/3,1/4,\dots\}$ does not sit nicely inside the closed unit interval.
We introduced the word ”cofibration” in [74] (1984) in the context of $C^{*}$-algebras so that the map
$f:X\to Y$
is a cofibration if and only if the map
$f^{*}:C(Y)\to C(X)$
is a cofibration. However, several authors have chosen to quote [74] but call this a ”fibration” (actually ”Schochet fibration”).  [^22] For example, Bunke [31] states the definition in much more elegant terms. He says that a map $A\to B$ is a Schochet fibration if the map $f_{*}:Hom(D,A)\longrightarrow Hom(D,B)$ is a Serre fibration of topological spaces for every $C^{*}$-algebra $D$. If $f:A\to B$ is a Schochet fibration with $J=Ker(f)$ then the associated short exact sequence $0\longrightarrow J\longrightarrow A\longrightarrow B\longrightarrow 0$ is a Schochet exact sequence. We will stick with ”cofibration.” The results in [74] are presented for complex $C^{*}$-algebras, but the parts having to do with cofibration apply to the real setting with no serious modification needed, so we will state facts but offer no proofs.
Here are some basic examples of cofibrations. We mention the item and then insert the reference in [74] where the facts are verified.
  1.
  2. (1)
    The map $p_{t}:IB\longrightarrow B$ is a cofibration for each $f\in[0,1]$. (Lemma 1.3)
  3. (2)
    In general, the map $A^{+}\to A^{+}/A\cong{\mathbb{R}}$ is not a cofibration.
  4. (3)
    If $p:A\to B$ is a cofibration and $f:D\to B$ is a map
[…]
    then the pullback $f^{*}(p):C\to D$ is a cofibration. (Prop. 1.5)
  5. (4)
    Given $f:A\to B$ then the mapping cone
    $Cf\overset{p}{\longrightarrow}A$
    is a cofibration. (Definition 2.1)
5. Homology Theories on Real $C^{*}$-algebras
In this section we give axioms for a homology theory for real $C^{*}$-algebras. We also point out some important and powerful consequences of these axioms. In truth, our main interest at the moment is $KO_{*}$. However, we believe that using the axioms makes the properties clearer. Certainly it makes the parallel with $K_{*}$ for complex $C^{*}$-algebras clearer and it makes the discussion of $KO$-theory with coefficients much easier, as we shall see.
In addition, we introduce axioms for a cofibre homology theory for real $C^{*}$-algebras. This is not immediately relevant for $KO_{*}$-theory but we believe that it belongs in this work as we shall explain.
Note: Nothing in this section uses Bott periodicity.
Definition 5.1.
A homology theory is a sequence $\{h_{n}\}$ of covariant functors from an admissible category [^23] Roughly speaking, a category is admissible if it is closed under the various operations (direct sum, direct limits, quotients, pullbacks, etc.) that we need to make our constructions. For instance, the natural numbers are closed under addition and multiplication but not under subtraction. For now, take the category to be all real $C^{*}$-algebras. One could also, for example, take the category of all real separable $C^{*}$-algebras. When working with homology theories on complex $C^{*}$-algebras we would use the category of complex $C^{*}$-algebras or the category of complex separable $C^{*}$-algebras, or something smaller such as the bootstrap category. $\mathcal{C}$} of real $C^{*}$-algebras to abelian groups which satisfies the following axioms:
Homotopy Axiom. Let $H:A\to IB$ be a homotopy from $f=p_{0}H$ to $g=p_{1}H$ in $\mathcal{C}$}. Then
$f_{*}=g_{*}:h_{n}(A)\longrightarrow h_{n}(B)\quad\text{for all n}.$
Exactness Axiom. Let
$0\longrightarrow J\overset{i}{\longrightarrow}A\overset{j}{\longrightarrow}A/J\longrightarrow 0$
be a short exact sequence in $\mathcal{C}$}. Then there is a map
$\partial:h_{n}(A/J)\longrightarrow h_{n-1}(J)$
and a long exact sequence
$\dots\to h_{n}(J)\overset{i_{*}}{\longrightarrow}h_{n}(A)\overset{j_{*}}{\longrightarrow}h_{n}(A/J)\overset{\partial}{\longrightarrow}h_{n-1}(J)\to\dots$
The map $\partial$ is natural with respect to morphisms of short exact sequences.
The axioms imply that if $A\cong\oplus_{i=1}^{k}A_{i}$ in $\mathcal{C}$ with $k$ finite then the natural map
$\oplus_{i=1}^{k}h_{n}(A_{i})\longrightarrow h_{n}(A)$
is an isomorphism. A further axiom is sometimes assumed. If it holds then the homology theory is said to be additive.
Additivity Axiom Let $A=\oplus_{i=1}^{\infty}A_{i}$ in $\mathcal{C}$}. Then the natural maps $A_{j}\to A$ induce an isomorphism
$\oplus_{i}h_{n}(A_{i})\overset{\cong}{\longrightarrow}h_{n}(A).$
for each $n$.
Theorem 5.2.
([74], Theorem 5.1) Suppose that $h_{*}$ is an additive homology theory and we are given a sequence of real $C^{*}$-algebras
$A_{1}\longrightarrow A_{2}\longrightarrow\dots\longrightarrow A_{n}\longrightarrow\dots$
[…]
Corollary 5.3.
Suppose that $h_{*}$ is an additive homology theory and we are given a sequence of real $C^{*}$-algebras
$A_{1}\longrightarrow A_{2}\longrightarrow\dots\longrightarrow A_{n}\longrightarrow\dots$
[…]
$h_{*}(A_{j}\otimes B)\to h_{*}(A\otimes B)$
induce an isomorphism
[…]
Definition 5.4.
Suppose that $A$ is a real $C^{*}$ -algebra. We say that $A$ is $h$-contractible if $h_{n}(A)=0$ for all $n$.
For example, the cone $CA$ of a real $C^{*}$-algebra is $h$-contractible for any homology theory $h$.
Proposition 5.5.
Suppose that $h_{*}$ is a homology theory on real $C^{*}$-algebras. Then the boundary map induces an isomorphism
$\partial:h_{n}(A)\overset{\cong}{\longrightarrow}h_{n-1}(SA)$
for all $n$.
Proof.
There is a short exact sequence
$0\longrightarrow SA\longrightarrow CA\longrightarrow A\longrightarrow 0$
with associated homology long exact sequence
$\dots\rightarrow h_{n}(CA)\longrightarrow h_{n}(A)\overset{\partial}{\longrightarrow}h_{n-1}(SA)\longrightarrow h_{n-1}(CA)\rightarrow\dots.$
Since $h_{*}(CA)=0$, the long exact sequence degenerates to short sequences of the form
$\dots\rightarrow 0\longrightarrow h_{n}(A)\overset{\partial}{\longrightarrow}h_{n-1}(SA)\longrightarrow 0\rightarrow\dots$
which implies the result ∎
Recall:
  • $\mathcal{T}_{\mathbb{R}}$ is the real Toeplitz algebra
  • $\mathcal{T}_{{\mathbb{R}}0}$ is the real anonymous Toeplitz algebra
  • $VA=C_{o}^{{\mathbb{R}}}(i{\bf{R}})\otimes A$ is the desuspension of AA
Theorem 5.6.
Suppose that $h_{*}$ is a homology theory on real $C^{*}$-algebras, and let $A$ be an arbitrary real $C^{*}$-algebra. Then
  1. (1)
    The homotopy equivalence $\pi:\mathcal{T}_{\mathbb{R}}\overset{\simeq}{\longrightarrow}{\mathbb{R}}$ extends to a homotopy equivalence
    $\pi\otimes 1:\mathcal{T}_{\mathbb{R}}\otimes A\overset{\simeq}{\longrightarrow}A.$
  2. (2)
    There is a natural isomorphism
    $(\pi\otimes 1)_{*}:h_{*}(\mathcal{T}_{\mathbb{R}}\otimes A)\overset{\cong}{\longrightarrow}h_{*}(A).$
  3. (3)
    The real $C^{*}$-algebra $\mathcal{T}_{{\mathbb{R}}0}\otimes A$ is $h_{*}$-contractible.
  4. (4)
    We have shown in the previous section that there is a short exact sequence of the form
    $0\longrightarrow{\mathcal{K}}_{R}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\longrightarrow V{\mathbb{R}}\longrightarrow 0.$
    This produces, for any $A$, the short exact sequence
    $0\longrightarrow A\otimes{\mathcal{K}}_{R}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\otimes A\longrightarrow VA\longrightarrow 0.$
    The long exact homology sequence associated to this short exact sequence degenerates to isomorphisms
    $\partial:h_{n}(VA)\overset{\cong}{\longrightarrow}h_{n-1}(A).$
Proof.
Part (1) is immediate from the definitions, and of course a homotopy equivalence of $C^{*}$-algebras induces an isomorphism on homology, establishing Part (2).
For Part (3) we recall that there is an exact sequence
$0\to\mathcal{T}_{{\mathbb{R}}0}\longrightarrow\mathcal{T}_{{\mathbb{R}}}\overset{\pi}{\longrightarrow}{\mathbb{R}}\to 0$
and when we tensor with $A$ the resulting sequence
$0\to\mathcal{T}_{{\mathbb{R}}0}\otimes A\longrightarrow\mathcal{T}_{{\mathbb{R}}}\otimes A\overset{\pi\otimes 1}{\longrightarrow}A\to 0$
remains exact. The associated homology long exact sequence looks in part like
$h_{n+1}(\mathcal{T}_{{\mathbb{R}}}\otimes A)\overset{\cong}{\longrightarrow}h_{n+1}(A)\overset{\partial}{\longrightarrow}h_{n}(\mathcal{T}_{{\mathbb{R}}0}\otimes A)\to h_{n}(\mathcal{T}_{{\mathbb{R}}}\otimes A)\overset{\cong}{\longrightarrow}h_{n}(A)$
and this implies [^24] This is a matter of pure algebra: suppose given an exact sequence of abelian groups $G_{1}\overset{\alpha_{1}}{\longrightarrow}G_{2}\overset{\alpha_{2}}{\longrightarrow}G_{3}\overset{\alpha_{3}}{\longrightarrow}G_{4}\overset{\alpha_{4}}{\longrightarrow}G_{5}$ with $\alpha_{1}$ and $\alpha_{4}$ isomorphisms. Then $Ker(\alpha_{2})\cong Im(\alpha_{1})=G_{2}$ since $\alpha_{1}$ is onto, which says that $\alpha_{2}\equiv 0$. Similarly, $Im(\alpha_{3})=Ker(\alpha_{4})=0$ since $\alpha_{4}$ is 1-1, which says that $\alpha_{3}\equiv 0$. Then $\alpha_{2}=\alpha_{3}\equiv 0$ implies that $G_{3}=0$. that $h_{n}(\mathcal{T}_{{\mathbb{R}}0}\otimes A)=0$ for all $n$, establishing Part (3).
For Part (4) we simply write out the associated homology long exact sequence and then observe that, by Part (3), every third term vanishes. We are thus left with
$0\longrightarrow h_{n}(VA)\overset{\partial}{\longrightarrow}h_{n-1}(A)\longrightarrow 0$
and this implies that $\partial$ is an isomorphism in each dimension.
The following corollary describes both suspension (which also is available in the complex $K$-theory situation) and desuspension (which is totally without precedent in the complex $K$-theory world).
Corollary 5.7.
For any homology theory $h_{*}$ on real $C^{*}$-algebras and for any real $C^{*}$-algebra $A$, there is a natural suspension isomorphism
$h_{n}(A)\cong h_{n-1}(SA)$
and a natural desuspension isomorphism
$h_{n}(A)\cong h_{n+1}(VA).$
Next we turn our attention to the creation of new homology theories. This technique has been used before, for example in [75] to define $K_{*}(A;{\mathbb{Z}}_{n})$, complex $K$-theory with coefficients ${\mathbb{Z}}_{n}$ from $K_{*}(-)$.
Proposition 5.8.
Suppose that $h_{*}$ is a homology theory and $N$ is a real $C^{*}$-algebra such that $N_{\mathbb{C}}$ is nuclear. Then $h_{*}(-\otimes N)$ is a homology theory as well. If $h_{*}$ is additive then so is $h_{*}(-\otimes N)$
Proof.
We need to check the homotopy and exactness axioms.
Suppose first that $H:A\to IB$ is a homotopy from $f=p_{0}H$ to $g=p_{1}H$. Then there is a natural continuous map
$H_{N}:A\otimes N\longrightarrow I(B\otimes N)$
and this implies that $f\otimes I_{N}$ is homotopic to $g\otimes I_{N}$. Hence
$(f\otimes I_{N})_{*}=(g\otimes I_{N})_{*}:h_{*}(A\otimes N)\longrightarrow h_{*}(B\otimes N)$
as required. So the homotopy axiom does hold.
Exactness is easy. Suppose given a short exact sequence
$0\to J\longrightarrow A\longrightarrow A/J\to 0.$
Tensor the sequence with $N$ and the resulting sequence
$0\to J\otimes N\longrightarrow A\otimes N\longrightarrow(A/J)\otimes N\to 0$
is still exact, since $N_{\mathbb{C}}$ is exact. Apply $h_{*}$ and obtain a long exact sequence.
Finally, suppose that $h_{*}$ is additive and given $A=\oplus_{1}^{\infty}A_{j}$. Then the natural maps $A_{j}\to A$ yield maps $A_{j}\otimes N\to A\otimes N$ and hence to
$h_{*}(A_{j}\otimes N)\longrightarrow h_{*}(A\otimes N)$
so the additivity of $h_{*}$ implies that there is a natural isomorphism
$\oplus_{j}h_{n}(A_{j}\otimes N)\overset{\cong}{\longrightarrow}h_{n}(A\otimes N).$
Corollary 5.9.
If $h_{*}$ is a homology theory on a category of real $C^{*}$-algebras then so is the theory $h_{*}(-\otimes{\mathbb{H}})$.
This will be of interest when we take $h_{*}=KO_{*}$.
Next we discuss cofibre homology theories. These are thoroughly discussed in the complex setting in [74] and so, as before, we simply give specific references to that paper.
Definition 5.10.
([74] p. 437) A cofibre homology theory is a sequence $\{h_{n}\}$ of covariant functors from an admissible category $\mathcal{C}$ of real $C^{*}$-algebras to abelian groups which satisfies the following axioms:
Homotopy Axiom (same as for homology theories)
Suspension Axiom. For all $A$, the natural suspension map
$\sigma_{A};h_{n}(A)\longrightarrow h_{n-1}(SA)$
is an isomorphism.
Cofibre Exactness Axiom. If $f:A\to B$ is a cofibration with mapping cone $Cf$ then for all $n$ the sequence
$h_{n}(Cf)\longrightarrow h_{n}(A)\longrightarrow h_{n}(B)$
is exact.
The exactness and homotopy axioms of a homology theory imply the cofibration exactness axiom ([74], Prop. 3.7), and hence any homology theory is a cofibre homology theory. The converse is false: see [74], Remark 8.6.
Here are two basic properties of cofibre homology theories.
Proposition 5.11.
([74], 3.7 and 3.8) Suppose that $h_{*}$ is a cofibre homology theory on an admissible category of real $C^{*}$-algebras. Then:
  1. (1)
    If $f:A\to B$ is a cofibration with $J=Ker(f)$ then there is a natural long exact sequence
    $\dots h_{n}(J)\longrightarrow h_{n}(A)\overset{f_{*}}{\longrightarrow}h_{n}(B)\longrightarrow h_{n-1}(J)\dots$
  2. (2)
    if $A_{1},\dots A_{n}$ are in the category then for each $k$ the natural map
    $\oplus_{j=1}^{n}h_{k}(A_{j})\longrightarrow h_{k}(A_{1}\oplus\dots\oplus A_{n})$
    is an isomorphism.
Remark 5.12.
([74] 8.5) Here is a family of examples of cofibre homology theories which are generally not homology theories. All $C^{*}$-algebras are in an admissible category. Fix some $C^{*}$-algebra $D$. Define
[…]
Then $D_{*}(A)$ is a cofibre homology theory.
Remark 5.13.
Why discuss cofibre homology theories if they seem unrelated to $KO$-theory? The basic point to understand is that, in classical algebraic topology, most homology and cohomology theories are in fact only cofibre theories. Topological $K$ and $KO$-theories are actually quite exceptional. We anticipate that in years to come that there will arise other theories to help us understand real $C^{*}$-algebras. Given the precedent in classical algebraic topology, there is no reason to suppose that these theories will satisfy the axioms that we have stated for homology theories, but they will be cofibre theories. It is already the case that cofibrations252525known there as ”Schochet fibrations” as discussed in Section 4. have appeared in several papers, for instance in Bunke [31].
6. $KO_{*}$-theory for real $C^{*}$-Algebras: Definitions and Basic Properties
Note: Nothing in this section uses Bott periodicity in any manner.
Sometimes matters of taste intervene in mathematical research. We are at that point. $K$-theory can be developed in terms of projections or in terms of projective modules. Analysts tend to like projections and algebraists and topologists like projective modules. Here’s how the connection goes.
Let $A$ be a unital ring (for example, a unital real $C^{*}$-algebra). A finitely-generated free $A$-module is a module of the form $A\oplus\dots\oplus A$, with a finite number (say $k$) of copies of $A$. Think of this as the identity function
$A\oplus\dots\oplus A\overset{1_{k}}{\longrightarrow}A\oplus\dots\oplus A$
which in turn we may regard as the identity matrix $1_{k}\in M_{k}(A)$. Conversely, if we start with $1_{k}\in M_{k}(A)$, regard it as a function and take its range, then we obtain $A\oplus\dots\oplus A$. Along the way we have chosen a basis.
A projection $p$ is an element of $M_{k}(A)$ with $p=p^{*}=p^{2}$. It too induces a map
$A^{k}\overset{p}{\longrightarrow}A^{k}$
but in general the map will not be 1-1 or onto. Let us denote by $P$ the image of the projection. Then $P$ is a finitely-generated $A$-module. The orthogonal complement of $p$ (call it $q$) has image that is orthogonal to $P$ and $P\oplus Q\cong A^{k}$. This is exactly what is required to say that $P$ is a finitely-generated projective $A$-module. Again, the process may be reversed.
The upshot is that projections in finite-dimensional matrix algebras over $A$ are exactly the same as finitely-generated projective $A$-modules. When developing $K$-theory in terms of projective modules, we regard two abstractly isomorphic finitely-generated projective $A$-modules as equivalent. When developing $K$-theory in terms of projections (elements in $A$ or in $M_{n}(A)$) we also need to consider certain equivalence classes.
Definition 6.1.
Suppose that $A$ is a unital real $C^{*}$-algebra. Then we define three equivalence relations on the set of projections of $A$ as follows. Let $p$ and $q$ be projections in $A$.
  1. (1)
    $p$ and $q$ are unitarily equivalent, written $p\equiv_{u}q$, if there exists a unitary [^26] A unitary is an element which satisfies $uu^{*}=u^{*}u=1$. $u\in A$ such that $q=upu^{*}$.
  2. (2)
    $p$ and $q$ are Murray - von Neumann equivalent, written $p\equiv_{m}q$, if there exists a partial isometry [^27] A partial isometry is an element $a$ for which $a^{*}a$ is a projection. $a\in A$ such that $p=aa^{*}$ and $a^{*}a=q$.
  3. (3)
    $p$ and $q$ are homotopic, written $p\equiv_{h}q$, if there exists a continuous path $R_{t}$ of projections in $A$ such that $R_{0}=p$ and $R_{1}=q$.
Proposition 6.2.
For any unital real $C^{*}$-algebra $A$, the relations $\equiv_{u}$, $\equiv_{m}$, and $\equiv_{h}$ are each equivalence relations. Futhermore, they are related by the following implications for any projections $p,q\in A$:
  • if $p\equiv_{u}q$ then $p\equiv_{m}q$,
  • if $p\equiv_{h}q$ then $p\equiv_{u}q$,
  • if $p\equiv_{m}q$ then $\begin{pmatrix}p&0\\ 0&0\end{pmatrix}\equiv_{u}\begin{pmatrix}q&0\\ 0&0\end{pmatrix}$ in $M_{2}(A)$,
  • if $p\equiv_{u}q$ then $\begin{pmatrix}p&0\\ 0&0\end{pmatrix}\equiv_{h}\begin{pmatrix}q&0\\ 0&0\end{pmatrix}$ in $M_{2}(A)$,
See Section 4.2 of [14] or Section 2.2 of [69] for details. The definition of $KO_{0}(A)$ in what follows is based on unitary equivalence, but because of Proposition 6.2 we could use either of the other equivalence relations with the same result.
Definition 6.3.
Suppose that $A$ is a unital real $C^{*}$-algebra. Let $\mathcal{P}(A)$ denote the set of unitary equivalence classes of projections in $A$. Let $M_{n}(A)=A\otimes M_{n}({\mathbb{R}})$ and $\mathcal{P}_{n}(A)=\mathcal{P}(M_{n}(A))$. The inclusion $M\to M\oplus 0$ of $M_{n}(A)$ into $M_{n+1}(A)$ induces inclusions
$\mathcal{P}_{n}(A)\subset\mathcal{P}_{n+1}(A).$
Let $\mathcal{P}_{\infty}(A)$ be the union of the $\mathcal{P}_{n}(A)$.
Note that if $p\in M_{r}(A)$ and $q\in M_{s}(A)$ then we can form $p\oplus q\in M_{r+s}(A)$ by putting $p$ and $q$ into the upper left and lower right corners of a square $r+s$ matrix respectively, with 0’s in the upper right and lower left. We do this even if $p=q$. This gives $\mathcal{P}_{\infty}(A)$ the structure of an abelian semigroup. [^28] An abelian semigroup is a set with a binary operation that we denote ++ that is associative and commutative. There are bad classic jokes connected with this concept. What is purple and commutes? An abelian grape. What is lavender and commutes? An abelian semigrape.
If $H$ is an abelian semigroup then there is an associated abelian group $Groth(H)$, the Grothendieck group of $H$. It may be defined as the quotient of $H\times H$ under the equivalence relation
$(h_{1},h_{2})\sim(k_{1},k_{2})\quad\text{whenever $\exists z\in H$ with}\quad h_{1}+k_{2}+z=h_{2}+k_{1}+z$
This can also be understood as taking formal differences of elements of $H$. The classic example is the natural numbers $\mathbb{N}$} with $Groth({\mathbb{N}})={\mathbb{Z}}$. In terms of formal differences, we think of 3-3 as [4][7][4]-[7]. Of course we want that to be the same as [5][8][5]-[8] and the equivalence relation does just that, since $4+8=7+5$.
The Grothendieck construction is natural and is a covariant functor from abelian semigroups to abelian groups. If $H$ is an abelian semiring (all the ring axioms except elements may not have additive inverses) then $Groth(H)$ is a commutative ring. We will see this happen when we restrict attention to $\mathcal{P}_{\infty}(C^{\mathbb{R}}(X))$.
Definition 6.4.
Suppose that $A$ is a real unital $C^{*}$-algebra. Define
$KO_{0}(A)=Groth(\mathcal{P}_{\infty}(A)).$
Thus an element of $KO_{0}(A)$ may be represented as a formal difference [p][q][p]-[q], where $p$ and $q$ are projections in matrix rings over $A$, with the following relations:
  1. (1)
    For any orthogonal matrix $u$,
    $p]=[upu^{-1}].$
  2. (2)
    For any projection $p$,
    $[p]+[0]=[p].$
  3. (3)
    $p]+[q]=[p\oplus q].$
Given any homomorphism $A\to B$, there is induced a homomorphism
$f_{*}:KO_{0}(A)\longrightarrow KO_{0}(B)$
where the map is given on generators by
$f_{*}([p])=[f(p)].$
It is easy to see that
  1. (1)
    If $i:A\to A$ is the identity map then $i_{*}:KO_{0}(A)\to KO_{0}(A)$ is the identity map.
  2. (2)
    Given real $C^{*}$-algebra maps
    $A\overset{f}{\longrightarrow}B\overset{g}{\longrightarrow}C,$
    then
    $(gf)_{*}=g_{*}f_{*}:KO_{0}(A)\longrightarrow KO_{0}(C).$
Thus $KO_{0}$ is a covariant functor from unital $C^{*}$-algebras and maps to abelian groups.
Remark 6.5.
It is easy to be tempted to think that one can use a similar procedure to produce a ring structure on $KO_{0}(A)$. The problem is the following: if $P$ and $Q$ are left modules over a non-commutative ring $A$, then $P\otimes_{A}Q$ is an abelian group but not necessarily an $A$- module. Thus $Groth(\mathcal{P}_{\infty}(A))$ does not have a product and the construction grinds to a halt. If $A$ is commutative then the product does make sense and we discuss this in future sections.
We wish to make the relationship between $KO_{0}(A)$ and $K_{0}(A)$ clear from the beginning, and so we take a short detour.
Definition 6.6.
Suppose that $A$ is a unital ring. Then we define the algebraic $K$-theory group $K_{0}^{alg}(A)$ by forming the category $\mathcal{P}_{\infty}(A)$ of isomorphism classes of finitely generated projective left $A$-modules, or, equivalently, the category of equivalence classes of projections in matrix rings over $A$, and then defining
$K_{0}^{alg}(A)=Groth(\mathcal{P}_{\infty}(A)).$
Note that this definition makes sense for any unital ring. There is a $K_{1}^{alg}(A)$ and (by high-tech methods) $K_{j}^{alg}(A)$ for all $j$, but these functors are NOT periodic in general.
If $A$ is a real $C^{*}$-algebra then
$KO_{0}(A)\cong K_{0}^{alg}(A)$
simply by comparing definitions. If $A$ is a complex $C^{*}$-algebra then
$K_{0}(A)\cong K_{0}^{alg}(A).$
This is true for $K_{0}$ but not for other algebraic $K_{j}$. For example, the topological $K_{1}$ of a $C^{*}$-algebra is a quotient of the algebraic $K_{1}$ of the algebra.
The simplest case is, of course, $A={\mathbb{R}}$. Then $M_{n}({\mathbb{R}})$ is simply the ring of $n\times n$ real matrices, a projection $p$ is simply an $n\times n$ matrix with $p=p^{2}=p^{*}$. If $p\in M_{n}({\mathbb{R}})$ is a projection then there is an orthogonal matrix $u$ such that
$upu^{-1}=diag[1,1,1,\dots 1,0,0,0,\dots 0],$
the diagonal matrix with $j$ 1’s down the main diagonal and njn-j 0’s following. This has rank $j$ with $0\leq j\leq n$ and defines a dimension function
$dim:\mathcal{P}_{n}({\mathbb{R}})\longrightarrow\{0,1,\dots,n\}.$
These dimension functions coalesce to form an isomorphism
$dim:KO_{0}({\mathbb{R}})\overset{\cong}{\longrightarrow}{\mathbb{Z}}.$
(Exactly the same argument gives $K_{0}({\mathbb{C}})\cong{\mathbb{Z}}.$ )
So far we have defined $KO_{0}(A)$ only for unital $A$. Note that if $A$ is a real $C^{*}$-algebra with unitization $A^{+}$ then there is a canonical short exact sequence
$0\longrightarrow A\longrightarrow A^{+}\overset{\pi}{\longrightarrow}(A^{+}/A)\,\cong{\mathbb{R}}\longrightarrow 0.$
Definition 6.7.
For any real $C^{*}$-algebra $A$, we define
$KO_{0}(A)=ker\big[KO_{0}(A^{+})\overset{\pi}{\longrightarrow}KO_{0}({\mathbb{R}})\cong{\mathbb{Z}}\big]$
[…]
$0\longrightarrow KO_{0}(A)\longrightarrow KO_{0}(A^{+})\longrightarrow{\mathbb{Z}}\longrightarrow 0.$
which splits unnaturally, and hence there is an unnatural splitting
$KO_{0}(A^{+})\cong KO_{0}(A)\oplus{\mathbb{Z}}.$
Using exactly the same idea, for any complex $C^{*}$-algebra $A$, we define
$K_{0}(A)=ker\big[K_{0}(A^{+})\overset{\pi}{\longrightarrow}K_{0}({\mathbb{C}})\cong{\mathbb{Z}}\big].$
so there is a natural split short exact sequence
$0\longrightarrow K_{0}(A)\longrightarrow K_{0}(A^{+})\longrightarrow{\mathbb{Z}}\longrightarrow 0$
and an (unnatural) isomorphism
$K_{0}(A^{+})\cong K_{0}(A)\oplus{\mathbb{Z}}.$
A homomorphism $f:A\to B$ induces a homomorphism $f_{*}:KO_{0}(A)\to KO_{0}(B)$ and $KO_{0}$ is a covariant functor from the category of real $C^{*}$-algebras and homomorphisms to abelian groups. The same idea carries through in the complex setting. [^30] Karoubi conjectured in the late 1970’s that if $A$ is a complex $C^{*}$-algebra then the algebraic and $C^{*}$-algebraic $K$-theory groups of $A\otimes{\mathcal{K}}$ were isomorphic, and similarly for the real setting. Brown and the second author proved this for $A={\mathbb{C}}$ in 1976 [30]. The general case was established by Suslin and Wodzicki (1990 and 1994)[54]. To get a feeling for this, look at [30]. It turns out that $K_{1}^{alg}({\mathcal{K}})\cong K_{1}({\mathcal{K}})=0.$ It is easy to show that $K_{1}({\mathcal{K}})=0$ but quite difficult to show that $K_{1}^{alg}({\mathcal{K}})=0.$
Remark 6.8.
For each real $C^{*}$-algebra $A$ there is a natural complexification map
$c:KO_{0}(A)\longrightarrow K_{0}(A_{\mathbb{C}})$
which produces a natural transformation of theories. To start with, suppose that $P$ is a finitely generated projective $A$-module with associated isomorphism
$P\oplus Q\cong A^{k}.$
By complexifying we obtain the isomorphism
$P_{\mathbb{C}}\oplus Q_{\mathbb{C}}\cong(A_{\mathbb{C}})^{k}$
and hence $P_{\mathbb{C}}$ is a finitely generated projective $A_{\mathbb{C}}$-module. This induces a natural map
$c:KO_{0}(A)\longrightarrow K_{0}(A_{\mathbb{C}})$
for $A$ unital by sending $[P]$ to $P_{\mathbb{C}}]$. The map is obviously a natural homomorphism. As the definition of the $KO$ and $K$-theory groups for nonunital algebras are essentially identical, it is easy to show that
$c:KO_{0}(A)\longrightarrow K_{0}(A_{\mathbb{C}})$
is well-defined and well-behaved for all real $C^{*}$-algebras.
We can describe this map in a different way. For any real $C^{*}$-algebra there is an inclusion map, which is a homomorphism $A\hookrightarrow A_{\mathbb{C}}$. If we regard $A_{\mathbb{C}}$ as a real $C^{*}$-algebra this induces a natural transformation
$c:KO_{0}(A)\longrightarrow KO_{0}(A_{\mathbb{C}}).$
A moment’s thought about projections or about finitely-generated projective modules will convince the reader that
$KO_{0}(A_{\mathbb{C}})\cong K_{0}(A_{\mathbb{C}})$
and so the composition gives
$c:KO_{0}(A)\longrightarrow K_{0}(A_{\mathbb{C}}).$
Here’s an important example: $A=C_{o}^{{\mathbb{R}}}({\mathbb{R}})$ so that $A^{+}=\real(S^{1})$. Then there is a natural short exact sequence
$0\longrightarrow C_{o}^{{\mathbb{R}}}({\bf{R}})\longrightarrow\real(S^{1})\longrightarrow{\mathbb{R}}\longrightarrow 0$
and we shall see that this induces a short exact sequence
[…]
We move on now to the higher $KO$-groups.
Definition 6.9.
If $A$ is a unital real $C^{*}$-algebra, then define $GL_{j}(A)$ to be the topological group of invertible matrices in $M_{j}(A)$. If $A$ is not unital, define
$GL_{j}(A)=\{x\in GL_{j}(A^{+}):x\equiv 1_{j}\,\,mod\,\,M_{j}(A)\}.$
Then $GL_{j}(A)$ is a closed normal subgroup of $GL_{j}(A^{+})$. We may embed
$GL_{j}(A)\to GL_{j+1}(A)$
by $x\to diag(x,1)$ and let
[…]
It is easy to see that $GL_{j}(-)$ is a covariant functor from real $C^{*}$-algebras to topological groups.
We note that it is straight-forward to prove313131using CW-complexes, which we will discuss in §9. that the natural maps
$GL_{j}(A)\longrightarrow GL_{\infty}(A)$
induce an isomorphism
[…]
There are two ways to define $KO_{n}(A)$ for $n=1,2,3,\dots$ and both ways have their pros and cons.
Definition 6.10.
FIRST DEFINITION For $n\geq 1$, define [^32] We always take the identity of the topological group to be the basepoint for homotopy purposes, and we always use based spaces and base-point preserving maps when dealing with compact spaces. In general $\pi_{n}(X)$ is an abelian group for $n>1$ and only a group for $n=1$. However, $\pi_{1}$ of a topological group is always abelian.
$KO_{n}(A)=\pi_{n-1}(GL_{\infty}(A)).$
If $f:A\to B$ is a homomorphism of real $C^{*}$-algebras then it induces a map
$f:GL_{\infty}(A)\to GL_{\infty}(B)$
and hence
$f_{*}:KO_{n}(A)\to KO_{n}(B).$
Then each $KO_{n}(-)$ is a covariant functor from real $C^{*}$-algebras to abelian groups. We write $KO_{*}(A)$ for the graded abelian group $\{KO_{n}(A):n\geq 0\}$. The same construction for complex $C^{*}$-algebras yields $K_{j}(A)$.
For many purposes it is convenient to use the fact that $O_{n}(A)$, the group of orthogonal $n\times n$ matrices, is a strong deformation retraction of $GL_{n}(A)$. [^33] For example, the set $O_{1}({\mathbb{R}})=\{\pm 1\}$ is a strong deformation retraction of $GL_{1}({\mathbb{R}})={\mathbb{R}}-\{0\}$, and $O_{2}$ (rotations and reflections of the plane - topologically two circles) is a strong deformation retraction of $GL_{2}({\mathbb{R}})$. The compact Lie group $O_{n}({\mathbb{R}})$ is naturally contained in $GL_{n}({\mathbb{R}})$ and there is a strong deformation retraction343434also known as Gram-Schmidt orthogonalization which deforms $GL_{n}({\mathbb{R}})$ down to $O_{n}({\mathbb{R}})$. Thus the two spaces are homotopy equivalent. This implies that $O_{\infty}({\mathbb{R}})$ has the same homotopy type as $GL_{\infty}({\mathbb{R}})$, and an analogous argument demonstrates that $O_{\infty}(A)$ has the same homotopy type as $GL_{\infty}(A)$ for any real (or complex!) $C^{*}$-algebra $A$. Thus we may as well define
$KO_{n}(A)=\pi_{n-1}(O_{\infty}(A))$
and this is the way it is done in many places in the literature.
Definition 6.11.
SECOND DEFINITION For $n\geq 1$, define
$KO_{n}(A)=KO_{0}(S^{n}A).$
If $f:A\to B$ is a homomorphism of real $C^{*}$-algebras then it induces
$f:S^{n}A\longrightarrow S^{n}B$
and hence
$f_{*}:KO_{n}(A)\to KO_{n}(B).$
Then each $KO_{n}(-)$ is a covariant functor from real $C^{*}$-algebras to abelian groups. We write $KO_{*}(A)$ for the graded abelian group $\{KO_{n}(A):n\geq 0\}$.
Theorem 6.12.
For all real unital $C^{*}$-algebras $A$,
$\pi_{n-1}(GL_{\infty}(A))\cong KO_{0}(S^{n}A)$
and hence the two definitions of $KO_{n}(A)$ agree for all $A$.
This result has been “well-known” for fifty years or so, but it is remarkably difficult to put together a rigorous proof. [^35] Jon Rosenberg (private communication) suggested the following approach. First, define bigraded $K$-theory via Clifford algebras and use the periodicity for Clifford algebras to show that $KO_{*}(A)$ is periodic of period 8. The resulting homology theory may be represented by an infinite loop space, and you then show that this infinite loop space may be taken to be determined by $GL_{\infty}(A)$. For information on Clifford algebras and periodicity please see future sections. The technical details of this approach are well beyond the scope of this work . Schröder (cf. Schröder, [76] Theorem 1.4.6 ) briefly discusses the proof of this theorem and then refers to Karoubi’s book [52] III, 1.6 (it is actually 1.11). The book does prove periodicity for $K_{*}$ but not for $KO_{*}$. For a detailed proof of that you have to go to an earlier paper of Karoubi, namely [45].
Theorem 6.13.
$KO_{*}$ is an additive homology theory on the category of real $C^{*}$-algebras.
Proof.
Schröder [76] establishes the homotopy axiom (in somewhat greater generality) as Proposition 1.4.2 and the additivity axiom as Proposition 1.4.8. ∎
Remark 6.14.
We previously defined the complexification natural transformation
$c:KO_{0}(A)\longrightarrow K_{0}(A_{\mathbb{C}})$
for all real $C^{*}$-algebras $A$. Using the second definition (and its analog for complex $K$-theory), it is easy to show that this natural transformation extends to a natural transformation
$c:KO_{n}(A)\longrightarrow K_{n}(A_{\mathbb{C}})$
of homology theories.
We note a very important and immediate consequence of this result. It justifies our use of “desuspension” to describe $VA$.
Corollary 6.15.
Suppose that $A$ is a real $C^{*}$-algebra. Then:
  1. (1)
    the natural short exact sequence
    $0\longrightarrow SA\longrightarrow CA\longrightarrow A\longrightarrow 0$
    induces a natural suspension isomorphism
    $KO_{n}(A)\cong KO_{n-1}(SA)$
    and, more generally,
    $KO_{n}(A)\cong KO_{n-k}(S^{k}A).$
  2. (2)
    There is a natural desuspension isomorphism [^36] The boundary map in the $KO$-theory short exact sequence associated to the Toeplitz extension $0\longrightarrow A\otimes{\mathcal{K}}_{R}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\otimes A\longrightarrow VA\longrightarrow 0$ has the form $KO_{n+1}(VA)\overset{\cong}{\longrightarrow}KO_{n}(A\otimes{\mathcal{K}}_{R})\cong KO_{n}(A)$ which we write right to left to make the comparison with suspension more dramatic.
    $KO_{n}(A)\cong KO_{n+1}(VA)$
    and, more generally,
    $KO_{n}(A)\cong KO_{n+k}(V^{k}A).$
Proof.
This is immediate from Corollary 5.7. ∎
Of course at the moment this corollary holds only when all of the subscripts are non-negative. We will remedy that defect in the next section.
Corollary 6.16.
For any real $C^{*}$-algebra $A$, there are natural isomorphisms
$KO_{n}(A)\cong KO_{n-1}(SA)\cong KO_{n}(VSA)$
and similarly
$KO_{n}(A)\cong KO_{n+1}(VA)\cong KO_{n}(SVA)$
Corollary 6.17.
For all real $C^{*}$-algebras $A$ and integers $k\geq 1$ and $n\geq 0$,
$KO_{n}(A)\cong KO_{n}(M_{k}(A))\cong KO_{n}(A\otimes{\mathcal{K}}_{\mathbb{R}})$
Proof.
It suffices to consider the case $n=0$, since by the previous corollary, $KO_{n}(A)\cong KO_{0}(S^{n}A)$. The group $KO_{0}(A)$ is generated by projections in matrix rings over $A$, and we use the simple observation that
$M_{i}(M_{j}(A))\cong M_{ij}(A)$
to show that
$\mathcal{P}(A)_{\infty}=\mathcal{P}_{\infty}(M_{k}(A))=\mathcal{P}_{\infty}(A\otimes\mathcal{K}_{\mathbb{R}}).$
Theorem 6.18.
(classical) Here are the $KO$-groups of the real $C^{*}$-algebras $\mathbb{R}$}, $\mathbb{H}$} and $\mathbb{C}$}.
n= 0 11 22 33 44 55 66 77
$KO_{n}({\mathbb{R}})$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} 0 0 0
$KO_{n}({\mathbb{H}})$ $\mathbb{Z}$} 0 0 0 $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0
$KO_{n}({\mathbb{C}})$ $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0
Let’s look at some examples. Steenrod [80] §22 is an excellent reference for concrete calculations. Of course he was calculating homotopy groups of classical Lie groups, as $K$-theory had not been invented then.
  1. (1)
    The group $KO_{0}({\mathbb{R}})={\mathbb{Z}}$. We discussed this previously: a projection in $M_{k}({\mathbb{R}})$ is determined up to equivalence by its rank. The group is generated by formal differences of projections in matrix algebras over $\mathbb{R}$}, and so the dimension (rank) function gives the isomorphism.
  2. (2)
    We have
[…]
    The topological group $O_{n}$ has two path components, namely the rotations $SO_{n}$ and the reflections, and the inclusions $O_{n}\subset O_{n+1}$ respect this, so
[…]
  3. (3)
    For subsequent groups we need only look at the path component of the identity of $GL_{\infty}({\mathbb{R}})$ and hence really only at the special orthogonal group $SO_{n}$. So
    $KO_{2}({\mathbb{R}})\cong\pi_{1}(SO_{n})\qquad n\,\,large$
    Now $SO_{2}\cong S^{1}$ and so $\pi_{1}(SO_{2})\cong{\mathbb{Z}}$, but $SO_{3}\cong RP^{3}$ as a topological space (where $RP^{3}$ is real projective space) and hence $\pi_{1}(SO_{3})={\mathbb{Z}}_{2}$. For $n>2$, any map $S^{1}\to SO_{n}$ is homotopic373737by the Cellular Approximation Theorem, cf. G. Whitehead [86] p. 77 and see §9 ahead. to a map taking values in $SO_{3}$ and hence $\pi_{1}(SO_{n})={\mathbb{Z}}_{2}$ for $n>2$, confirming that
    $KO_{2}({\mathbb{R}})\cong{\mathbb{Z}}_{2}.$
  4. (4)
    To calculate
    $KO_{3}({\mathbb{R}})\cong\pi_{2}(SO_{n})\qquad n\,\,large$
    we may use again the fact that $SO_{3}\cong RP^{3}$ to deduce that $\pi_{2}(SO_{3})=0$. Steenrod shows (p. 116) that $\pi_{2}(SO_{4})=0$. Cellular Approximation tells us that $\pi_{2}(SO_{n})=0$ for $n>4$ and hence
    $KO_{3}({\mathbb{R}})\cong 0.$
    There is an alternate approach which is much simpler: a 1936 result of É. Cartan tells us that if $G$ is any compact Lie group then $\pi_{2}(G)=0.$
  5. (5)
    Steenrod calculates $\pi_{j}(SO_{n})$ for $j=3,4$ in §§22-24, and these imply that
    $KO_{4}({\mathbb{R}})\cong{\mathbb{Z}}\qquad and\qquad KO_{5}({\mathbb{R}})\cong 0.$
  6. (6)
    The story of the computation of the group
    $KO_{6}({\mathbb{R}})\cong\pi_{5}(SO_{n})=0$
    is very interesting historically. Its computation depended upon calculating $\pi_{5}(S^{3})$ and that group was the subject of controversy. Pontrjagin had announced that $\pi_{5}(S^{3})=0$, which conflicted with predictions stemming from Bott’s work on the Periodicity Theorem - the second author heard Bott mention once that he withheld publication of his results on periodicity for that reason. Eventually Pontrjagin and (independently) G. Whitehead determined that $\pi_{n+2}(S^{n})\cong{\mathbb{Z}}_{2}$. This all went on during the publication of Steenrod’s book [80] (1951) - you see mention of it in §24.11 and the end of §21.
7. External Products, the Künneth Theorem, and Internal Products
In this section we show that both real and complex $K$-theory have external products and that they fit together coherently. We also state simplified versions of the Künneth Theorem in both the real and complex settings. Jumping ahead a bit, we include in the presentation of the Künneth Theorem the restriction of the result to the (real and complex) $K$-theory for compact spaces.
External Products
Definition 7.1.
An homology theory $h_{*}$ on real or complex $C^{*}$-algebras is said to have external products if for any (perhaps with nuclearity assumed) algebras $A$ and $B$ in the category there is a family of natural transformations
$h_{m}(A)\otimes h_{n}(B)\longrightarrow h_{m+n}(A\otimes B)$
which are associative and graded commutative: if $x\in h_{m}(A)$ and $y\in h_{n}(B)$ then
$yx=(-1)^{mn}xy.$
Proposition 7.2.
  1. (1)
    For real $C^{*}$-algebras $A$ and $B$ (and nuclearity assumed), there exists a natural external product map
    $\mu_{O}\colon KO_{m}(A)\otimes KO_{n}(B)\rightarrow KO_{m+n}(A\otimes B).$
  2. (2)
    Similarly, for complex $C^{*}$-algebras $A$ and $B$ (and nuclearity assumed), there exist a natural external product map
    $\mu_{U}\colon K_{m}(A)\otimes K_{n}(B)\rightarrow K_{m+n}(A\otimes B).$
  3. (3)
    These product maps respect complexification. That is, the diagram
[…]
    commutes, where the vertical maps are complexification.
Proof.
(Sketch) If $p,q$ are projections in $M_{j}(A)$ and $M_{k}(B)$, respectively, then the Kronecker product [^38] If $M$ and $N$ are matrices representing linear transformations $V_{1}\to W_{1}$ and $V_{2}\to W_{2}$ with respect to fixed choices of bases $\{v_{i}^{1}\}$, $\{v_{j}^{2}\}$, $\{w_{k}^{1}\}$, $\{w_{l}^{2}\}$, then their Kronecker product $M\otimes N$ is the matrix representing the linear transformation $V_{1}\otimes V_{2}\to W_{1}\otimes W_{2}$ with respect to bases $\{v_{i}^{1}\otimes v_{j}^{2}\}$ and $\{w_{k}^{1}\otimes w_{l}^{2}\}$. gives $p\otimes q$, a projection in $M_{j+k}(A\otimes B)$. This induces a map
$KO_{0}(A)\otimes KO_{0}(B)\rightarrow KO_{0}(A\otimes B)$
(skipping over details regarding how this is defined vis a vis the Grothendieck construction). Taking suspensions, we then obtain maps
$KO_{0}(S^{m}A)\otimes KO_{0}(S^{n}B)\rightarrow KO_{0}(S^{m+n}(A\otimes B))$
for all $m,n\geq 0$, which gives $\mu_{O}$.
The analogous construction gives the complex pairing.
The Künneth Theorem
Künneth theorems deal in general with determining $h_{*}(A\otimes B)$ in terms of $h_{*}(A)$ and $h_{*}(B)$. One has to put some constraints on the algebras in order for the theorem to hold. The full Künneth Theorem for $K$-theory involves some serious homological algebra. For $K_{*}(A)$ it was established by Schochet [73] (1984) where the bootstrap category was first introduced, and for $KO_{*}(A)$ it is due to Boersema [15] (2002), utilizing united $K$-theory. The following results are special cases of these general theorems. Anticipating events, we also state the corresponding results for the $K$-theory of topological spaces. See Section 11 for $K^{*}(X)$.
Theorem 7.3.
The Künneth Theorem (special cases)
  1. (1)
    Suppose that $A$ and $B$ are complex $C^{*}$-algebras, $A$ is in the bootstrap category (which includes all commutative complex $C^{*}$-algebras) and $K_{*}(B)$ is torsion-free. Then the natural external product map induces an isomorphism of graded groups [^39] If $G_{*}$ and $H_{*}$ are graded abelian groups then $(G_{*}\otimes H_{*})_{5}$ (for example) is understood to be the direct sum of the groups $G_{j}\otimes H_{5-j}$. This makes sense as long as only a finite number of these groups are non-zero. Bott periodicity makes things more complicated. So let us stipulate that whenever we are tensoring $K$-theory groups, that $\in\{0,1\}$ in the complex case and $\in\{0,1,\dots 6,7\}$ in the real case.
    $K_{*}(A)\otimes K_{*}(B)\overset{\cong}{\longrightarrow}K_{*}(A\otimes B).$
    In particular, if $X$ and $Y$ are compact spaces and $K^{*}(Y)$ is torsion-free then the natural external product map induces an isomorphism of graded groups
    $K^{*}(X)\otimes K^{*}(Y)\overset{\cong}{\longrightarrow}K^{*}(X\times Y).$
  2. (2)
    Suppose that $A$ and $B$ are real $C^{*}$-algebras, $A_{\mathbb{C}}$ is in the bootstrap category (this includes all commutative real $C^{*}$-algebras $A$), and $B$ has the property that $K_{*}(B_{\mathbb{C}})$ is torsion-free. Then the natural external product map induces an isomorphism of graded $KO_{*}({\mathbb{R}})$-modules [^40] We are casually assuming that $KO_{*}({\mathbb{R}})$ is a graded commutative ring that acts on $KO_{*}(A)$ for any real $C^{*}$-algebra $A$. We will establish this rigorously subsequently. Actually the analogous statement in the complex case is also true, but since $K_{*}({\mathbb{C}}))$ is just $\mathbb{Z}$} and 0 in degrees 0 and $1$ respectively, the action has little significance here.
    $KO_{*}(A)\otimes_{KO_{*}({\mathbb{R}})}KO_{*}(B)\overset{\cong}{\longrightarrow}KO_{*}(A\otimes B).$
    This includes cases where $B=C^{{\mathbb{R}}}(iS^{1})$, $C_{o}^{{\mathbb{R}}}(i{\bf{R}})$, $\real(S^{n})$. In particular, if $X$ and $Y$ are compact spaces and $Y$ has the homotopy type of a sphere or a product of spheres414141or a finite or countable disjoint union of such spaces, then there is a natural isomorphism
    $KO^{*}(X)\otimes_{KO^{*}(pt)}KO^{*}(Y)\overset{\cong}{\longrightarrow}KO^{*}(X\times Y).$
Internal Products
Internal products present a major challenge, since in general they may not exist.
Definition 7.4.
An homology theory $h_{*}$ on a category $\mathcal{A}$ of real or complex $C^{*}$-algebras is said to have internal products if for any (perhaps with nuclearity assumed) algebras $A$ in the category there is a family of natural transformations
$h_{m}(A)\otimes h_{n}(A)\longrightarrow h_{m+n}(A)$
which is associative and graded commutative: if $x\in h_{m}(A)$ and $y\in h_{n}(A)$ then
$yx=(-1)^{mn}xy.$
In other words, $h_{*}$ is a functor from $\mathcal{A}$ to the category of graded commutative rings.
Proposition 7.5.
  1. (1)
    Restrict $K_{*}$ to the category of commutative complex $C^{*}$-algebras. Then $K_{*}$ has an internal product.
  2. (2)
    Restrict $KO_{*}$ to the category of commutative real $C^{*}$-algebras. Then $KO_{*}$ has an internal product.
Proof.
  1. (1)
    The natural map $m:A\otimes A\to A$ given by $x\otimes y\to xy$ is a map of $C^{*}$-algebras since $A$ is commutative, and hence it induces a map
    $K_{m+n}(A\otimes A)\longrightarrow K_{m+n}(A)$
    which is all we need to convert the standard external product into an internal product. To see this more geometrically, we can use the Gelfand transform: write $A=C^{{\mathbb{C}}}(X)$ for some compact space $X$ and then the diagonal map $X\to X\times X$ corresponds to $m$ under the transform.
  2. (2)
    The proof in the real case is identical to the complex case if we stick to algebras. The geometric argument is also available, based on the diagonal map
    $(X,\tau)\longrightarrow(X,\tau)\times(X,\tau).$
Proposition 7.6.
  1. (1)
    Complex $K$-theory $K_{*}$ is a functor from the category of commutative complex $C^{*}$-algebras to the category of graded commutative rings. That is, for any commutative complex $C^{*}$-algebra $A$, there is a natural graded commutative ring structure on $K_{*}(A)$, and if $A\to B$ is a map of commutative complex $C^{*}$-algebras then the induced map on $K_{*}$ is a map of graded commutative rings.
  2. (2)
    Real $K$-theory $KO_{*}$ is a functor from the category of commutative real $C^{*}$-algebras to the category of graded commutative rings. That is, for any commutative real $C^{*}$-algebra $A$, there is a natural graded commutative ring structure on $KO_{*}(A)$, and if $A\to B$ is a map of commutative real $C^{*}$-algebras then the induced map on $KO_{*}$ is a map of graded commutative rings.
We have noted previously that $KO^{*}(X)$ has the structure of a graded ring for all compact spaces $X$, and this includes the case $X=pt$ a point. It is very important to specify the grading here, and so we say explicitly that
$KO^{-n}(X)\cong KO_{n}(\real(X))$
and, in particular,
$KO^{-n}(pt)\cong KO_{n}({\mathbb{R}}).$
Corollary 7.7.
$KO_{*}({\mathbb{R}})$ is a graded commutative ring. For any real $C^{*}$-algebra $A$, $KO_{*}(A)$ is a graded module over $KO_{*}({\mathbb{R}})$. Similarly, $K_{*}({\mathbb{C}})$ is a graded ring and for any complex $C^{*}$-algebra $A$, $K_{*}(A)$ is a graded module over $K_{*}({\mathbb{C}})$.
Proof.
Since $\mathbb{R}$$\otimes{\mathbb{R}}\cong{\mathbb{R}}$, there are maps $KO_{m}({\mathbb{R}})\otimes KO_{n}({\mathbb{R}})\rightarrow KO_{m+n}({\mathbb{R}})$ which provide the multiplicative structure.
Furthermore for any real $C^{*}$-algebra $A$, we have ${\mathbb{R}}\otimes A=A$ which give maps $KO_{m}({\mathbb{R}})\otimes KO_{n}(A)\rightarrow KO_{m+n}(A)$. This is the $KO_{*}({\mathbb{R}})$-module structure on $KO_{*}(A)$.
The complex case is essentially identical. ∎
Remark 7.8.
We shall describe the ring structure of $KO_{*}({\mathbb{R}})$ and $KO_{*}({\mathbb{R}})\otimes{\mathbb{Q}}$ in some detail in Section 15. Note that
$KO_{n}({\mathbb{R}})\otimes{\mathbb{Q}}=\begin{cases}{\mathbb{Q}}&n=0,\,4\\ 0&\text{otherwise}\end{cases}$
It turns out that $KO_{n}(A)\otimes{\mathbb{Q}}$ is periodic of period $4$.
Next we introduce one of the other homology theories that may be constructed from $KO_{*}$.
Definition 7.9.
(Karoubi) Quaternionic $K$-theory is defined by
$KH_{*}(A)=KO_{*}(A\otimes{\mathbb{H}})$
Remark 7.10.
(Karoubi) We know from Cor. 5.9 that $KH_{*}$ is indeed an additive homology theory. There is a natural exterior pairing
$KH_{r}(A)\otimes KH_{s}(A)=KO_{r}(A\otimes{\mathbb{H}})\otimes KO_{s}(A\otimes{\mathbb{H}})\longrightarrow KO_{r+s}(A\otimes A\otimes{\mathbb{H}}\otimes{\mathbb{H}}).$
Since $H\otimes H\cong M_{r}({\mathbb{R}})$ as remarked in §1, this simplifies to
$KH_{r}(A)\otimes KH_{s}(A)\longrightarrow KO_{r+s}(A\otimes A)$
and if $A$ is real commutative we obtain a pairing
$KH_{r}(A)\otimes KH_{s}(A)\longrightarrow KO_{r+s}(A)$
which, as we shall see, is related to real Bott periodicity.
For the relationship between $KH_{*}$, $KSp_{*}$ and other related theories, please see [54], particularly Example 2.16.
8. Bott Periodicity
In this section we show how Bott Periodicity provides real power to $KO_{*}$ and $K_{*}$. Complex $K$-theory is periodic of period 2. In contrast,
Theorem 8.1.
(Bott Periodicity) [21], [45], [89]. For any real $C^{*}$-algebra $A$, there is a natural isomorphism
$KO_{j+8}(A)\cong KO_{j}(A).$
This theorem was initially established by Bott [21] in 1957 in the context of the homotopy groups of the classical Lie groups, i.e., before there really was a $KO$-theory. There are many, many other proofs of Bott’s theorem, originating in several different areas of math, illustrating its importance.
In the late 1960’s, Wood [89] and Karoubi [45] (see also Milnor [62]) proved it in the context of real (and also complex) Banach algebras and Banach categories. Here is a vague outline of Wood’s proof. (We shall return to this in much greater detail in Part 3.) Recall from Chapter 6 that
$KO_{n}(A)\cong\pi_{n-1}(GL_{\infty}(A)).$
Wood proves that the connected component of the identity of $GL_{\infty}(A)$ is homotopy equivalent to the connected component of the identity of the 8th loop space $\Omega^{8}GL_{\infty}(A)$ of $GL_{\infty}(A)$. [^42] See the next section for the nice properties of loop spaces $\Omega^{k}X$. This implies that, for each $n$, that there is an isomorphism
$\pi_{j}(GL_{\infty}(A))\cong\pi_{j}(\Omega^{8}GL_{\infty}(A))$
But
$\pi_{j}(\Omega^{8}GL_{\infty}(A))\cong\pi_{j+8}(GL_{\infty}(A))$
which translates to
$KO_{j+8}(A)\cong KO_{j}(A).$
Note that just as $GL_{n}({\mathbb{R}})$ deformation retracts down to $O_{n}({\mathbb{R}})$, the group $GL_{n}(A)$ deformation retracts down to $O_{n}(A)$ and so we may use orthogonal groups rather than the larger general linear groups.
Corollary 8.2.
Using periodicity, we may extend $KO_{*}(A)$ for all $\in{\mathbb{Z}}.$
[…]
Corollary 8.3.
For each $n>0$,
$KO_{n}({\mathbb{R}})\cong\pi_{n-1}(\mathbf{O})\cong\pi_{0}(\Omega^{n-1}\mathbf{O}).$
This is a useful way to think about $KO_{n}({\mathbb{R}})$, since each of the spaces $\Omega^{k}\mathbf{O}$ may be constructed rather concretely from classical Lie groups, as we shall see.
We now recap the principal properties of $KO_{*}$ for future use. Let $\mathscr{C}^{*}_{real}$ denote the category of real $C^{*}$-algebras and their homomorphisms.
Theorem 8.4.
Real $K$-theory $KO_{*}(A)$ is a $\mathbb{Z}$}-graded additive homology theory defined on the category $\mathscr{C}^{*}_{real}$ of all real $C^{*}$-algebras and taking values in $\mathbb{Z}$}-graded abelian groups. That is, it satisfies the following axioms:
  1.
  2. (1)
    Homotopy Axiom. Let $H:A\to IB$ be a homotopy from $f=p_{0}H$ to $g=p_{1}H$ in $\mathscr{C}^{*}_{real}$. Then
    $f_{*}=g_{*}:KO_{n}(A)\longrightarrow KO_{n}(B)\quad\text{for all n}.$
  3. (2)
    Exactness Axiom. Let
    $0\longrightarrow J\overset{i}{\longrightarrow}A\overset{j}{\longrightarrow}A/J\longrightarrow 0$
    be a short exact sequence in $\mathscr{C}^{*}_{real}$. Then for each $n$ there is a map
    $\partial:KO_{n}(A/J)\longrightarrow KO_{n-1}(J)$
    and a long434343very long! It extends infinitely in both directions. exact sequence
    $\dots\to KO_{n}(J)\overset{i_{*}}{\longrightarrow}KO_{n}(A)\overset{j_{*}}{\longrightarrow}KO_{n}(A/J)\overset{\partial}{\longrightarrow}KO_{n-1}(J)\to\dots$
    The map $\partial$ is natural with respect to morphisms of short exact sequences.
  4. (3)
    Additivity Axiom. Let $A_{i}\in\mathscr{C}^{*}_{real}$ and let
    $A=\oplus_{i=1}^{\infty}A_{i}.$
    Then the natural maps $A_{j}\to A$ induce an isomorphism
    $\oplus_{i}KO_{n}(A_{i})\overset{\cong}{\longrightarrow}KO_{n}(A)$
    for each $n$.
    In addition, $KO_{*}$ satisfies the following properties:
  5. (4)
    Matrix Stability: For all real $C^{*}$-algebras $A$, positive integers $k$ ,and integers $n$,
    $KO_{n}(A)\cong KO_{n}(M_{k}(A))\cong KO_{n}(A\otimes{\mathcal{K}}_{\mathbb{R}}).$
  6. (5)
    Periodicity: For any real $C^{*}$-algebra $A$ and $j\in{\mathbb{Z}}$, there is a natural Bott Periodicity isomorphism
    $KO_{j+8}(A)\cong KO_{j}(A).$
    Stated in homotopy terms, Periodicity asserts that $\Omega^{8}\mathbf{O}\simeq\mathbf{O}$.
Here are the analogous properties for complex $C^{*}$-algebras and their $K$-theory. Let $\mathscr{C}^{*}_{cx}$ denote the category of complex $C^{*}$-algebras and their homomorphisms.
Theorem 8.5.
Complex $K$-theory $K_{*}(A)$ is a $\mathbb{Z}$}-graded additive homology theory defined on the category $\mathscr{C}^{*}_{cx}$ of all complex $C^{*}$-algebras and taking values in $\mathbb{Z}$}-graded abelian groups. That is, it satisfies the following axioms:
  1.
  2. (1)
    Homotopy Axiom. Let $H:A\to IB$ be a homotopy from $f=p_{0}H$ to $g=p_{1}H$ in $\mathscr{C}^{*}_{cx}$. Then
    $f_{*}=g_{*}:K_{n}(A)\longrightarrow K_{n}(B)\quad\text{for all n}.$
  3. (2)
    Exactness Axiom. Let
    $0\longrightarrow J\overset{i}{\longrightarrow}A\overset{j}{\longrightarrow}A/J\longrightarrow 0$
    be a short exact sequence in $\mathscr{C}^{*}_{cx}$. Then for each $n$ there is a map
    $\partial:K_{n}(A/J)\longrightarrow K_{n-1}(J)$
    and a long exact sequence
    $\dots\to K_{n}(J)\overset{i_{*}}{\longrightarrow}K_{n}(A)\overset{j_{*}}{\longrightarrow}K_{n}(A/J)\overset{\partial}{\longrightarrow}K_{n-1}(J)\to\dots$
    The map $\partial$ is natural with respect to morphisms of short exact sequences.
  4. (3)
    Additivity Axiom. Let $A=\oplus_{i=1}^{\infty}A_{i}$ in $\mathscr{C}^{*}_{cx}$. Then the natural maps $A_{j}\to A$ induce an isomorphism
    $\oplus_{i}K_{n}(A_{i})\overset{\cong}{\longrightarrow}K_{n}(A)$
    for each $n$.
    In addition, $K_{*}$ satisfies the following properties:
  5. (4)
    Matrix Stability: For all complex $C^{*}$-algebras $A$ and positive integer $k$ and integer $j$,
    $K_{j}(A)\cong K_{j}(M_{k}(A))\cong K_{j}(A\otimes{\mathcal{K}}_{\mathbb{C}})$
  6. (5)
    Periodicity: For any complex $C^{*}$-algebra $A$, there is a natural Bott Periodicity isomorphism
    $K_{j+2}(A)\cong K_{j}A).$
    Stated in homotopy terms, Periodicity asserts that $\Omega^{2}\mathbf{U}\simeq\mathbf{U}$.
Remark 8.6.
(M. Karoubi)
We know from Cor. 5.9 that $KH_{*}$ is indeed a homology theory. It is periodic of period 4; the isomorphism is given by tensoring with the class of the identity in $KO_{0}({\mathbb{H}})$. In Remark 7.10 we noted that if $A$ is a real commutative $C^{*}$-algebra then there is a natural map
$KH_{r}(A)\otimes KH_{s}(A)\longrightarrow KO_{r+s}(A).$
Take $A={\mathbb{R}}$ and $r=s=4.$ Then we have
$KH_{4}(A)\otimes KH_{4}(A)\longrightarrow KO_{8}(A).$
Now
$KH_{4}({\mathbb{R}})\cong\pi_{3}(GL_{\infty}({\mathbb{H}}))$
by the first definition of $KO_{*}$. By CW-approximation,
$\pi_{3}(GL_{\infty}({\mathbb{H}})\cong\pi_{3}(GL_{1}({\mathbb{H}})\cong\pi_{3}(S^{3})\cong{\mathbb{Z}}$
The resulting map
${\mathbb{Z}}\cong KH_{4}(A)\otimes KH_{4}(A)\longrightarrow KO_{8}(A)\cong{\mathbb{Z}}.$
sends $1\otimes 1$ to the Bott generator of $KO_{8}({\mathbb{R}})$.
For further information on this matter, and in particular the relationship of $KH_{*}$ to symplectic $K$-theory, please see Karoubi and Weibel [53].
Part II $KO^{*}$-theory for Spaces
9. The Basics: Topological Spaces
In this section we review the information we need about topological spaces, loop spaces and classifying spaces of $\mathbf{O}$ and $\mathbf{U}$, in order to present $KO$-theory for spaces.
Definition 9.1.
A based topological space is a topological space $X$ together with a choice of basepoint $x_{o}\in X$. A based map $f:X\to Y$ of based topological spaces is a continuous map from $X$ to $Y$ with the property that $f(x_{0})=y_{0}$. We denote the set of based maps from $X$ to $Y$ by $F(X,Y)$ and based homotopy classes of based maps by $[X,Y]$. In particular, we may define the k’th homotopy group of $X$ by
$\pi_{k}(X)=[S^{k},X].$
Note that on occasion we will want to deal with spaces and maps without specified basepoints, in which case we will denote the set of maps by $F(X,Y)^{u}$, where $u$ refers to “unbased.”  [^44] We considered using the term “baseless” but that seemed too derogatory. Following Bourbaki, we understand compact spaces to be compact and Hausdorff.
In order to proceed, and as noted in the Introduction, we must put a small point-set topological restriction on our topological spaces: we follow G. Whitehead [86] and May [59].
Definition 9.2.
A space $X$ is compactly generated if $X$ is Hausdorff and if it has the following property: if $Y$ is a subset of $X$ and $Y\cap C$ is closed in $X$ for each compact subset $C$ of $X$, then $Y$ itself is closed in $X$.
Every locally compact space is compactly generated. So is every metrizable space, so this includes smooth manifolds and in particular all of the topological groups that we will need as well as the homogeneous spaces such as the space $O_{2k}/(O_{k}\times O_{k})$. All CW complexes, including spaces like $\mathbf{O}$ are compactly generated. If you confront a space $X$ that is not compactly generated then it is easy to retopologize it canonically to get a new compactly generated space $kX$ and a map $kX\to X$ which is the identity function and is continuous. [^45] See May [59] p. 37) for details.
We assume henceforth that all spaces are compactly generated.
Next we need some basic homotopy constructions. Some of the constructions need to be retopologized (as just described) to make them stay in the category of compactly generated spaces, and we will assume that that automatically happens when needed. For example, if we take the function space $F(X,Y)$ of all based maps from $X$ to $Y$ with the compact-open topology then that requires a slight adjustment to make it compactly generated. The basepoint of $F(X,Y)$ will be the constant function $f(x)=y_{0}$ for all $x\in X$.
Definition 9.3.
Suppose that $X$ and $Y$ are based compactly generated spaces. Then:
  1.
  2. (1)
    The product $X\times Y$ is given by
    $X\times Y=\{(x,y):x\in X,y\in Y\}$
    with basepoint $(x_{0},y_{0})$ and the product topology.
  3. (2)
    The join (or one-point union) of $X$ and $Y$, denoted $X\vee Y$, is the disjoint union of $X$ and $Y$ with $x_{0}$ identified with $y_{0}$. Alternatively, think of this as a subset of $X\times Y$, namely
    $X\vee Y=\{(x,y)\in X\times Y:x=x_{0}\,\,or\,\,y=y_{0}\,\,(or\,\,both)\}$
    topologized as a closed subset of $X\times Y$. For example, the join of two circles is a figure 8 and the point that the two circles have in common is the basepoint.
  4. (3)
    The smash product $X\wedge Y$ is given by
    $X\wedge Y=(X\times Y)/(X\vee Y),$
    the product $X\times Y$ modulo the collapsing of the join $X\vee Y$ to a point, which becomes the basepoint. This is topologized as a quotient of $X\times Y$. For example, $S^{1}\times S^{1}$ is the 2-torus, whereas $S^{1}\wedge S^{1}$ is the 2-sphere $S^{2}$.
  5. (4)
    The suspension of a topological space $X$ is the space
    $SX=S^{1}\wedge X.$
    We let
    $S^{n}X=S^{n}\wedge X$
    and note that
    $S(S^{n}X)\cong S^{n+1}X.$
    Sometimes in the literature this is called the reduced suspension of $X$, with the term “suspension” reserved for $S^{1}\times X$.
  6. (5)
    The loop space of a topological space $X$ is defined to be
    $\Omega X=F(S^{1},X)$
    and higher loop spaces are given by
    $\Omega^{n}X\cong F(S^{n},X).$
    Note that
    $\Omega(\Omega^{n}X)\cong\Omega^{n+1}X.$
  7. (6)
    The pullback construction is defined exactly as for $C^{*}$-algebras. That is, given a diagram
[…]
    then the pullback $P$ is defined to be
    $P=\{(x,y)\in X\times Y:f(x)=g(y)\}$
    with the subspace topology and with associated commuting diagram
[…]
  8. (7)
    The pushout construction is dual. That is, given a diagram
[…]
    then the pushout $Q$ is defined by taking the disjoint union of $X$ and $Y$ and identifying $f(w)$ with $g(w)$ for each $w\in W$.
[…]
    There is a very important special case . Suppose that $Y=D^{n}$, the $n$-dimensional disk with boundary sphere $S^{n-1}$, $\iota$ is the inclusion map
    $\iota:S^{n-1}\to D^{n}$
    and we are given a map $f:S^{n-1}\to X$. Then we have a diagram
[…]
    and the associated pushout is denoted $X\cup_{f}D^{n}$. This construction is called adding an $n$-cell to $X$. Here’s a specific example: if $f:S^{3}\to S^{2}$ is the Hopf map, then
    $S^{2}\cup_{f}D^{4}\cong{\mathbb{C}}P^{2}$
    the projective plane.
Both $S$ and $\Omega$ are covariant functors. If $f:X\to Y$ is a based map then there are natural induced maps $Sf:SX\to SY$ and $\Omega f:\Omega X\to\Omega Y$ with the requisite properties. Better yet, we have the following classical result, which is basic in homotopy theory and of critical importance in understanding the physics invariants to come.
Theorem 9.4.
Suppose that $X$ and $Y$ are based and compactly generated. Then the functors $S$ and $\Omega$ are adjoint functors. That is, there is a natural homeomorphism
$F(SX,Y)\overset{\cong}{\longrightarrow}F(X,\Omega Y).$
and, by iterating,
$F(S^{n}X,Y)\overset{\cong}{\longrightarrow}F(X,\Omega^{n}Y).$
Proof.
We will prove something stronger, namely that if $W$ is also based and compactly generated, then there is a natural homeomorphism
$F(W\wedge X,Y)\overset{\cong}{\longrightarrow}F(X,F(W,Y)).$
Suppose that $f:W\wedge X\to Y$. Define $\phi(f):X\to F(W,Y)$ by
$\phi(f)(x)(w)=f(w\wedge x).$
One checks that this map is well-defined, continuous, preserves basepoints, and then write down its inverse as follows. If $g:X\to F(W,Y)$ then define $\psi(g):W\wedge X\to Y$ by
$\psi(g)(w\wedge x)=g(x)(w).$
Then check that $\psi$ is well-defined, continuous, preserves basepoints, and that $\Psi$ is the inverse for $\phi$. ∎
Next, we must briefly discuss CW-complexes. This is a big classical subject and is very nicely developed in many texts, such as G.W. Whitehead [86] Ch. 2, and J. P. May [59] Ch. 10. We will give a small taste with no details, which will suffice for our needs.
We have explained what is meant by adding an $n$-cell above. There are easy examples. If $X$ is a point then $X\cup_{f}D^{n}\cong S^{n}$. If $X=S^{2}$ and $f:S^{3}\to S^{2}$ is the Hopf map then $S^{2}\cup_{f}D^{4}\cong CP^{2}$.
Definition 9.5.
A CW-complex $X$ is a topological space $X$ which is the union of spaces $X^{0}\subseteq X^{1}\subseteq X^{2}\subseteq X^{3}\subseteq\dots$ where $X^{0}$ is a discrete set of points, called vertices, and inductively, $X^{n}$ (called the n-skeleton of $X$) is obtained from $X^{n-1}$ by adding on some [^46] Perhaps none and perhaps an infinite number n-cells. A map $f:X\to Y$ of CW-complexes is said to be cellular if $f(X^{n})\subseteq Y^{n}$ for each n. If only a finite number of cells are added in total then $X$ is called a finite CW-complex and, of course, $X^{k}=X$ for $k$ sufficiently large.
It is easy to give examples of CW-complexes, though the structure may not be obvious at first. For instance, any compact connected manifold $M$ is homotopy equivalent to a finite CW-complex with one vertex. Thus spaces like $U_{n}$ and $SO_{n}$ fit this description, as does $O_{n}$ if we allow two vertices. On a much deeper level, we may, for example, regard the sequence
$O_{1}\subset O_{2}\subset O_{3}\subset\dots\subset O_{n}\subset\dots\mathbf{O}$
as a sequence of skeleta with $X^{n}=O_{n+1}$, showing that $\mathbf{O}$ is a CW-complex. Similarly, the sequence
$RP^{1}\subset RP^{2}\subset RP^{3}\subset\dots\subset RP^{n}$
tells us how real projective spaces are constructed.
How does all of this help us? Here’s a nice consequence.
Theorem 9.6.
Suppose that $X$ is a CW-complex with skeleta $X^{n}$ and associated maps $\rho_{n}:X^{n}\to X$. Then
  1. (1)
    The map
    $(\rho_{n})_{*}:\pi_{k}(X^{n})\longrightarrow\pi_{k}(X)$
    is an isomorphism for $k<n$ and is onto for $k=n$.
  2. (2)
    The natural induced map $\rho$
[…]
    is an isomorphism.
For example, it is not hard to show that $\pi_{1}(RP^{2})={\mathbb{Z}}_{2}.$ and hence $\pi_{1}(RP^{n})={\mathbb{Z}}_{2}$ for all $n\geq 2$ and $\pi_{1}(RP^{\infty})={\mathbb{Z}}_{2}$
Proof.
This is an easy consequence of the Cellular Approximation Theorem (cf. G. W. Whitehead [86] II. 4.5) which says, in part, that if $f:Y\to X$ is a map of CW-complexes, then $f$ is homotopic to a cellular map $f^{\prime}$, a map that has the property that $f^{\prime}(Y^{j})\subseteq X^{j}$ for each $j$. Taking $Y=S^{j}$, a sphere, this tells us that a map $f:S^{j}\to X$ is homotopic to a map $S^{j}\to X^{j}$. So, for example, if $n>2$ then any map $S^{2}\to S^{n}$ is null-homotopic. ∎
10. The Basics: Fibrations, Fibre Bundles, Principal Bundles, and $\mathbf{BO}$
Fibrations and fibre bundles underlie large portions of mathematics and physics, albeit sometimes in disguise. For example, real and complex vector bundles are going to be prominent for us later in this work. For now we sketch out the basics.
There are several fine sources for this material, and we would particularly mention Steenrod [80], G. W. Whitehead [86], Husemoller [44], and, more recently, J. P. May [59]. We present here merely a brief outline based upon May [59] Chapter 7, and urge the interested reader to explore further in those sources.
Definition 10.1.
A surjection $p:E\to B$ is a (Hurewicz) fibration if it satisfies the covering homotopy property (CHP). Given a commuting diagram
[…]
with $\iota(y)=(y,0)$, then there exists a map $\tilde{h}$ making the diagram commute. In other words, if one end of the homotopy $h$ may be covered (by $f$) to $E$ then the whole homotopy may be covered by $\tilde{h}$.
For each $b\in B$, the space $F_{b}=p^{-1}(b)$ is called the fibre over the point $b$.
We shall see many examples in future sections; we mention only a few here.
  1.
  2. (1)
    The trivial fibration
    If $E=B\times F$ with $p=\pi_{1}:B\times F\to B$ the projection on the first factor, then $E\overset{p}{\longrightarrow}B$ is a fibration, and then $F_{b}\cong F$ for each $b\in B$.
  3. (2)
    The Möbius band
    Form the rectangle $0,17]\times[0,1]$. If we then identify $(0,t)$ with $(17,t)$ we obtain a cylinder of width $1$ and circumference $17$ and this is a trivial fibration over the circle of circumference $17$. However, if we identify $(0,t)$ with (17,1t)(17,1-t) then we obtain a non-trivial fibration over the circle, with total space the Möbius band.
  4. (3)
    Coset spaces
    If $G$ is a Lie group and $H$ a closed subgroup with coset space $G/H$ then the sequence
    $H\longrightarrow G\longrightarrow G/H$
    is a fibration with fibre $H$.
Proposition 10.2.
If $p:E\to B$ is a fibration and $f:B^{\prime}\to B$ is some map then the pullback $p^{\prime}:p^{-1}(E)\to B^{\prime}$ in the associated pullback diagram
[…]
is a fibration.
Definition 10.3.
An open cover $\mathcal{U}=\{U_{\alpha}\}$ of a space $B$ is said to be numerable if for each open set $U$ in the cover there is a continuous map $\lambda_{U}:B\to[0,1]$ such that
$\lambda_{U}^{-1}(0,1]=U$
and the cover is locally finite; that is, each $b\in B$ appears in only finitely many open sets of the cover.
Any open cover of a paracompact [^47] A space is paracompact if every open cover of the space has a locally finite refinement. Examples include compact spaces and metrizable spaces. space has a numerable refinement.
Theorem 10.4.
(cf. [59], p. 49.) Suppose that $\mathcal{U}$ is a numerable open cover of $B$. Then the following are equivalent:
  1. (1)
    The map $p:E\longrightarrow B$ is a fibration.
  2. (2)
    For every open set $U\in\mathcal{U}$, the pullback
[…]
    is a fibration.
Theorem 10.5.
(cf. [59] p. 52) If the base space $B$ of a fibration $p:E\to B$ is path-connected then any two fibres $F_{x}$ and $F_{y}$ are homotopy equivalent.
Next, we record what will be for us one of the most useful properties of fibrations.
Theorem 10.6.
(cf. May [59] p. 64). Suppose that $p:E\to B$ is a fibration, $B$ is path-connected with basepoint $b_{0}$. Let $F=p^{-1}(b_{0})$ and fix some basepoint in $F$. Then there is a natural long exact sequence of homotopy groups
$\to\pi_{n}(F)\to\pi_{n}(E)\overset{p_{*}}{\longrightarrow}\pi_{n}(B)\overset{\delta}{\longrightarrow}\pi_{n-1}(F)\to\dots\to\pi_{0}(B)\to 0.$
Proposition 10.7.
Suppose that
$F\longrightarrow E\longrightarrow B$
is a fibration of CW-complexes and $E$ is contractible. Then $F$ is homotopy equivalent to $\Omega B$.
Proof.
The long exact homotopy sequence for fibrations degenerates, since $\pi_{j}(E)=0$ for all $j$, and we are left with
$0\longrightarrow\pi_{j+1}(B)\overset{\cong}{\longrightarrow}\pi_{j}(F)\longrightarrow 0$
and so
$\pi_{j}(F)\cong\pi_{j+1}(B)\cong\pi_{j}(\Omega B)\qquad\forall j.$
If $F$ and $B$ are CW-complexes then so is $\Omega B)$, and this implies by Whitehead’s theorem that they are homotopy equivalent. ∎
Definition 10.8.
A map $p:E\to B$ is a fibre bundle [^48] Some people spell the word “fiber” but we follow Norman Steenrod, the second author’s mathematical ancestor who, quite literally, wrote the (first!) book on “Fibre Bundles” [80] in 1951.
if it possesses a numerable open cover $\mathcal{U}$ and there is a fixed topological space $F$ and homeomorphisms $\phi_{U}$ for each open set $U$ in the cover such that the diagram
[…]
commutes. Every fibre bundle is a fibration.
Perhaps the most well-known examples of fibre bundles are real and complex vector bundles, such as the tangent bundle of a smooth manifold. We will be discussing those in future sections. For now we concentrate upon a different class of examples, namely principal $G$-bundles. This will be very sketchy and we recommend Husemoller [44] starting on page 42 for a detailed treatment.
Definition 10.9.
Suppose that $G$ is a topological group and $p:E\to B$ is a fibre bundle together with a continuous right action $E\times G\to G$. Then $p:E\to B$ is a principal $G$-bundle and is homeomorphic to $E\to E/G$ if the following properties hold:
  1.
  2. (1)
    For each $b\in B$, the action of $G$ restricts to an action on the fibre
    $F_{b}\times G\to F_{b}.$
  3. (2)
    $G$ acts freely on $F_{b}$ ; that is, $xg=x$ implies that $g$ is the identity element of $G$.
  4. (3)
    $G$ acts transitively on $F_{b}$; that is, if $x,y\in F_{b}$ then there exists some $g\in G$ with $xg=y$.
  5. (4)
    For each $b\in B$ and $x\in F_{b}$, the map $\phi_{b,x}:G\to F_{b}$ given by
    $\phi_{b,x}(g)=xg$
    is a homeomorphism.
Hence $F_{b}\cong G$ for each $b\in B$ and the orbit space $E/G$ is homeomorphic to $B$. So, up to isomorphism, we have
$G\longrightarrow E\longrightarrow E/G.$
If $G$ is a Lie group, the action on $E$ is smooth, and $E\to B$ is a map of smooth manifolds then the construction may be carried out smoothly.
Examples of principal $G$ bundles that will occur many times in this work are easy to state. Take $G$ to be a compact Lie group, $H$ a closed subgroup. For example, we have the coset fibrations
$O_{n}\longrightarrow O_{n+1}\longrightarrow O_{n+1}/O_{n}\cong S^{n}$
and
$O_{n}\times O_{n}\longrightarrow O_{2n}\longrightarrow O_{2n}/(O_{n}\times O_{n}).$
These are both principal bundles.
Here’s another example. Take $G={\mathbb{Z}}_{2}$ and let ${\mathbb{Z}}_{2}$ act on the sphere $S^{n}$ by the antipodal action. Then there is a natural principal ${\mathbb{Z}}_{2}$ bundle, namely
${\mathbb{Z}}_{2}\longrightarrow S^{n}\longrightarrow RP^{n}$
with base space real projective space. There is an obvious map of principal ${\mathbb{Z}}_{2}$- bundles
[…]
and taking direct limits we obtain
${\mathbb{Z}}_{2}\longrightarrow S^{\infty}\longrightarrow RP^{\infty}$
The space $S^{\infty}$ is contractible, and so this presents us with an example of a universal $G$-bundle:
Definition 10.10.
(cf. May [59] 196-198) A principal $G$-bundle
$G\longrightarrow E\longrightarrow B$
is a universal principal $G$-bundle if the total space $E$ is contractible.
Any two universal principal $G$-bundles are homotopy equivalent. The usual notation for the universal principal $G$-bundle is
$G\longrightarrow EG\longrightarrow{{BG}}$
The space $BG$ is called the classifying space of $G$-bundles for reasons that will become apparent.
So, for example,
${\mathbb{Z}}_{2}\longrightarrow S^{\infty}\longrightarrow RP^{\infty}$
is the universal principal ${\mathbb{Z}}_{2}$-bundle and $RP^{\infty}\cong{{B{\mathbb{Z}}_{2}}}$.
Remark 10.11.
The universal principal $G$-bundles play a critical role in the classification of $G$-bundles. The basic fact is that the pullback of a principal $G$-bundle is itself a principal $G$-bundle. Then for some fixed space $B$ (e.g. paracompact to simplify matters) and some continuous function $f:B\to BG$, the pullback diagram
[…]
produces a principal $G$-bundle $E\to B$. This gives us a map
$B,BG]^{u}\longrightarrow Prin_{G}(B)$
from homotopy classes of maps $B\to BG$ to isomorphism classes of principal $G$-bundles over $B$. This map is an isomorphism (cf. Husemoller [44] 12.2 ). So this justifies calling $BG$ a ”classifying space” since homotopy classes of maps into it classify principal $G$-bundles.
The cases that are of greatest interest to us are when $G=O_{n}$ or $U_{n}$, because with some additional work one can show that maps into $BO_{n}$ classify real vector bundles of real dimension $n$ and similarly maps into $BU_{n}$ classify complex vector bundles of complex dimension $n$. We will come to this when we discuss real and complex vector bundles.
There is one other construction that will be very useful to us.
Definition 10.12.
Suppose that $E\longrightarrow B$ is a principal $G$-bundle and $F$ is a $G$-space with an effective [^49] This means that if $gf=f$ for all $f$ then $g$ is the identity element of the group. left $G$-action. Then $G$ acts on $E\times F$ by
$g(e,f)=(eg^{-1},gf).$
Define the associated bundle
$E\times_{G}F=(E\times F)/G\longrightarrow B.$
The associated bundle is a fibre bundle over $B$ with fibre $F$ and group $G$. Every fibre bundle with group $G$ arises in this manner.
Here are a very important couple of examples. If we take $G=O_{n}$ and $F={\mathbb{R}}^{n}$ with the usual action of $O_{n}$ on ${\mathbb{R}}^{n}$ then for any $G$-space $Y$,
$Y\times_{G}{\mathbb{R}}^{n}\longrightarrow B$
will be a real vector bundle of dimension $n$. Similarly, if $G=U_{n}$ acts on ${\mathbb{C}}^{n}$ then we obtain the associated complex vector bundle
$Y\times_{G}{\mathbb{C}}^{n}\longrightarrow B$
We need to carefully construct the topological space $\mathbf{BO}$. Hatcher [40] and Husemoller [44] are excellent references for this section and May [59] gives a very useful presentation. This is a very brief summary. The construction of the spaces $\mathbf{BU}$ and $\mathbf{BSp}$ in the complex and symplectic settings is done in the same way and we omit discussion of them. Along the way we will discover some very important fibre bundles.
Form the Grassmann manifold $G_{k}({\bf{R}}^{n})$, which consists of the space of all real $k$-dimensional vector subspaces through the origin in ${\mathbb{R}}^{n}$. It is a based compact Hausdorff space and a finite CW-complex. There are natural inclusion maps
$G_{k}({\mathbb{R}}^{n})\longrightarrow G_{k}({\mathbb{R}}^{n+1})$
and we define $BO_{k}$ to be their union. The name is not accidental:
Proposition 10.13.
(May [59] pp. 184-5)
  1. (1)
    There is a contractible $O_{n}$-space $EO_{n}$ and a universal principal $O_{n}$-bundle
    $O_{n}\longrightarrow EO_{n}\longrightarrow BO_{n}$
    and similarly for $BU_{n}$ and $BSp_{n}$. [^50] For example, we have shown that there is a fibre bundle ${\mathbb{Z}}_{2}\longrightarrow S^{\infty}\longrightarrow RP^{\infty}$ with $S^{\infty}$ contractible, and thus $B{\mathbb{Z}}_{2}\cong RP^{\infty}$. There is an analogous construction that shows that $BS^{1}\cong CP^{\infty}$.
  2. (2)
    There are natural inclusion maps
    $G_{k}({\mathbb{R}}^{n})\longrightarrow G_{k+1}({\mathbb{R}}^{n+1})$
    which induce inclusion maps
    $BO_{k}\longrightarrow BO_{k+1}.$
[…]
    $\mathbf{O}\longrightarrow E\mathbf{O}\longrightarrow\mathbf{BO}.$
  3. (3)
    The analogous constructions for the unitary groups yields a universal principal $\mathbf{U}$- bundle
    $\mathbf{U}\longrightarrow E\mathbf{U}\longrightarrow\mathbf{BU}.$
  4. (4)
    The analogous constructions for the symplectic groups yields a universal principal $\mathbf{Sp}$-bundle
    $\mathbf{Sp}\longrightarrow E\mathbf{Sp}\longrightarrow\mathbf{BSp}.$
Corollary 10.14.
The natural maps
[…]
and
[…]
are isomorphisms.
This implies by Proposition 10.7 that $\Omega BO_{n}\simeq O_{n}$ and hence
$\pi_{j}(BO_{n})\cong\pi_{j-1}(O_{n})\qquad{\text{for each}}\,\,j.$
This implies by Proposition 10.7 that
$\Omega\mathbf{BO}\simeq\mathbf{O}\qquad and\qquad\pi_{n}(\mathbf{BO})\cong\pi_{n-1}(\mathbf{O}).$
It turns out that $\mathbf{BO}$ has natural addition and multiplication maps which turn any $X,\mathbf{BO}]^{u}$ into a commutative semiring. Here is a sketch of the construction.
The addition operation ultimately arises from the maps
$O_{j}\times O_{k}\longrightarrow O_{j+k}$
given by
$(A,B)\longrightarrow\left[\begin{matrix}A&0\\ 0&B\end{matrix}\right]$
which leads to a map $\mathbf{O}\times\mathbf{O}\to\mathbf{O}$ and then to
$\mathbf{BO}\times\mathbf{BO}\overset{\alpha}{\longrightarrow}\mathbf{BO}.$
The multiplication map is more complicated. It arises from the Kronecker product of matrices
$O_{j}\times O_{k}\longrightarrow O_{jk}\qquad denoted\qquad M\times N\longrightarrow M\otimes N$
which gives the matrix of the tensor product of linear transformations. This leads to a map
$\mathbf{BO}\times\mathbf{BO}\overset{\mu}{\longrightarrow}\mathbf{BO}.$
11. Bott Periodicity Revisited; First View of the Homotopy Road Map
One of the principal goals of this work is to plot a path that takes us from classical Bott Periodicity to physics, stopping along the way at Clifford algebras, symmetric spaces and Cartan classification, and the TCS physics symmetries. Our guide will be the Homotopy Road Map which we will construct along the way. Step One takes us from Bott Periodicity to the Big Ten spaces.
We have glossed over the proofs of real and complex Bott Periodicity. There are many proofs and we have referenced several. In the previous sections we have seen that real Bott Periodicity can be stated in terms of
$KO_{j+8}(A)\cong KO_{j}(A)$
for all real $C^{*}$-algebras, or equivalently as
$\Omega^{8}\mathbf{O}\simeq\mathbf{O}.$
Similarly, complex Bott periodicity can be stated as
$K_{j+2}(A)\cong KO_{j}(A)$
for all complex $C^{*}$-algebras, or equivalently as
$\Omega^{\mathbf{U}}\simeq\mathbf{U}.$
We need to take a much more careful look at the individual spaces $\Omega^{k}\mathbf{O}$ and $\Omega^{k}\mathbf{U}$. We are going to simply state the result that we need, and we refer the reader to Theorem 18.3 and to Part IV of Milnor, Morse Theory [62] for details.
Definition 11.1.
  1. (1)
    Define
    ${\mathcal{R}}_{0}=\mathbf{BO}\times{\mathbb{Z}}\qquad\qquad{\mathcal{R}}_{1}=\mathbf{O}$
    and for $j=2,3,\dots 7$ define ${\mathcal{R}}_{j}=\Omega^{j-1}\mathbf{O}$.
  2. (2)
    Similarly, define
    ${\mathcal{R}}_{0}^{c}=\mathbf{BU}\times{\mathbb{Z}}\qquad\qquad{\mathcal{R}}_{1}^{c}=\mathbf{U}$
Here is the Homotopy Road Map, followed by an explanation of its entries.
Table 11.2.
The Homotopy Road Map: Step I
finite approx Big Ten Bott ${\mathcal{R}}_{k}$ $\pi_{0}({\mathcal{R}}_{k})$
$_{n} 𝐔\mathbf{U}$ $\Omega^{0}\mathbf{U}$ ${\mathcal{R}}_{1}^{c}$ 0
$U_{2n}/(U_{n}\times U_{n})$ $\mathbf{BU}\times{\mathbb{Z}}$ $\Omega^{1}\mathbf{U}$ ${\mathcal{R}}_{0}^{c}$ $\mathbb{Z}$}
$_{n} 𝐎\mathbf{O}$ $\Omega^{0}\mathbf{O}$ ${\mathcal{R}}_{1}$ ${\mathbb{Z}}_{2}$
$O_{2n}/U_{n}$ $\mathbf{O}/\mathbf{U}$ $\Omega^{1}\mathbf{O}$ ${\mathcal{R}}_{2}$ ${\mathbb{Z}}_{2}$
$U_{n}/Sp_{n}$ $\mathbf{U}/\mathbf{Sp}$ $\Omega^{2}\mathbf{O}$ ${\mathcal{R}}_{3}$ 0
$Sp_{2n}/(Sp_{n}\times Sp_{n})$ $\mathbf{BSp}\times{\mathbb{Z}}$ $\Omega^{3}\mathbf{O}$ ${\mathcal{R}}_{4}$ $\mathbb{Z}$}
$_{n} 𝐒𝐩\mathbf{Sp}$ $\Omega^{4}\mathbf{O}$ ${\mathcal{R}}_{5}$ 0
$Sp_{n}/U_{n}$ $\mathbf{Sp}/\mathbf{U}$ $\Omega^{5}\mathbf{O}$ ${\mathcal{R}}_{6}$ 0
$U_{n}/O_{n}$ $\mathbf{U}/\mathbf{O}$ $\Omega^{6}\mathbf{O}$ ${\mathcal{R}}_{7}$ 0
$O_{2n}/(O_{n}\times O_{n})$ $\mathbf{BO}\times{\mathbb{Z}}$ $\Omega^{7}\mathbf{O}$ ${\mathcal{R}}_{0}$ $\mathbb{Z}$}
  1. (1)
    The first column indicates the family of Lie groups or homogeneous spaces associated with the row. These are all compact spaces.
  2. (2)
    The second column indicates description of the limiting Big Ten [^51] The second author grew up following Big Ten football and couldn’t resist the opportunity to call the ten spaces by their true name. homogeneous space in terms of the iterated loop space of the orthogonal group (or the unitary group, for the first two rows). Note that in the three cases $\mathbf{BU}\times{\mathbb{Z}}$, $\mathbf{BSp}\times{\mathbb{Z}}$, and $\mathbf{BO}\times{\mathbb{Z}}$, the space mentioned is the limit of the unstable space times the integers. So what is true, for example in the bottom row, is that
    1. (a)
      $O_{2n}/(O_{n}\times O_{n})\longrightarrow\mathbf{O}/(\mathbf{O}\times\mathbf{O})\simeq\mathbf{BO}$
      and
    2. (b)
      $\mathbf{BO}\times{\mathbb{Z}}={\mathcal{R}}_{0}\simeq\Omega^{7}\mathbf{O}$
  3. (3)
    The third column gives the identification of the Big Ten speces in terms of iterated loop spaces of $\mathbf{U}$ and of $\mathbf{O}$ and illustrates their essential tie with Bott Periodicity. This will be explained in §17. For the moment we ask that it be taken on faith.
  4. (4)
    The fourth column indicates the name of the space in the physics literature. We will use this terminology with the hope that it will make the connection with physics clearer.
  5. (5)
    The fifth column indicates $\pi_{0}({\mathcal{R}}_{j})$. Remember that saying that $\pi_{0}=0$ means that the space is path connected. Saying that $\pi_{0}={\mathbb{Z}}_{2}$ says that the space has two connected components. If $\pi_{0}={\mathbb{Z}}$ then there are an infinite number of path components.
Guideposts along the way:
  1. (1)
    Bott Periodicity:
    $\Omega^{2}\mathbf{U}\simeq\mathbf{U}\qquad\qquad\Omega^{8}\mathbf{O}\simeq\mathbf{O}$
  2. (2)
    . Notation:
    $\Omega{\mathcal{R}}_{j}\simeq{\mathcal{R}}_{j+1}\qquad mod\,8\qquad\qquad\Omega{\mathcal{R}}_{j}^{c}\simeq{\mathcal{R}}_{j+1}^{c}\qquad mod\,2$
  3. (3)
    Restatement of Bott Periodicity:
    ${\pi_{j}({\mathcal{R}}_{k})=\pi_{0}({\mathcal{R}}_{j+k})}\cong\pi_{j+k}({\mathcal{R}}_{0})\qquad mod\,8\,\,$
    and
    ${\pi_{j}({\mathcal{R}}_{k}^{c})=\pi_{0}({\mathcal{R}}_{j+k}^{c})}\cong\pi_{j+k}({\mathcal{R}}_{0}^{c})\qquad mod\,2\,\,$
  4. (4)
    There are natural maps
    ${\mathcal{R}}_{i}\wedge{\mathcal{R}}_{j}\longrightarrow{\mathcal{R}}_{i+j}$
    in the real case, (with indexing mod 8), and
    ${\mathcal{R}}_{i}^{c}\wedge{\mathcal{R}}_{j}^{c}\longrightarrow{\mathcal{R}}_{i+j}^{c}$
    in the complex case (with indexing mod 2). These maps correspond to tensor product of vector bundles.
Take for example the fourth row. The group $U_{n}$ may be included naturally in the group $O_{2n}$ with homogeneous space $O_{2n}/U_{n}$. As $n\to\infty$ we obtain the homogeneous space $\mathbf{O}/\mathbf{U}$. It turns out that $\mathbf{O}/\mathbf{U}$ is homotopy equivalent to $\Omega^{1}\mathbf{O}={\mathcal{R}}_{2}$ and so
$\pi_{0}(\mathbf{O}/\mathbf{U})=\pi_{0}({\mathcal{R}}_{2})\cong{\mathbb{Z}}_{2}.$
The identification of the spaces ${\mathcal{R}}_{j}$ with the homogeneous spaces mentioned follows from some of the proofs of Bott periodicity. We shall explore this in much more detail when we reach Clifford algebras (see Theorem 18.3 in particular). So for the moment consider this theorem as a taste of the Homotopy Road Map, which will stretch from homotopy land and Bott periodicity to the 10-fold classification of symmetric spaces and to the 10-fold way in physics.
Each of the spaces ${\mathcal{R}}_{j}$ is a CW-complex and that means, for example, that the map
[…]
is an isomorphism. However, that does NOT mean that $\pi_{j}(O_{n})\to\pi_{j}(\mathbf{O})$ is an isomorphism for all $n$. Here’s an example: we calculate $\pi_{5}(O_{k})$ for all $k$.
[…]
12. Next Contact with Physics: Low Homotopy Groups of the Ten-Fold Way
In this section we will compute $\pi_{0}$, $\pi_{1}$ and $\pi_{2}$ of the classical Lie groups and homogeneous spaces that arise in the Ten-Fold Way. This section contains material that is very well-known in the math world but not in the physics world (yet). [^54] Lest we give the wrong impression, a great deal more is known about the higher homotopy groups of compact Lie groups and their associated homogeneous spaces. We are simply looking at the lowest dimensional calculations. It is useful there because some physicists (cf. Orion and Akkermans [66] ) need not just the “stable” results provided by the Bott Periodicity theorem but also the low-dimensional results to which this section is devoted. It is fascinating to see how soon the Bott results take shape here.
Suppose that $G$ is a compact Lie group and $H$ is a closed subgroup. Then
$H\longrightarrow G\longrightarrow G/H$
is a fibre bundle and there is an associated long exact sequence in homotopy that in low dimensions takes the form
$0=\pi_{2}(G)\to\pi_{2}(G/H)\to\pi_{1}(H)\to\pi_{1}(G)\to\pi_{1}(G/H)\to\qquad(*)$
$\to\pi_{0}(H)\to\pi_{0}(G)\to\pi_{0}(G/H)\to 0$
The group $\pi_{2}(G)=0$ for all compact Lie groups as previously mentioned. If the group $H$ is path-connected [^55] A space $X$ is path-connected if $\pi_{0}(X)=0$. so that $\pi_{0}(H)=0$, then the sequence breaks apart into two simpler sequences-
$0\to\pi_{2}(G/H)\to\pi_{1}(H)\to\pi_{1}(G)\to\pi_{1}(G/H)\to 0\qquad(**)$
and
$0\to\pi_{0}(G)\to\pi_{0}(G/H)\to 0$
so that $\pi_{0}(G)\cong\pi_{0}(G/H)$.
These sequences are natural. That is, given a map (i.e. a continuous homomorphism) of compact groups $f:G\to G^{\prime}$, closed subgroups $H\subset G$ and $H^{\prime}\subset G^{\prime}$ and a commuting diagram
$HfG\begin{CD}$$V{}V{}V\\ H^{\prime}@>{f}>{}>G^{\prime}\end{CD}$
then there is a commuting diagram of long exact sequences.
It’s easy to show that if $X$ and $Y$ are topological spaces then
$\pi_{j}(X\times Y)=\pi_{j}(X)\oplus\pi_{j}(Y)\qquad\forall j.$
We will use this fact many times in what comes below. Another fact to keep in mind is that homotopy groups revolve about basepoint-preserving maps. This implies that for $j>0$ that $\pi_{j}(G)$ depends only on the path-component of $G$ that contains the basepoint, which we always take to be the identity when dealing with topological groups. In particular, the inclusion $SO_{n}\to O_{n}$ induces isomorphisms
$\pi_{j}(SO_{n})\overset{\cong}{\longrightarrow}\pi_{j}(O_{n})\qquad\qquad j>0.$
Here is an outline of this section. Note that there are ten groups of spaces and these correspond to the spaces ${\mathcal{R}}_{j}$ discussed in the previous section.
  1.
  2. (1)
    The Groups $U_{n}$ and $SU_{n}$
  3. (2)
    The Homogenous Spaces $U_{n}/(U_{k}\times U_{n-k})$ and $SU_{n}/(SU_{k}\times SU_{n-k})$
  4. (3)
    The Groups $O_{n}$ and $SO_{n}$
  5. (4)
    The Homogeneous Spaces $O_{2n}/U_{n}$, $SO_{2n}/SU_{n}$, and $SO_{2n}/U_{n}$
  6. (5)
    The Homogeneous Spaces $U_{n}/Sp_{n}$ and $SU_{n}/Sp_{n}$
  7. (6)
    The Homogeneous Spaces $Sp_{n}/(Sp_{k}\times Sp_{n-k})$
  8. (7)
    The Groups $Sp_{n}$
  9. (8)
    The Homogeneous Spaces $Sp_{n}/U_{n}$ and $Sp_{n}/SU_{n}$
  10. (9)
    The Homogeneous Spaces $U_{n}/O_{n}$, $SU_{n}/SO_{n}$ and $U_{n}/SO_{n}$
  11. (10)
    The Homogeneous Spaces $O_{n}/(O_{k}\times O_{n-k})$ and $SO_{n}/(SO_{k}\times SO_{n-k})$
1. The Groups $U_{n}$ and $SU_{n}$
The group $U_{n}$ is path-connected for all $n$. Note that $U_{1}\cong S^{1}$ and maps from $S^{1}$ into $U_{n}$ factor up to homotopy through the $2$-skeleton, which is $S^{1}$, by cellular approximation. (For example, $U_{2}$ looks like $U_{1}$ with an additional $S^{3}$ attached). Hence
$\pi_{1}(U_{n})=\pi_{1}(U_{1})={\mathbb{Z}}\qquad\forall n\geq 1.$
The group $SU_{n}$ is simply connected, [^56] A space $X$ is simply connected if $\pi_{1}(X)=0$. In the case of $SU_{n}$, even more is true. We have $SU_{1}=0$, $SU_{2}\cong S^{3}$ and, for larger $n$, $SU_{n}$ is obtained from $SU_{n-1}$ by adjoining even higher-dimensional cells. and hence for all $n$ we have
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$U_{n}$ 0 $\mathbb{Z}$} 0
$SU_{n}$ 0 0 0
These results connect up with the stable group ${\mathcal{R}}_{1}^{c}=\mathbf{U}$ since $\lim U_{n}=\mathbf{U}.$
2. The Homogenous Spaces $U_{n}/(U_{k}\times U_{n-k})$ and $SU_{n}/(SU_{k}\times SU_{n-k})$
To eliminate a trivial case and get rid of duplications, let us assume throughout that $0<k\leq n-k$.
The map
$\phi=diag(A,B):U_{k}\times U_{n-k}\to U_{n}$
takes a pair $(A,B)$ of matrices and places them in the upper left and lower right quarters of an $n\times n$ matrix (with zeros in the remaining two quarters).
First we consider the path-connected homogeneous spaces $U_{n}/(U_{k}\times U_{n-k})$. The map
$\phi=diag(A,B):U_{k}\times U_{n-k}\to U_{n}$
takes a pair $(A,B)$ of matrices and places them in the upper left and lower right parts of an $n\times n$ matrix (with zeros in the remaining two quarters). Sequence (**) applies, and so we have
$0\to\pi_{2}(U_{n}/(U_{k}\times U_{n-k}))\to\pi_{1}(U_{k}\times U_{n-k})\overset{\phi_{*}}{\longrightarrow}\pi_{1}(U_{n})\to\pi_{1}(U_{n}/(U_{k}\times U_{n-k}))\to 0$
Plugging in what we know gives
$0\to\pi_{2}(U_{n}/(U_{k}\times U_{n-k}))\to{\mathbb{Z}}^{2}\overset{\phi_{*}}{\longrightarrow}{\mathbb{Z}}\to\pi_{1}(U_{n}/(U_{k}\times U_{n-k}))\to 0$
Let $f:S^{1}\to U_{k}$ represent the generator of $\pi_{1}(U_{k})$ and let
$\delta:U_{k}\to U_{k}\times U_{n-k}$
be given by $\delta(x)=(x,1)$. Then the map $\phi_{*}\delta_{*}(f)$ generates $\pi_{1}(U_{n})\cong{\mathbb{Z}}$ and hence $\phi_{*}$ is onto. This implies that
$\pi_{1}(U_{n}/(U_{k}\times U_{n-k}))=0\qquad and\qquad\pi_{2}(U_{n}/(U_{k}\times U_{n-k}))={\mathbb{Z}}.$
Determining $\pi_{j}(SU_{n}/(SU_{k}\times SU_{n-k}))$ is easy: it is zero for $j\leq 2$ because $\pi_{j}(SU_{k})=0$ for $j\leq 2$.
We pause to note that what we are seeing here is a small manifestation of Bott periodicity. The space $U_{n}$ stabilizes to the infinite unitary group $\mathbf{U}$ as $n\to\infty$ , and the homogeneous spaces $U_{2n}/(U_{n}\times U_{n})$ stabilize to the classifying space $\mathbf{BU}$ and $\mathbf{BU}\times{\mathbb{Z}}\cong{\mathcal{R}}_{0}^{c}$ as $n\to\infty$. Periodicity implies that $\pi_{1}(\mathbf{U})\cong\pi_{0}(\mathbf{BU})$ and that is what we are seeing already at the earliest stage:
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$U_{n}$ 0 $\mathbb{Z}$} 0
$U_{n}/(U_{k}\times U_{n-k})\,\,\,\times{\mathbb{Z}}$ $\mathbb{Z}$} 0 $\mathbb{Z}$} $0<k<n$
$SU_{n}$ 0 0 0
$SU_{n}/(SU_{k}\times SU_{n-k})\,\,\,\times{\mathbb{Z}}$ $\mathbb{Z}$} 0 0 $0<k<n$
Bott periodicity gives us
$\pi_{n}(\mathbf{U})\cong\pi_{n+2}(\mathbf{U})$
and putting that together with the easy statement that $\Omega(\mathbf{BU})\simeq\mathbf{U}$ gets us the full complex periodicity statement. To see this when looking at $SU_{n}/(SU_{k}\times SU_{n-k})$ one has to go to higher homotopy groups.
3. The Groups $O_{n}$ and $SO_{n}$
The groups $O_{n}$ have two path components, rotations and reflections, and $SO(n)$ is the closed, index $2$ subgroup of rotations. Thus
$\pi_{0}(O_{n})={\mathbb{Z}}_{2}\qquad\qquad\pi_{0}(SO_{n})=0\qquad\forall n\geq 1,$
and
$\pi_{j}(O_{n})=\pi_{j}(SO_{n})\qquad\qquad\forall j\geq 1.$
Now $SO_{2}\cong S^{1}$ as is easy to see. Less obvious [^57] Steenrod ([80] §22) has a detailed exposition of this (22.3) and related matters. is the fact that $SO_{3}\cong RP^{3}$. This implies that
$\pi_{1}(O_{1})\cong\pi_{1}(SO_{1})=0\qquad\pi_{1}(O_{2})\cong\pi_{1}(SO_{2})={\mathbb{Z}}\qquad\pi_{1}(O_{3})\cong\pi_{1}(SO_{3})={\mathbb{Z}}_{2}$
and, by cellular approximation,
$\pi_{1}(O_{n})\cong\pi_{1}(SO_{n})={\mathbb{Z}}_{2}\quad\forall n\geq 3.$
To summarize:
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$O_{1}$ ${\mathbb{Z}}_{2}$ 0 0
$O_{2}$ ${\mathbb{Z}}_{2}$ $\mathbb{Z}$} 0
$O_{n}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $n\geq 3$
$SO_{1}$ 0 0 0
$SO_{2}$ 0 $\mathbb{Z}$} 0
$SO_{n}$ 0 ${\mathbb{Z}}_{2}$ 0 $n\geq 3$
These results tie to the group ${\mathcal{R}}_{1}=\mathbf{O}$, of course, since $\lim O_{n}\cong\mathbf{O}$.
As mentioned earlier, much more is known about these groups. For example, Steenrod shows [80] that we may extend the chart as follows. (Note $\pi_{j}(O_{n})=\pi_{j}(SO_{n})$ for $j>0$ since $SO_{n}$ is the path-component of the identity of $O_{n}$. )
$X$ $\pi_{3}(X)$ $\pi_{4}(X)$ $\pi_{5}(X)$
$O_{3}$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$
$O_{4}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}$
$O_{5}$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$
$O_{k}$ $\mathbb{Z}$} 0 $\mathbb{Z}$} $k>5$
4. The Homogeneous Spaces $O_{2n}/U_{n}$, $SO_{2n}/SU_{n}$, and $SO_{2n}/U_{n}$
Let
$\phi_{n}:U_{n}\to O_{2n}$
denote the standard inclusion. The groups $U_{n}$ are path-connected for all $n$. So we have
$0\to\pi_{2}(O_{2n}/U_{n})\to\pi_{1}(U_{n})\overset{(\phi_{n})_{*}}{\to}\pi_{1}(O_{2n})\to\pi_{1}(O_{2n}/U_{n})\to 0\to\pi_{0}(O_{2n})\to\pi_{0}(O_{2n}/U_{n})\to 0.\qquad(**)$
First, consider the case $n=1$. Then the sequence becomes
$0\to\pi_{2}(O_{2}/U_{1})\to{\mathbb{Z}}\overset{\phi_{*}}{\longrightarrow}{\mathbb{Z}}\to\pi_{1}(O_{2}/U_{1})\to 0\to{\mathbb{Z}}_{2}\to\pi_{0}(O_{2}/U_{1})\to 0$
The map $\phi_{*}$ is an isomorphism since $U_{1}=SO_{2}\subset O_{2}$ and this implies that
$\pi_{0}(O_{2}/U_{1})\cong{\mathbb{Z}}_{2}\qquad\pi_{1}(O_{2}/U_{1})=\pi_{2}(O_{2}/U_{1})=0.$
Now consider the case $n>1$. Then the sequence becomes
$0\to\pi_{2}(O_{2n}/U_{n})\to{\mathbb{Z}}\overset{\phi_{*}}{\longrightarrow}{\mathbb{Z}}_{2}\to\pi_{1}(O_{2n}/U_{n})\to 0\to{\mathbb{Z}}_{2}\to\pi_{0}(O_{2n}/U_{n})\to 0$
[…]
$\pi_{1}(O_{2n}/U_{n})=0\qquad and\qquad\pi_{2}(O_{2n}/U_{n})={\mathbb{Z}}$
for all $n$.
For $SO_{2n}/SU_{n}$ we apply sequence (**) since $SO_{2n}$ is path-connected. We get
$0\to\pi_{2}(SO_{2n}/SU_{n})\to\pi_{1}(SU_{n})\longrightarrow\pi_{1}(SO_{2n})\to\pi_{1}(SO_{2n}/SU_{n})\to 0$
which simplifies to
$0\to\pi_{2}(SO_{2n}/SU_{n})\to 0\longrightarrow\pi_{1}(SO_{2n})\to\pi_{1}(SO_{2n}/SU_{n})\to 0$
Thus $\pi_{2}(SO_{2n}/SU_{n})=0$ and
$\pi_{1}(SO_{2n}/SU_{n})\cong\pi_{1}(SO_{2n})$
which is $\mathbb{Z}$} for $n=1$ and ${\mathbb{Z}}_{2}$ for $n>1$.
Finally, we consider $SO_{2n}/U_{n}$. The case $n=1$ is easy, since $U_{1}=SO_{2}$ and so all of the homotopy groups vanish. So assume that $n>1$. As $SO_{2n}$ is path-connected we may use sequence **. It reads
$0\to\pi_{2}(SO_{2n}/U_{n})\longrightarrow\pi_{1}(U_{n})\overset{\phi_{*}}{\longrightarrow}\pi_{1}(SO_{2n})\longrightarrow\pi_{1}(SO_{2n}/U_{n})\to 0$
which simplifies to
$0\to\pi_{2}(SO_{2n}/U_{n})\longrightarrow{\mathbb{Z}}\overset{\phi_{*}}{\longrightarrow}{\mathbb{Z}}_{2}\longrightarrow\pi_{1}(SO_{2n}/U_{n})\to 0$
To determine the map $\phi$, we can consider the commutative diagram
[…]
which implies that the map $\phi$ is onto. Thus
$\pi_{1}(SO_{2n}/U_{n})=0\qquad and\qquad\pi_{2}(SO_{2n}/U_{n})\cong{\mathbb{Z}}.$
In summary,
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$O_{2}/U_{1}$ ${\mathbb{Z}}_{2}$ 0 0
$O_{2n}/U_{n}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} $n\geq 2$
$SO_{2}/SU_{1}$ $\mathbb{Z}$} $\mathbb{Z}$} 0
$SO_{2n}/SU_{n}$ 0 ${\mathbb{Z}}_{2}$ 0 $n\geq 2$
$SO_{2n}/U_{n}$ 0 0 $\mathbb{Z}$}
These results connect up with ${\mathcal{R}}_{2}=\lim O_{2n}/U_{n}$.
5. The Homogeneous Spaces $U_{n}/Sp_{n}$ and $SU_{n}/Sp_{n}$.
These are path-connected spaces. We have $\pi_{j}(Sp_{n})=0$ for $j\leq 2$ (see §7 ), so sequence (**) simplifies to
$0\to\pi_{2}(U_{n}/Sp_{n})\to 0\to{\mathbb{Z}}\to\pi_{1}(U_{n}/Sp_{n})\to 0.\qquad(**)$
Hence
$\pi_{0}(U_{n}/Sp_{n})=0\qquad\pi_{1}(U_{n}/Sp_{n})\cong{\mathbb{Z}}\qquad\pi_{2}(U_{n}/Sp_{n})\cong 0\qquad\forall n$
For $SU_{n}/Sp_{n}$, the sequence simplifies to
$0\to\pi_{2}(SU_{n}/Sp_{n})\to 0\to 0\to\pi_{1}(SU_{n}/Sp_{n})\to 0\qquad(**)$
and so
$\pi_{j}(SU_{n}/Sp_{n})=0\quad j\leq 2\qquad\forall\,\,n$
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$U_{n}/Sp_{n}$ 0 $\mathbb{Z}$} 0
$SU_{n}/Sp_{n}$ 0 0 0
These results tie in with ${\mathcal{R}}_{3}=\lim U_{n}/Sp_{n}$.
6. The Homogeneous Spaces $Sp_{n}/(Sp_{k}\times Sp_{n-k})$
We will see in the next section that $\pi_{j}(Sp_{n})=0$ for $j\leq 2$. This implies that in the sequence (*) that two out of every three groups is zero, and hence
$\pi_{j}(Sp_{n}/(Sp_{k}\times Sp_{n-k}))=0\qquad j=0,1,2\qquad\forall\,\,n$
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$Sp_{n}/(Sp_{k}\times Sp_{n-k})$ 0 0 0
These results connect with ${\mathcal{R}}_{4}={\bf{BSp}}\times{\mathbb{Z}}$ since $Sp_{2n}/(Sp_{n}\times Sp_{n})$ converges to $\bf{BSp}$}.
7. The Groups $Sp_{n}$
This is the easiest section to write. The groups $Sp_{n}$ have their lowest-dimensional cell in dimension 3, which implies that
$\pi_{j}(Sp_{n})=0\quad\forall\,\,n\quad j\leq 2.$
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$Sp_{n}$ 0 0 0
These groups connect up with ${\mathcal{R}}_{5}=\lim Sp_{n}$.
8. The Homogeneous Spaces $Sp_{n}/U_{n}$ and $Sp_{n}/SU_{n}$
The spaces $Sp_{n}/U_{n}$ are path-connected. Sequence (**) gives an exact sequence
$\to\pi_{2}(Sp_{n})\to\pi_{2}(Sp_{n}/U_{n})\to\pi_{1}(U_{n})\to\pi_{1}(Sp_{n})\to\pi_{1}(Sp_{n}/U_{n})\to 0$
which, after evaluating, becomes the sequence
$\to 0\to\pi_{2}(Sp_{n}/U_{n})\to{\mathbb{Z}}\to 0\to\pi_{1}(Sp_{n}/U_{n})\to 0$
from which we deduce that
$\pi_{1}(Sp_{n}/U_{n})=0\qquad\qquad\pi_{2}(Sp_{n}/U_{n})\cong{\mathbb{Z}}.$
For $Sp_{n}/SU_{n}$ sequence (**) gives an exact sequence
$0\to\pi_{2}(Sp_{n}/SU_{n})\to 0\to 0\to\pi_{1}(Sp_{n}/SU_{n})\to 0$
and so $\pi_{j}(Sp_{n}/SU_{n})=0$ for all $n$ and for $j\leq 2$.
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$Sp_{n}/U_{n}$ 0 0 $\mathbb{Z}$}
$Sp_{n}/SU_{n}$ 0 0 0
These groups relate to ${\mathcal{R}}_{6}=\lim Sp_{n}/U_{n}$.
9. The Homogeneous Spaces $U_{n}/O_{n}$, $SU_{n}/SO_{n}$ and $U_{n}/SO_{n}$
First consider $U_{n}/O_{n}$. As $O_{n}$ is not path-connected we use sequence (*) and obtain
$0\to\pi_{2}(U_{n}/O_{n})\to\pi_{1}(O_{n})\overset{\alpha}{\to}\pi_{1}(U_{n})\to\pi_{1}(U_{n}/O_{n})\to\pi_{0}(O_{n})\to\pi_{0}(U_{n})\to\pi_{0}(U_{n}/O_{n})\to 0$
The fact that $U_{n}$ is path-connected implies that $\pi_{0}(U_{n}/O_{n})=0$ for all $n$ and so the sequence simplifies to
$0\to\pi_{2}(U_{n}/O_{n})\to\pi_{1}(O_{n})\overset{\alpha}{\to}\pi_{1}(U_{n})\to\pi_{1}(U_{n}/O_{n})\to{\mathbb{Z}}_{2}\to 0$
Suppose first that $n=1$. Then $O_{1}\to U_{1}\to U_{1}/O_{1}$ is really just the double cover
${\mathbb{Z}}_{2}\longrightarrow S^{1}\overset{\pi}{\longrightarrow}U_{1}/O_{1}\cong S^{1}$
and so
$\pi_{0}(U_{1}/O_{1})=0\qquad\pi_{1}(U_{1}/O_{1})={\mathbb{Z}}\qquad\pi_{2}(U_{1}/O_{1})=0.$
Suppose next that $n=2$. The commuting diagram
[…]
shows that the map $\alpha:\pi_{1}(O_{2})\to\pi_{1}(U_{2})$ is an isomorphism. Thus the sequence (*) reduces down to
$0\to\pi_{2}(U_{n}/O_{n})\longrightarrow{\mathbb{Z}}\overset{\cong}{\longrightarrow}{\mathbb{Z}}\longrightarrow\pi_{1}(U_{n}/O_{n})\longrightarrow{\mathbb{Z}}_{2}\to 0$
which implies that
$\pi_{1}(U_{n}/O_{n})\cong{\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(U_{n}/O_{n})=0.$
Suppose next that $n>2$. Then we have
$0\to\pi_{2}(U_{n}/O_{n})\to{\mathbb{Z}}_{2}\overset{\alpha}{\to}{\mathbb{Z}}\overset{\beta}{\longrightarrow}\pi_{1}(U_{n}/O_{n})\to{\mathbb{Z}}_{2}\to 0$
The map $\alpha\equiv 0$ and so
$\pi_{2}(U_{n}/O_{n})={\mathbb{Z}}_{2}.$
We are left with
$0\to{\mathbb{Z}}\cong\pi_{1}(U_{n})\overset{\beta}{\longrightarrow}\pi_{1}(U_{n}/O_{n})\to{\mathbb{Z}}_{2}\to 0.$
This implies algebraically that $\pi_{1}(U_{n}/O_{n})$ is either $\mathbb{Z}$} with $\beta$ being multiplication by $2$,or ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ with $\beta$ being isomorphic inclusion onto the first factor. We will settle this point in a moment.
We turn next to $SU_{n}/SO_{n}$. The space $SU_{n}$ is path-connected and so
$\pi_{0}(SU_{n}/SO_{n})=0.$
Turning to sequence (**) we obtain
$0\to\pi_{2}(SU_{n}/SO_{n})\to\pi_{1}(SO_{n})\overset{\alpha}{\to}\pi_{1}(SU_{n})\to\pi_{1}(SU_{n}/SO_{n})\to 0$
which becomes
$0\to\pi_{2}(SU_{n}/SO_{n})\to\pi_{1}(SO_{n})\overset{\alpha}{\to}0\to\pi_{1}(SU_{n}/SO_{n})\to 0$
Hence $\pi_{1}(SU_{n}/SO_{n})=0$ and $\pi_{2}(SU_{n}/SO_{n})\cong\pi_{1}(SO_{n})$. This group is $\mathbb{Z}$} when $n=2$ and ${\mathbb{Z}}_{2}$ for $n>2$.
We return to the identification of $\pi_{1}(U_{n}/O_{n})$ above. There is a commutative diagram
[…]
with each vertical column and horizontal row a fibration. Taking homotopy groups we obtain the following diagram with exact rows and columns
[…]
The map $\bar{\beta}$ is multiplication by $2$ which implies that the map $\zeta$ must be an inclusion of $\pi_{1}(U_{n}/O_{n})$ into $\mathbb{Z}$} and so $\pi_{1}(U_{n}/O_{n})\cong{\mathbb{Z}}$. This settles the extension question mentioned above.
Finally, consider $U_{n}/SO_{n}$. The space $U_{n}$ is path-connected. If $n=2$ then sequence (**) gives us
$0\to\pi_{2}(U_{2}/SO_{2})\longrightarrow\pi_{1}(SO_{2})\overset{\cong}{\longrightarrow}\pi_{1}(U_{2})\longrightarrow\pi_{1}(U_{2}/SO_{2})\to 0$
which simplifies [^59] Exercise 1: If $0\to H\to G_{1}\overset{\cong}{\longrightarrow}G_{2}$ is an exact sequence of abelian groups, then $H=0$. Exercise 2: If $G_{1}\overset{\cong}{\longrightarrow}G_{2}\to H\to 0$ is an exact sequence of abelian groups ,then $H=0$.
to
$0\to\pi_{2}(U_{2}/SO_{2})\longrightarrow{\mathbb{Z}}\overset{\cong}{\longrightarrow}{\mathbb{Z}}\longrightarrow\pi_{1}(U_{2}/SO_{2})\to 0$
and hence $\pi_{j}(U_{2}/SO_{2})=0$ for $j\leq 2$.
If $n>2$ then the sequence
$0\to\pi_{2}(U_{n}/SO_{n})\longrightarrow\pi_{1}(SO_{n})\longrightarrow\pi_{1}(U_{n})\longrightarrow\pi_{1}(U_{n}/SO_{n})\to 0$
simplifies to
$0\to\pi_{2}(U_{n}/SO_{n})\longrightarrow{\mathbb{Z}}_{2}\longrightarrow{\mathbb{Z}}\longrightarrow\pi_{1}(U_{n}/SO_{n})\to 0.$
The map $Z_{2}\to Z$ is the trivial map, of course, and so
$\pi_{1}(U_{n}/SO_{n})\cong{\mathbb{Z}}\qquad and\qquad\pi_{2}(U_{n}/SO_{n})\cong{\mathbb{Z}}_{2}$
To summarize:
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$U_{1}/O_{1}$ 0 ${\mathbb{Z}}_{2}$ 0
$U_{2}/O_{2}$ 0 ${\mathbb{Z}}_{2}$ 0
$U_{n}/O_{n}$ 0 $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ $n\geq 3$
$SU_{2}/SO_{2}$ 0 0 $\mathbb{Z}$}
$SU_{n}/SO_{n}$ 0 0 ${\mathbb{Z}}_{2}$ $n\geq 3$
$U_{2}/SO_{2}$ 0 0 0
$U_{n}/SO_{n}$ 0 $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ $n\geq 3$
These groups relate to ${\mathcal{R}}_{7}=\lim U_{n}/O_{n}$.
10. The Homogeneous Spaces $O_{n}/(O_{k}\times O_{n-k})$ and $SO_{n}/(SO_{k}\times SO_{n-k})$.
We will need the following information over and over again in this section, so we copy it from Section 5.
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$O_{1}$ ${\mathbb{Z}}_{2}$ 0 0
$O_{2}$ ${\mathbb{Z}}_{2}$ $\mathbb{Z}$} 0
$O_{n}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $n\geq 3$
$SO_{1}$ 0 0 0
$SO_{2}$ 0 $\mathbb{Z}$} 0
$SO_{n}$ 0 ${\mathbb{Z}}_{2}$ 0 $n\geq 3$
In what follows we will always assume that $0<k\leq n-k$.
Consider path-components first. We have
$\to\pi_{0}(O_{k}\times O_{n-k})\overset{\gamma}{\longrightarrow}\pi_{0}(O_{n})\to\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))\to 0$
which simplifies to
$\to{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\overset{\gamma}{\longrightarrow}{\mathbb{Z}}_{2}\to\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))\to 0.$
The map $\gamma$ is onto and hence $\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))=0$. A similar and easier argument works for $SO$ and hence
$\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))=0\qquad and\qquad\pi_{0}(SO_{n}/(SO_{k}\times SO_{n-k}))=0$
for $0<k\leq n-k<n$.
Next we consider $\pi_{1}$ and $\pi_{2}$ for seven special cases.
  1. (1)
    The spaces $O_{2}/(O_{1}\times O_{1})$ and $SO_{2}/(SO_{1}\times SO_{1})$
    There is a natural identification
    $O_{2}/O_{1}\cong SO_{2}/SO_{1}=SO_{2}=S^{1}$
    with associated fibration
    $O_{1}\longrightarrow O_{2}/O_{1}\longrightarrow O_{2}/(O_{1}\times O_{1})$
    which we may identify with the double cover
    ${\mathbb{Z}}_{2}\longrightarrow S^{1}\longrightarrow O_{2}/(O_{1}\times O_{1})$
    Thus $O_{2}/(O_{1}\times O_{1})\cong S^{1}$ and hence
    $\pi_{1}(O_{2}/(O_{1}\times O_{1}))\cong{\mathbb{Z}}\qquad and\qquad\pi_{2}(O_{2}/(O_{1}\times O_{1})=0.$
    The fibration
    $SO_{1}\longrightarrow SO_{2}/SO_{1}\longrightarrow SO_{2}/(SO_{1}\times SO_{1})$
    trivializes since $SO_{1}$ is one point, so $SO_{2}/(SO_{1}\times SO_{1})\cong S^{1}$ and hence
    $\pi_{1}(SO_{2}/(SO_{1}\times SO_{1}))\cong{\mathbb{Z}}\qquad and\qquad\pi_{2}(SO_{2}/(SO_{1}\times SO_{1})=0.$
  2. (2)
    The space $O_{3}/(O_{1}\times O_{2})$
    We need some preliminary calculations. First of all,
    $O_{3}/O_{2}\cong SO_{3}/SO_{2}\cong RP^{3}/S^{1}.$
    The long exact sequence of the associated fibration has the form
    $\to\pi_{2}(RP^{3})\to\pi_{2}(RP^{3}/S^{1})\to\pi_{1}(S^{1})\overset{p}{\longrightarrow}\pi_{1}(RP^{3})\to\pi_{1}(RP^{3}/S^{1})\to 0$
    which simplifies to
    $0\to\pi_{2}(RP^{3}/S^{1})\to{\mathbb{Z}}\overset{p}{\longrightarrow}{\mathbb{Z}}_{2}\to\pi_{1}(RP^{3}/S^{1})\to 0$
    with $p$ onto, and this implies that
    $\pi_{1}(RP^{3}/S^{1})=0\qquad and\qquad\pi_{2}(RP^{3}/S^{1})={\mathbb{Z}}$
    and, translating back, we have
    $\pi_{1}(O_{3}/O_{2})=0\qquad and\qquad\pi_{2}(O_{3}/O_{2})={\mathbb{Z}}$
    and
    $\pi_{1}(SO_{3}/SO_{2})=0\qquad and\qquad\pi_{2}(SO_{3}/SO_{2})={\mathbb{Z}}$
    Now, apply these results to the long exact sequence associated to the fibration
    $O_{1}\longrightarrow O_{3}/O_{2}\longrightarrow O_{3}/(O_{1}\times O_{2})$
    and one obtains
    $0\to\pi_{2}(O_{3}/O_{2})\cong{\mathbb{Z}}\longrightarrow\pi_{2}(O_{3}/(O_{1}\times O_{2}))\to 0$
    and
    $0\longrightarrow\pi_{1}(O_{3}/(O_{1}\times O_{2})\to\pi_{0}(O_{1})\cong{\mathbb{Z}}_{2}\to 0$
    so that
    $\pi_{1}(O_{3}/(O_{1}\times O_{2})={\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(O_{3}/(O_{1}\times O_{2})\cong{\mathbb{Z}}.$
  3. (3)
    The space $SO_{3}/(SO_{1}\times SO_{2})$
    Similarly, we have
    $0\to\pi_{2}(SO_{3}/SO_{2})\cong{\mathbb{Z}}\longrightarrow\pi_{2}(SO_{3}/(SO_{1}\times O_{2}))\to 0$
    and
    $0\longrightarrow\pi_{1}(SO_{3}/(SO_{1}\times SO_{2})\to\pi_{0}(SO_{1})\cong 0\to 0$
    so that
    $\pi_{1}(SO_{3}/(SO_{1}\times SO_{2})=0\qquad and\qquad\pi_{2}(SO_{3}/(SO_{1}\times SO_{2})\cong{\mathbb{Z}}.$
  4. (4)
    The space $O_{4}/(O_{2}\times O_{2})$
    Applying the sequence (*) to $O_{4}/(O_{2}\times O_{2})$ gives us
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to\pi_{1}(O_{2}\times O_{2})\to\pi_{1}(O_{4})\to\pi_{1}(O_{4}/(O_{2}\times O_{2}))\to$
    $\to\pi_{0}(O_{2}\times O_{2})\to\pi_{0}(O_{4})\to\pi_{0}(O_{4}/(O_{2}\times O_{2}))\to 0$
    which, after applying what we already know, becomes
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}^{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\overset{p}{\longrightarrow}\pi_{1}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\to 0$
    This implies that $p$ must be the zero map, and then we have
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}^{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\to 0$
    and
    $0\to\pi_{1}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\to 0$
    so that
    $\pi_{1}(O_{4}/(O_{2}\times O_{2}))\cong{\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(O_{4}/(O_{2}\times O_{2}))\cong{\mathbb{Z}}^{2}$
  5. (5)
    The space $SO_{4}/(SO_{2}\times SO_{2})$
    The sequence for $SO_{4}/(SO_{2}\times SO_{2})$ is simpler. We wind up with
    $0\to\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\longrightarrow\pi_{1}(SO_{2}\times SO_{2})\overset{onto}{\longrightarrow}\pi_{1}(SO_{4})\longrightarrow\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\to 0$
    which simplifies to
    $0\to\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\longrightarrow{\mathbb{Z}}^{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\overset{p}{\longrightarrow}\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\to 0$
    The map $p$ must be the zero map, and so
    $\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\cong 0\qquad and\qquad\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\cong{\mathbb{Z}}^{2}$
  6. (6)
    The space $O_{n}/(O_{k}\times O_{n-k})$
    Suppose as usual that $2<k\leq n-k$. Then the sequence
    $0\to\pi_{2}(O_{n}/(O_{k}\times O_{n-k}))\to\pi_{1}(O_{k}\times O_{n-k})\overset{\gamma}{\to}\pi_{1}(O_{n})\to\pi_{1}(O_{n}/(O_{k}\times O_{n-k}))\to$
    $\to\pi_{0}(O_{k}\times O_{n-k})\to\pi_{0}(O_{n})\to\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))$
    simplifies to
    $0\to\pi_{2}(O_{n}/(O_{k}\times O_{n-k}))\to{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\overset{\bar{\gamma}}{\longrightarrow}{\mathbb{Z}}_{2}\overset{\zeta}{\longrightarrow}\pi_{1}(O_{n}/(O_{k}\times O_{n-k}))\to{\mathbb{Z}}_{4}\overset{\gamma}{\longrightarrow}{\mathbb{Z}}_{2}\to 0$
    The map $\bar{\gamma}$ is onto, and hence $\zeta$ is the zero map. The map $\gamma$ is also onto. Thus
    $\pi_{1}(O_{n}/(O_{k}\times O_{n-k}))\cong{\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(O_{n}/(O_{k}\times O_{n-k}))\cong{\mathbb{Z}}_{2}.$
  7. (7)
    The space $SO_{n}/(SO_{k}\times SO_{n-k})$
    The case of $SO_{n}/(SO_{k}\times SO_{n-k})$ is much simpler. The space $SO(n)$ is path-connected and sequence (**) gives us
    $0\to\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))\longrightarrow\pi_{1}(SO_{k}\times SO_{n-k})\longrightarrow\pi_{1}(SO_{n})\to\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))\to 0.$
    Since $k>2$, the sequence simplifies to
    $0\to\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))\longrightarrow{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\to\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))\to 0.$
    and so
    $\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))=0\qquad and\qquad\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))={\mathbb{Z}}_{2}.$
To summarize:
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$O_{2}/(O_{1}\times O_{1})$ 0 $\mathbb{Z}$} 0
$O_{3}/(O_{1}\times O_{2})$ 0 ${\mathbb{Z}}_{2}$ $\mathbb{Z}$}
$O_{4}/(O_{2}\times O_{2})$ 0 ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}^{2}$
$O_{n}/(O_{k}\times O_{n-k})$ 0 ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ $2<k\leq n-k$
$SO_{2}/(SO_{1}\times SO_{1})$ 0 $\mathbb{Z}$} 0
$SO_{3}/(SO_{1}\times SO_{2})$ 0 0 $\mathbb{Z}$}
$SO_{4}/(SO_{2}\times SO_{2})$ 0 0 ${\mathbb{Z}}^{2}$
$SO_{n}/(SO_{k}\times SO_{n-k})$ 0 0 ${\mathbb{Z}}_{2}$ $2<k\leq n-k$
Here the connection is to ${\mathcal{R}}_{8}={\bf{BO}}$ and $O_{2n}/(O_{n}\times O_{n})$ converges to $\bf{BO}$}.
Summary: The Eight Real Groups and Homogeneous Spaces
The chart below summarizes the information we have about the stable behavior of the first three homotopy groups of the eight real groups and homogeneous spaces. As before, we insert the $\times{\mathbb{Z}}$” as needed and adjust the results accordingly.
Note that the bottom four rows, marked with **, are simply a repetition of the top four rows, inserted so the diagonal pattern is more evident.
$X$ $\pi_{0}(X)$ $\pi_{1}(X)$ $\pi_{2}(X)$
$O_{n}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $n\geq 3$
$O_{2n}/U_{n}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} $n\geq 2$
$U_{n}/Sp_{n}$ 0 $\mathbb{Z}$} 0
$Sp_{n}/(Sp_{k}\times Sp_{n-k})\times{\mathbb{Z}}$ $\mathbb{Z}$} 0 0
$Sp_{n}$ 0 0 0
$Sp_{n}/U_{n}$ 0 0 $\mathbb{Z}$}
$U_{n}/O_{n}$ 0 $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ $n\geq 3$
$O_{n}/(O_{k}\times O_{n-k})\times{\mathbb{Z}}$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ $2<k\leq n-k$
** $O_{n}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $n\geq 3$
** $O_{2n}/U_{n}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} $n\geq 2$
** $U_{n}/Sp_{n}$ 0 $\mathbb{Z}$} 0
** $Sp_{n}/(Sp_{k}\times Sp_{n-k})\times{\mathbb{Z}}$ $\mathbb{Z}$} 0 0
Look carefully and you will see that parts of Bott Periodicity are revealed. Imagine lines rising at $45^{o}$ from the lower left to the upper right.
13. The Basics: Real Vector Bundles
In this basic section we provide the definitions and results needed to define $KO^{*}(X)$ for compact spaces in the following section.
The elementary theory of complex vector bundles and $K^{*}(X)$ has been explicated beautifully, particularly by Atiyah [7], Husemoller [44], Blackadar [25], and more recently by Hatcher [40] and by Rørdam, Larsen, and Lausten [69]. [^60] We shall refer to these illuminaries as the “complex authorities”. So one is tempted to say “read these sources but everywhere you see complex insert real.” This approach falls apart when one gets to graded $K$-theory and more sophisticated matters. So we will briefly sketch out the real facts of life and urge the reader to look at the complex sources as well.
A separate question arises to the practical physicist: why isn’t this Chapter 1? Why do we need all this operator algebra stuff? Our response to this is mostly practical. First, we need to use algebras such as $C(X,M_{n}({\mathbb{R}}))$. For a deeper answer, consider that if we stick to real vector bundles then we can suspend but we cannot desuspend. We need the real commutative algebras such as $C_{o}^{{\mathbb{R}}}(i{\bf{R}})$ and $\real(iS^{1})$ to perform this trickery, and they are NOT of the form $C_{o}^{{\mathbb{R}}}(X)$ or $\real(X)$.
Deeper yet, we know from studying complex theory that deep results such as the Künneth Theorem and the Universal Coefficient Theorem are quite useful in many places in physics, and we will see that analogous real results will help us as well. Thus we need $KO$ for real $C^{*}$-algebras and not just for real vector bundles.
Having said that, we recognize that $KO^{*}(X)$ is important in its own right, as its use of vector bundles brings us much closer to topological and geometric properties that we need.
Definition 13.1.
A real vector bundle $(E,p,X)$, frequently denoted by its total space $E$, consists of a continuous map $p:E\to X$ of topological spaces such that
  1. (1)
    The space $E_{x}=p^{-1}(x)$ is a real vector space for each $x\in X$ called the fibre over $x$, and
  2. (2)
    The following local triviality condition holds: for each $x\in X$ there is an open set $U_{x}\subset X$ with $x\in U_{x}$ and a homeomorphism $h$ such that the diagram
[…]
    commutes, where $p_{1}$ is projection on the first coordinate. In addition, for each $x$ the induced map $h:E_{x}\to{\mathbb{R}}^{k}$ must be linear.
Note that we could avoid some work here by using principal $O_{n}$-bundles, as their associated ${\mathbb{R}}^{n}$-bundles are essentially real vector bundles. That is, a real vector bundle $E\longrightarrow X$ of dimension $n$ over some compact space is essentially equivalent to an associated bundle of the form
$Y\times_{O_{n}}{\mathbb{R}}^{n}\longrightarrow X.$
The simplest examples of vector bundles over $X$ are the product bundles
$X\times{\mathbb{R}}^{k}\to X$
denoted $\theta_{k}$ when $X$ is understood. [^61] There is a bit of debate about whether $\theta_{0}$ exists; if you believe that there is a vector space of dimension 0 then your answer is “yes”.
Definition 13.2.
Suppose that $(E,p,X)$ and $(E^{\prime},p^{\prime},X^{\prime})$ are real vector bundles. A morphism or map of real vector bundles is a commuting diagram
[…]
of continuous maps, where
${\tilde{f}}:E_{y}^{\prime}\longrightarrow E_{f(y)}$
is a linear transformation for each $y\in X^{\prime}$. If $\tilde{f}$ is a homeomorphism then it is called an isomorphism of vector bundles. Any bundle isomorphic to a product bundle is called a trivial bundle.
If $X$ is a path-connected space then the local triviality condition insures that fibres of a vector bundle must all be of the same dimension: this is called the rank or dimension of the vector bundle. If $X$ has several path components then a vector bundle over $X$ may well have a different rank on each path component.
Perhaps the most important real vector bundles arise as tangent bundles of compact smooth manifolds $M$. Say that $M$ has dimension $k$. Each point $m\in M$ has a $k$-dimensional real vector space $(TM)_{m}$ associated to it, namely the vectors tangent to the manifold $M$ at the point $m$. These coalesce to form the $k$-dimensional tangent bundle $TM$ with projection map $p:TM\to M$ with $p^{-1}(m)=(TM)_{m}$. We denote the tangent bundle as $TM\to M$ or simply $TM$. Local triviality says that each point $m$ of the manifold has an open set $U$ surrounding it so that if we restrict the bundle to $U$ then, up to isomorphism, the bundle takes the form
$U\times{\mathbb{R}}^{k}\longrightarrow U.$
We may form the dual cotangent bundle $T^{*}M\to M$ similarly, with
$(T^{*}M)_{m}\cong Hom((TM)_{m},{\mathbb{R}})\qquad\forall m\in M.$
A further example. Think of the sphere $S^{k}\subset{\mathbb{R}}^{k+1}$. Then each point $m\in S^{k}$ has vectors based at the point $m$ and perpendicular to the sphere. This forms a one-dimensional bundle $N(S^{k})$ called the normal bundle of $S^{k}$.
Definition 13.3.
Suppose that $f:X\to Y$ and $E\to Y$ is a real vector bundle over $Y$. Then we define the real vector bundle $f^{*}E\to X$ to be the pullback in the commuting diagram
[…]
That is,
$f^{*}(E)=\{(x,e)\in X\times E:f(x)=p(e)\}.$
It is a nice exercise to prove that $f^{*}(E)\to X$ is really a real vector bundle, of the same dimension as $E$, and each $\widetilde{f}:f^{*}(E)_{x}\to E_{f(x)}$ is an isomorphism on fibres. If $X$ is a subset of $Y$ then $f^{*}(E)$ is called the restriction of the bundle to $X$ and sometimes it is written $E|_{X}$.
Definition 13.4.
Suppose that $p:E\to X$ is a real vector bundle of $X$. A cross-section or section of the bundle is a continuous function $s:X\to E$ with the property that $ps:X\to X$ is the identity. It is nowhere zero if $s(x)\neq 0\in E_{x}$ for every $x\in X$. The set of all sections of the real vector bundle $E$ is denoted $\Gamma(E)$.
For example, the trivial bundle $\theta_{n}\to X$ has $n$ linearly independent nowhere zero sections $s_{1},\dots s_{n}$. That is, for each $x\in X$, the set of vectors
$\{s_{1}(x),\dots s_{n}(x)\}$
is linearly independent.
On the other hand, the tangent bundle of $S^{2}$ has the property that every section $s:S^{2}\to T(S^{2})$ must be zero at some point. [^62] If this were not the case then you could comb the hair of a coconut.
Definition 13.5.
Suppose that $E$ and $F$ are real vector bundles over the compact space $X$. We define their sum to be the real vector bundle $E\oplus F$ where the fibre over the point $x$ is $E_{x}\oplus F_{x}$. Similarly, the tensor product of the bundles $E\otimes F$ is the real vector bundle where the fibre over the point $x$ is $E_{x}\otimes_{{\mathbb{R}}}F_{x}$. [^63] This requires more careful explanation but is identical to the complex situation, so please refer to the complex authorities. For example, if $E$ is a real vector bundle over $X$ and $\theta_{1}$ is the one-dimensional trivial bundle, then $E\otimes\theta_{1}\cong E$, generalizing the fact that ${\mathbb{R}}^{n}\otimes_{\mathbb{R}}{\mathbb{R}}\cong{\mathbb{R}}^{n}$.
Finally let $Vect_{k}^{\mathbb{R}}(X)$ denote the set of isomorphism classes of real vector bundles of dimension $k$ and $Vect^{\mathbb{R}}(X)$ denote the union of the $Vect_{k}^{\mathbb{R}}(X)$.
It is easy to see that $Vect^{\mathbb{R}}(X)$ has natural addition and multiplication operations, and these give $Vect^{\mathbb{R}}(X)$ the structure of an abelian semiring with unit $\theta_{1}$. It turns out that we can calculate it as well in homotopy terms.
It is best to restrict at this point to bundles over compact spaces. One very important consequence of this restriction is the following fact: If $E$ is any finite-dimensional real vector bundle over a compact space $X$ then there exists another finite-dimensional real vector bundle $F$ over $X$ such that $E\oplus F$ is a trivial bundle.646464cf. Atiyah [7] p. 27, Cor. 1.4.14.
Over a compact space, isomorphism classes of real vector bundles correspond to homotopy classes of certain maps. Here is the result:
Theorem 13.6.
(cf. Husemoller [44], page 34, Theorem 7.2 ) For each $k$, there is a canonical universal $k$-dimensional real vector bundle $\gamma_{k}:EO_{k}\to BO_{k}$ such that for any compact space $X$ there is an isomorphism of commutative semirings
$X,BO_{k}]^{u}\overset{\cong}{\longrightarrow}Vect_{k}^{\mathbb{R}}(X)$
induced by sending a function $f:X\to BO_{k}$ to $f^{*}\gamma_{k}$.
Remark 13.7.
One should compare this result to Remark 10.11 where the classification of principal $G$-bundles is discussed. The vector bundle $\gamma_{k}$ is strongly related to the universal principal $O_{n}$-bundle.
Definition 13.8.
Suppose that $\pi:E\to X$ is a real vector bundle over the compact space $X$. Then the set of sections $\Gamma(E)$ is a module over the commutative ring $C^{\mathbb{R}}(X)$ with the following operations:
  1. (1)
    Addition defined by
    $(s_{1}+s_{2})(x)=s_{1}(x)+s_{2}(x)$
  2. (2)
    Scalar multiplication defined by
    $(rs)(x)=r(s(x))\qquad r\in{\mathbb{R}},\,\,\,s\in\Gamma(E).$
  3. (3)
    The action of $C^{\mathbb{R}}(X)$ on $\Gamma(E)$ is defined by
    $(f\circ s)(x)=f(x)s(x)\qquad for\,\,\,f\in C^{\mathbb{R}}(X),\quad s\in\Gamma(E).$
It is easy to see that a real bundle $E$ of dimension $n$ has $n$ linearly independent non-vanishing sections if and only if $\Gamma(E)\cong C^{\mathbb{R}}(X)^{n}$ and hence $E$ is isomorphic to the trivial bundle $\Theta_{n}$ of dimension $n$.
As remarked above, every real vector bundle $E\to X$ over a compact space is a direct summand of a trivial bundle:
$E\oplus F\cong\theta_{n}.$
Take sections on both sides and note that
$\Gamma(E)\oplus\Gamma(F)\cong\Gamma(E\oplus F)\cong\Gamma(\theta_{n})\cong\real(X)\oplus\overset{n}{\dots}\oplus\real(X).$
Thus $\Gamma(E)$ is a direct summand of a finitely generated free $\real(X)$-module and hence it is a finitely-generated projective $\real(X)$-module. This implies that the map
$p:\real(X)^{n}\cong\Gamma(E\oplus F)\longrightarrow\Gamma(E)\hookrightarrow\real(X)^{n}$
is an orthogonal projection:
$p=p^{2}=p^{*}\in M_{n}(\real(X)).$
Thus to each real vector bundle $E$ over a compact space $X$ we have associated a projection $p=p(E)\in M_{n}(\real(X))$. We denote this map as
$\mathcal{P}:Vect^{\mathbb{R}}(X)\longrightarrow\mathcal{P}_{\infty}(\real(X))$
from the category $Vect^{\mathbb{R}}(X)$ of isomorphism classes of vector bundles over $X$ to the category $\mathcal{P}_{\infty}(\real(X))$ of projections in matrix rings over $C(X)$.
Remark 13.9.
We note in passing all of the analogous properties of complex vector bundles, all established in the complex authorities.
  1. (1)
    Locally, a complex bundle has the structure $U\times{\mathbb{C}}^{k}\longrightarrow U$.
  2. (2)
    A complex bundle of complex dimension $n$ has an associated $U_{n}$-bundle
    $E\times_{U_{n}}{\mathbb{C}}^{n}\to B.$
  3. (3)
    The space $\Gamma(E)$ of complex sections of a complex bundle $E\to B$ over a compact space $X$ is a module over the ring $C^{\mathbb{C}}(X)$.
  4. (4)
    $X,BU_{k}]^{u}\cong Vect_{k}^{\mathbb{C}}(X).$
Here’s a similar construction that we will need in the next section.
Definition 13.10.
Suppose that $p:E\to X$ is a real vector bundle of real dimension $n$. Then we may form the associated complexified vector bundle
$E\otimes{\mathbb{C}}=E\times_{{\mathbb{R}}^{n}}{\mathbb{C}}^{n}$
of complex dimension $n$. On each fibre $E_{x}\cong{\mathbb{R}}^{n}$ we are simply forming
$E_{x}\otimes_{{\mathbb{R}}^{n}}{\mathbb{C}}\cong{\mathbb{C}}^{n}.$
This construction induces the complexification map
$c:Vect_{\mathbb{R}}^{n}(X)\longrightarrow Vect_{\mathbb{C}}^{n}(X)$
which is a map of semirings.
Principal bundles may be pulled back exactly as is done with vector bundles, and in fact, for $X$ compact, isomorphism classes of principal $G$-bundles correspond exactly to elements of $[X,BG]$.
14. The Basics: $KO$-theory for Spaces
Definition 14.1.
. For a compact space $X$, define
$KO^{0}(X)=Groth(Vect^{\mathbb{R}}(X))$
the Grothendieck group of $Vect^{\mathbb{R}}(X)$. Elements of $KO^{0}(X)$ may be represented as formal differences of vector bundles [E][F][E]-[F] over $X$. Then $KO^{0}(X)$ is a commutative ring with identity $1=[\theta_{1}]$, the class of the one-dimensional trivial bundle over $X$.
If $f:X\to Y$ then there is an induced map
$f^{*}:KO^{0}(Y)\longrightarrow KO^{0}(X)$
defined on generators by
$f^{*}[E]=[f^{*}(E)]$
and then $KO^{0}$ is a contravariant [^65] A functor $F$ is contravariant if a map $f:A\to B$ induces a map $f^{*}:FB\to FA$ with the usual naturality conditions. functor from compact spaces to commutative unital rings.
Here are a few examples.
  • If $X$ is contractible, such as the closed unit ball, then $X$ has only trivial bundles, and hence $KO^{0}(X)\cong{\mathbb{Z}}$. In particular, if $X=\{x_{o}\}$ is a single point then $KO^{0}(x_{o})\cong{\mathbb{Z}}$ as commutative unital rings. Thus if $x_{o}$ is the basepoint of $X$ then there is an associated “dimension” map
    $KO^{0}(X)\overset{dim}{\longrightarrow}KO^{0}(pt)\cong{\mathbb{Z}}$
  • If $X$ is the circle $S^{1}$ then the tangent bundle is trivial, as there is an obvious nowhere zero section. However, it turns out that there is a non-trivial real bundle $M$ of dimension $1$. It is essentially the Mobius strip! Note that if you go around the Mobius strip twice then you get a trivial bundle $\theta_{2}=2\theta_{1}$. So $KO^{0}(S^{1})$ is generated by $\theta_{1}]$ and by $[M]$ with $2([M]-\theta_{1})=0$. Then we have the split exact sequence
    $0\longrightarrow{\mathbb{Z}}_{2}\longrightarrow KO^{0}(S^{1})\overset{dim}{\longrightarrow}{\mathbb{Z}}\longrightarrow 0$
    and hence $KO^{0}(S^{1})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ with the summands being generated by $\theta_{1}$ and $M]-\theta_{1}$ respectively and
    $dim([M]-\theta_{1})=0.$
  • (Kaplansky) Suppose that $X=S^{n}$ some sphere. Then we have the tangent bundle $TS^{n}$ and also the normal bundle $N(S^{n})\cong\theta_{1}$, consisting of all vectors at a point of the sphere that are perpendicular to the tangent plane of the sphere. It is easy to see that
    $TS^{n}\oplus N(S^{n})\cong\theta_{n+1}$
    so, in particular, $TS^{n}$ plus a trivial bundle is isomorphic to a trivial bundle, whether $TS^{n}$ is trivial or not. Such a bundle is called stably trivial. So the tangent bundle of $S^{n}$ is stably trivial for all $n$. It is a rather deep result that $TS^{n}$ is isomorphic to a trivial bundle if and only if n=1, 3,n=1,\,3, or $7$. (This is related to the existence of the complex numbers, the quaternions, and the Cayley numbers in dimensions 2, 4,2,\,4, and $8$ and the fact that there are no other such algebras in other dimensions.)
We are particularly interested in relating $KO$ for spaces to $KO$ for $C^{*}$-algebras. Here is the main result.
Theorem 14.2.
(Swan [83]) Let $X$ be a compact space. Then there is a natural equivalence of categories
$Vect^{\mathbb{R}}(X)\simeq\it{Proj^{R}}(C^{\mathbb{R}}(X))$
between the category $Vect^{\mathbb{R}}(X)$ of isomorphism classes of vector bundles over $X$ and the category $\it{Proj^{R}}(C^{\mathbb{R}}(X))$ of projections in matrix algebras over $C^{\mathbb{R}}(X)$. This induces a natural isomorphism of functors
$KO^{0}(X)\cong KO_{0}(C(X)).$
We will put off a discussion of the proof of Swan’s Theorem and related results until the next section of this article.
Here’s an immediate consequence.
Definition 14.3.
Let $X$ be a compact space. Then the groups $KO^{-n}(X)$ are defined by
$KO^{-n}(X)=KO_{n}(\real(X)).$
If $f:X\to Y$ then this definition yields a map
$f^{*}:KO^{-n}(Y)\longrightarrow KO^{-n}(X)$
and each $KO^{-n}$ is a contravariant, homotopy-invariant functor from compact spaces to abelian groups.
Definition 14.4.
  1.
  2. (1)
    For a locally compact space $X$ with one-point compactification $X^{+}$, define
    $KO^{0}(X)=ker\big[KO^{0}(X^{+})\longrightarrow KO^{0}(+)\cong{\mathbb{Z}}\big]$
    so there is a natural short exact sequence
    $0\longrightarrow KO^{0}(X)\longrightarrow KO^{0}(X^{+})\longrightarrow{\mathbb{Z}}\longrightarrow 0$
    and an (unnatural) isomorphism
    $KO^{0}(X^{+})\cong KO^{0}(X)\oplus{\mathbb{Z}}$
    A proper [^66] Proper maps of locally compact spaces are those that extend uniquely to one-point compactifications map $X\to Y$ of locally compact spaces induces a homomorphism $KO^{0}(Y)\to KO^{0}(X)$ and $KO^{0}$ is a contravariant functor from the category of locally compact spaces and proper maps and homomorphisms to abelian groups.
  3. (2)
    For $n>0$ and a locally compact space $X$, we define
    $KO^{-n}(X)=KO_{n}(C_{o}^{\mathbb{R}}(X))$
  4. (3)
    As $KO_{n}(A)$ is periodic of period 8 by Bott periodicity for all real $C^{*}$-algebras, we conclude that the same holds for the special case of $KO^{*}(X)$ and hence we may extend $KO^{*}(X)$ to a $\mathbb{Z}$}-graded collection of contravariant functors which satisfy
    $KO^{n}(X)\cong KO^{n-8}(X)\qquad\forall n.$
Definition 14.5.
For $X$ a compact space with basepoint $x_{o}$, we define the real reduced $KO$-theory groups by
$\widetilde{KO}^{n}(X)=ker\big[KO^{n}(X)\longrightarrow KO^{n}(x_{o})\big].$
Theorem 14.6.
Suppose that $X$ is a based compact space. Then
  1. (1)
    $\widetilde{KO}^{0}(X)\cong[X,\mathbf{BO}\times{\mathbb{Z}}]\cong[X,{\mathcal{R}}_{0}].$
  2. (2)
    For $0\leq n\leq 7$,
    $\widetilde{KO}^{-n}(X)\cong\widetilde{KO}^{0}(S^{n}X)\cong[S^{n}X,\mathbf{BO}\times{\mathbb{Z}}]\cong$
    $\cong[S^{n}X,{\mathcal{R}}_{0}]\cong[X,\Omega^{n}{\mathcal{R}}_{0}]\cong[X,{\mathcal{R}}_{n}].$
Here is a comprehensive statement of the main properties of the reduced cohomology theory $\widetilde{KO}^{*}(X)$.
Theorem 14.7.
([1], pages 66-76) $\widetilde{KO}^{*}(X)$ is a reduced cohomology theory. [^67] We are looking at theories on compact spaces, as that is the category most closely tied to functional analysis. For example, we want to include the spectrum of a bounded operator, and that means that we must include every closed subset of the plane. Most algebraic topology textbooks use categories closely tied to CW-complexes. In that setting one expects different behavior from cohomology theories. For example, we require a long exact sequence for every $Y\subset X$ where $Y$ is a closed subset of $X$. In most algebraic topology settings the requirement would be that $Y$ sits nicely inside of $X$; you can see the difference by letting $Y$ be a Cantor set inside the unit interval $X=[0,1]$. This would not be acceptable in most algebraic topology contexts. That is, for each $j\geq 0$,
  1. (1)
    $\widetilde{KO}^{-j}(-)$ is a contravariant functor from based compact spaces to abelian groups.
  2. (2)
[…]
  3. (3)
    If $f\simeq g$ are based homotopic maps $X\to Y$ then
    $f^{*}=g^{*}:\widetilde{KO}^{-j}(Y)\to\widetilde{KO}^{-j}(X).$
  4. (4)
    There is a suspension isomorphism
    $\widetilde{KO}^{-j}(SX)\cong\widetilde{KO}^{-j-1}(X).$
  5. (5)
    If $\iota:Y\to X$ is an inclusion of a closed subset of $X$ with quotient space $X/Y$ and associated sequence
    $Y\overset{\iota}{\longrightarrow}X\overset{\pi}{\longrightarrow}X/Y$
    then there is a long exact sequence
    $\dots\longrightarrow\widetilde{KO}^{-j-1}(X)\overset{\iota^{*}}{\longrightarrow}\widetilde{KO}^{-j-1}(Y)\longrightarrow\widetilde{KO}^{-j}(X/Y)\overset{\pi^{*}}{\longrightarrow}\widetilde{KO}^{-j}(X)\overset{\iota^{*}}{\longrightarrow}\dots$
  6. (6)
    For $X$ and $Y$ based compact spaces, there is a natural short exact sequence
    $0\longrightarrow\widetilde{KO}^{-j}(X\wedge Y)\longrightarrow\widetilde{KO}^{-j}(X\times Y)\longrightarrow\widetilde{KO}^{-j}(X)\oplus\widetilde{KO}^{-j}(Y)\longrightarrow 0$
    which splits unnaturally.
  7. (7)
[…]
    and similarly for $\widetilde{KO^{*}}(X)$.
  8. (8)
    Bott Periodicity: There is a natural isomorphism
    $\widetilde{KO}^{*}(X)\cong\widetilde{KO}^{*-8}(X).$
  9. (9)
    Künneth Theorem (special case) [15]. If $X$ and $Y$ are compact spaces and $Y$ has the homotopy type of a sphere or a product of spheres, then the internal product induces a natural isomorphism
    $\widetilde{KO}^{*}(X)\otimes_{KO^{*}(pt)}\widetilde{KO}^{*}(Y)\overset{\cong}{\longrightarrow}\widetilde{KO}^{*}(X\wedge Y).$
Remark 14.8.
If $X$ is an infinite CW-complex such as $RP^{\infty}$, $BO_{n}$, or $\mathbf{BO}$, then there are several ways to generalize $KO$-theory to deal with them. However, we generally must give up the idea of describing elements of the resulting group via real vector bundles.
Remark 14.9.
$KO^{*}$ is thus a cohomology theory in the classical Eilenberg-Steenrod sense, and so it has a spectrum [^68] The word “spectrum” has many diverse meanings within science. Here it is important to note that this use of the word has nothing at all to do with the spectrum of an operator on Hilbert space. The word spectrum here means a sequence of spaces with related properties. For example, it might be the sequence $\{\Omega^{n}X\}$ for some space $X$ but with no assumption of periodicity. and an associated homology theory (cf. May [59]). Unfortunately, this homology theory does not have a concrete, simple characterization and one cannot, for instance, easily write down elements of the $KO$-homology of nice spaces. The Kasparov framework, using
$KO_{*}(X)\cong KK_{*}^{\mathbb{R}}(\real(X),{\mathbb{R}})$
doesn’t seem easy to use either. Perhaps it will be of some use in the future to physicists, but that is out of our vision for now.
Both complex and real $K$-theory for spaces are related to the cohomology of the space. This connection is useful but complicated. The following is a minimal sketch of the idea, first in the complex case and then in the real case.
Each complex vector bundle $E\to X$ has associated to it certain cohomology classes
$c_{i}(E)\in H^{2i}(X;{\mathbb{Z}})$
called Chern classes (cf. Milnor-Stasheff [63], Chapter 14). These may be defined using differential geometry or purely topologically. Here’s a quick sketch. Suppose $E\to X$ is a complex bundle of dimension $n$. We use the fact that the integral cohomology of the space $BU_{n}$ is a polynomial algebra over the integers with generators $\{c_{1},\dots,c_{n}\}$ with $dim(c_{j})=2j$:
$H^{*}(BU_{n};{\mathbb{Z}})\cong{\mathbb{Z}}[c_{1},\dots,c_{n}].$
Assuming $X$ is compact, then there is a continuous map $f:X\to BU_{n}$ such that $E\cong f^{*}(\gamma_{n})$ is isomorphic to the pullback of the universal bundle $\gamma_{n}$ over $BU_{n}$ and $f$ is unique up to based homotopy. Define the Chern class $c_{j}(E)$ of the vector bundle $E$ by
$c_{j}(E)=f^{*}(c_{j})\in H^{2j}(X;{\mathbb{Z}}).$
Next, there is a certain rational combination of the classes $c_{j}(E)$ called the Chern character and written
$ch(E)\in\bigoplus_{k}H^{2k}(X;{\mathbb{Q}}).$
Here are its principal properties on the category of compact spaces.
Proposition 14.10.
(cf. Hatcher [40], Section 4.1.)
  1. (1)
    The Chern character is a natural transformation
    $ch:K^{0}(X)\longrightarrow\bigoplus_{k}H^{2k}(X;{\mathbb{Q}})$
    and using suspension this may be extended to
    $ch:K^{-1}(X)\longrightarrow\bigoplus_{k}H^{2k-1}(X;{\mathbb{Q}}).$
  2. (2)
    The Chern character is a homomorphism of graded commutative rings. Specifically,
    $ch(E\oplus F)=ch(E)+ch(F)$
    and
    $ch(E\otimes F)=ch(E)ch(F).$
  3. (3)
    The Chern character induces an isomorphism of graded commutative rings
    $ch:K^{*}(X)\otimes{\mathbb{Q}}\overset{\cong}{\longrightarrow}\bigoplus_{k}H^{**}(X;{\mathbb{Q}}).$
It is very important to note that in general the Chern character takes values generally in rational cohomology $\oplus_{k}H^{2k}(X;{\mathbb{Q}})$. Only rarely does it take values in integer cohomology $\oplus_{k}H^{2k}(X;{\mathbb{Z}})$.
Here’s a good example. Take $X=S^{2n}$, the sphere of even dimension. Then its cohomology vanishes except in dimensions 0 and $2n$, so the only Chern class of a complex bundle $E\to S^{2n}$ that has a chance of being non-zero is $c_{n}(E)$. Here is what happens.
For the even spheres, the Chern character simplifies considerably. If $E$ is any complex bundle over $S^{2n}$, then
$ch(E)=dim(E)+\frac{(-1)^{n+1}c_{n}(E)}{(n-1)!}.$
Using this information, we have the following result.
Proposition 14.11.
(cf. Hatcher [40], 4.4) A class in $H^{2n}(S^{2n};{\mathbb{Z}})\cong{\mathbb{Z}}$ occurs as a Chern class $c_{n}(E)$ if and only if it is divisible by (n1)!(n-1)!.
Moving on, we come to the case of $KO^{*}(X)$, and for this we refer to Milnor-Stasheff [63] Chapter 15 and to Hatcher [40]. If $E\to X$ is a real vector bundle with complexification $E\otimes{\mathbb{C}}$. It turns out that
$2c_{2j+1}(E\otimes{\mathbb{C}})=0\quad\forall j$
so the only elements of the form $c_{k}(E\otimes{\mathbb{C}})$ that may be nonzero after tensoring with $\mathbb{Q}$} have $k$ even, so that the dimension of the element is some multiple of $4$.
Define the $i$’th Pontrjagin class $p_{i}(E)$ of the real bundle $E\to X$ by
$p_{i}(E)=(-1)^{i}c_{2i}(E\otimes{\mathbb{C}})\in H^{4i}(X;{\mathbb{Z}}).$
Define the Pontrjagin character $Ph$ to be the composite
$Ph:KO^{0}(X)\overset{c}{\longrightarrow}K^{0}(X)\overset{ch}{\longrightarrow}\bigoplus_{k}H^{4k}(X;{\mathbb{Q}}).$
Here are its principal properties on the category of compact spaces. For much more, see Grady and Sati [38] Section 2.1.
Proposition 14.12.
(cf. Hatcher [40], Section 4.1.)
  1. (1)
    The Pontrjagin character is a natural transformation
    $Ph:KO^{0}(X)\longrightarrow\bigoplus_{k}H^{4k}(X;{\mathbb{Q}})$
    and using suspension this may be extended to
    $Ph:KO^{*}(X)\longrightarrow\bigoplus_{k}H^{**}(X;{\mathbb{Q}}).$
  2. (2)
    The Pontrjagin character is a homomorphism of graded commutative rings. Specifically,
    $Ph(E\oplus F)=Ph(E)+Ph(F)$
    and
    $Ph(E\otimes F)=Ph(E)Ph(F).$
  3. (3)
    The Pontrjagin character induces an isomorphism of graded commutative rings
    $Ph:K^{0}(X)\otimes{\mathbb{Q}}\overset{\cong}{\longrightarrow}\bigoplus_{k}H^{4k}(X;{\mathbb{Q}}).$
Remark 14.13.
(Grady and Sati [38], §2.1.) It’s interesting to look at the simplest case, when $X$ is a point $pt$. Then $KO^{-*}(pt)=KO_{*}({\mathbb{R}})$ and so $KO^{*}(pt)\otimes{\mathbb{Q}}={\mathbb{Q}}$ when =0,4*=0,4 and is 0 otherwise. The generator of $KO^{0}(pt)$ is the identity of the ring, and is represented by a one-dimensional trivial bundle over the point. It goes to the generator of $H^{0}(pt;{\mathbb{Z}})\cong{\mathbb{Z}}\subset H^{0}(pt;{\mathbb{Q}})$ since $Ph$ is a ring map. The group $KO^{-4}(pt)$ is more interesting. We have a natural isomorphism
$KO^{-4}(pt)\cong KO^{0}(S^{4})$
and then
${\mathbb{Z}}\cong\pi_{4}(S^{4})\overset{\cong}{\longrightarrow}KO^{0}(S^{4})\overset{Ph}{\longrightarrow}H^{4}(S^{4};{\mathbb{Q}})$
and the image of the identity is twice the generator of $H^{4}(S^{4};{\mathbb{Z}})$, as may be seen by pairing it with the fundamental class of $S^{4}$. The upshot of this is that the map
$KO^{*}(pt)\otimes{\mathbb{Q}}\overset{Ph}{\longrightarrow}\bigoplus_{k}H^{4k}(pt;{\mathbb{Q}})$
actually is integral, taking the form [^69] One may reasonably ask why don’t we see $H^{-4}(pt;{\mathbb{Z}})$ rather than $H^{4}(pt;{\mathbb{Z}})$. The answer is that we should! The problem is that ordinary homology and cohomology were invented first- H. Poincaré’s 1895 paper Analysis Situs is a reasonable starting point, but he used Betti numbers rather than homology groups per se, and then in the 1920’s and 1930’s several people started to use our present notation. There is no reason to think that they were troubled by pairing two classes of degree $n$ and winding up with a class of degree 0.
$KO^{0}(pt)\overset{\cong}{\longrightarrow}H^{0}(pt;{\mathbb{Z}})\cong{\mathbb{Z}}\qquad\text{and}\qquad KO^{-4}(pt)\overset{\times 2}{\longrightarrow}H^{4}(pt;{\mathbb{Z}})\cong{\mathbb{Z}}.$
Finally, we touch on Chern numbers. We do this gingerly, since the name Chern number is used in certain areas of physics for pretty much any number that can be calculated via some sort of topological construct. Here is the classical definition.
Definition 14.14.
Suppose that $M^{n}$ is an oriented connected manifold and the tangent bundle $TM$ can be given the structure of a complex vector bundle with associated Chern classes $c_{j}(TM)\in H^{2j}(M^{n};{\mathbb{Z}})$. Choose some polynomial
$p=p(c_{1}(TM),\dots c_{n}(TM))$
where each term has total degree nn. [^70] For example, we could have $p=17c_{1}(TM)^{4}+37c_{2}(TM)^{2}$ as each term has total degree $8$. Let $M]\in H_{n}(M;{\mathbb{Z}})$ denote the fundamental class of the manifold. Then evaluating $p\in H^{n}(M;{\mathbb{Z}})$ against the fundamental class $M]\in H_{n}(M;{\mathbb{Z}})$ gives a Chern number
$p,[M]>\,\in\,{\mathbb{Z}}.$
More generally, we can replace the tangent bundle by any complex bundle over $M$ and define Chern numbers associated to that new bundle.
15. Relating $K$-theory for $C^{*}$-Algebras and $K$-theory for Spaces
In this section, we prove theorems that connect the topological $KO$-theory of a compact or locally compact space $X$ to the $C^{*}$-algebraic $KO$-theory of the $C^{*}$-algebra $A=\real(X)$ or $A=C_{o}^{{\mathbb{R}}}(X)$ respectively, and similarly in the complex setting. Some of these theorems are well-known among operator algebraists, but references are hard to find.
Theorem 15.1.
Let $X$ be a compact topological space. Then:
  1. (1)
    There is a natural isomorphism
    $KO_{0}(\real(X))\cong KO^{0}(X).$
    Similarly, if $X$ is a locally compact topological space, then there is a natural isomorphism
    $KO_{0}(C_{o}^{{\mathbb{R}}}(X))\cong KO^{0}(X).$
  2. (2)
    There is a natural isomorphism
    $K_{0}(C^{{\mathbb{C}}}(X))\cong K^{0}(X).$
    Similarly, if $X$ is a locally compact topological space, then there is a natural isomorphism
    $K_{0}(C_{o}^{{\mathbb{C}}}(X))\cong K^{0}(X).$
This theorem is an immediate consequence of a theorem of R.G. Swan [^71] CS: Swan was on faculty at the U. of Chicago where I studied. He was very shy - would time his lectures so that he was standing by the exit door at the end. One day in my homological algebra class the instructor, Saunders MacLane, got stuck on a proof. The class ended and he told us he would finish the proof next time. At the next class MacLane said that, being still stuck, he went for help to the person that everyone in the Department went to when they needed help - to Dick Swan. Swan gave him the answer. that we mentioned previously:
Theorem 15.2.
(R. G. Swan [83]) Let $X$ be a compact space and let $\mathbb{F}$} denote either the real or the complex numbers. [^72] Swan also includes the quaternions. Then there is a natural isomorphism
$\Theta:\mathcal{P}_{\infty}(C^{\mathbb{F}}(X))\simeq Vect^{\mathbb{F}}(X)$
between $Vect^{\mathbb{F}}(X)$, isomorphism classes of $\mathbb{F}$} - vector bundles over $X$, and $\mathcal{P}_{\infty}(C^{\mathbb{F}}(X))$, projections in matrix rings over $C^{\mathbb{F}}(X)$.
Swan’s introduction is worth quoting:
Serre …has shown that there is a one-to-one correspondence between algebraic vector bundles over an affine variety and finitely generated projective modules over its coordinate ring. For some time, it has been assumed that a similar correspondence exists between topological bundles over a compact Hausdorff space $X$ and finitely generated projective modules over the ring of continuous real-valued functions on XX… However, no rigorous treatment of the correspondence seems to have been given. I will give such a treatment here…
Theorem 15.2 implies Theorem 15.1 by simply applying the Grothendieck construction. If $\mathbb{F}$$\mathbb{R}$} then
$\rm{Groth}(\mathcal{P}_{\infty}(A))=KO_{0}(A).$
and
$\rm{Groth}(\rm{Vect}^{\mathbb{R}}(X))=KO^{0}(X).$
Then use the fact that the Grothendieck construction is natural with respect to homomorphisms of abelian semigroups.
There are several nice proofs of the complex case of these theorems, but proofs of the real case are scarce. We shall supply a new proof of the real case of Theorem 15.2 which is somewhat more analytic than Swan’s proof.
We first focus on the case that $X$ is compact. For any projection
$p\in\mathcal{P}_{n}(A)\subseteq C(X,M_{n}({\mathbb{R}})),$
let
$E_{p}={\rm range~}p$ $=\{(x,v)\in X\times{\mathbb{R}}^{n}\mid v\in p(x){\mathbb{R}}^{n}\}$
$=\{(x,v)\in X\times{\mathbb{R}}^{n}\mid p(x)v=v\}\;.$
Then $E_{p}$ is a real vector bundle over $X$ with associated projection map $\pi\colon E_{p}\rightarrow X$ given by $\pi(x,v)=x$. It is clear that $E_{p}\cong E_{p\oplus 0}$ where
$p\oplus 0=\mathrm{diag}(p,0)\in\mathcal{P}_{n+1}(A).$
Lemma 15.3.
Let $X$ be a compact space and let $A=\real(X)$. If $p,q\in\mathcal{P}_{n}(A)$ are unitarily equivalent, then there is an isomorphism of real bundles $E_{p}\cong E_{q}$.
Proof.
Let
$E_{p}$ $=\{(x,v)\in X\times{\mathbb{R}}^{n}\mid p(x)v=v\}$
$E_{q}$ $=\{(x,v)\in X\times{\mathbb{R}}^{n}\mid q(x)v=v\}$
and let $u\in C(X,M_{n}({\mathbb{R}}))$ be a unitary satisfying $upu^{*}=q$. There is an homomorphism $\phi:E_{p}\to E_{q}$ defined by $\phi(x,v)=(x,u(x)v)$. Indeed, we check that if $p(x)v=v$ then
$q(x)\cdot u(x)v=u(x)p(x)v=u(x)v.$
The inverse is defined using $u^{*}(x)$. Furthermore, $\phi$ clearly commutes with the projection map $\pi$. Thus $E_{p}$ and $E_{q}$ are isomorphic objects in the category of real vector bundles. ∎
It follows that the construction $p\mapsto E_{p}$ gives a well-defined map
$\Theta\colon\mathcal{P}_{\infty}(A)\rightarrow\rm{Vect}^{\mathbb{R}}(X)$
and it is easy to show that this is a semigroup homomorphism, that is
$\Theta(p\oplus q)\cong\Theta(p)\oplus\Theta(q).$
Lemma 15.4.
The map $\Theta$ is an isomorphism.
Proof.
To prove that $\Theta$ is injective, assume that $p,q$ in $M_{n}(A)$ are projections that satisfy $E_{p}\cong E_{q}$. So there exists $a\in M_{n}(A)=C(X,M_{n}({\mathbb{R}}))$ where $a_{x}$ is an isomorphism from $p(x){\mathbb{R}}^{n}$ to $q(x){\mathbb{R}}^{n}$ for each $x$ (and $a_{x}$ vanishes on $p^{\perp}$). Thus we have $ap=a$ and $qa=a$. Using continuous functional calculus, let $s=f(a^{*}a)a^{*}$, where
$f(t)=\begin{cases}0&t=0\\ t^{-1/2}&t>0.\end{cases}$
Since $X$ is compact, the non-zero part of the spectrum of $a$ is bounded below. So $f$ is continuous on $\rm{sp}(a)$ and $s\in M_{n}(A)$. Then it can be checked that $s$ is a partial isometry that satisfies $ss^{*}=q$ and $s^{*}s=p$.
Now let
$u=\begin{bmatrix}s&1-q\\ 1-p&s^{*}\end{bmatrix}\;$
in $M_{2n}(A)=C(X,M_{2n}({\mathbb{R}}))$. Check that $u$ is a unitary and that it satisfies
$u\begin{bmatrix}p&0\\ 0&0\end{bmatrix}u^{*}=\begin{bmatrix}q&0\\ 0&0\end{bmatrix}\;$.
Thus $\mathrm{diag}(p,0)$ is unitarily equivalent to $\mathrm{diag}(q,0)$ in $M_{2n}(A)$, so $[p]=[q]$ in $\rm{Proj}(A)$.
To show that $\Theta$ is surjective, let $E$ be any locally trivial vector bundle with projection map $\pi\colon E\rightarrow X$. Since $E$ is locally trivial, there is a finite open cover $\{U_{i}\}_{i=1}^{k}$ of $X$ such that for each $i$ the restriction $E_{U_{i}}$ is isomorphic to $U_{i}\times{\mathbb{R}}^{n}$. Let $\phi_{i}\colon E_{U_{i}}\rightarrow U_{i}\times{\mathbb{R}}^{n}$ be this isomorphism. Now, let $\{\psi_{i}\}_{i=1}^{k}$ be a partition of unity subordinate to $\{U_{i}\}$, a collection of continuous functions from $X$ to $[0,1]$ such that $\sum_{i=1}^{k}\psi_{i}=1$ and $\rm{supp}(\psi_{i})\subseteq U_{i}$. Then the function
$f(v)=(\pi(v),\psi_{1}(v)\phi_{1}(v),\dots,\psi_{k}(v)\phi_{k}(v))$
defines an injective map of vector bundles from $E$ to $X\times{\mathbb{R}}^{nk}$. Using the standard inner product on ${\mathbb{R}}^{nk}$, for each $x\in X$, let $p_{x}$ be the orthogonal projection from ${\mathbb{R}}^{nk}$ to the range of $f$ over $x$. Then $p_{x}$ is a continuous function of $x$ and is a projection in the algebraic sense that $p^{2}=p$ and $p^{*}=p$. Thus $p]\in\mathcal{P}_{\infty}(A)$ and by construction we have $\Theta([p])\cong E$. ∎
Proof of Theorem 15.2..
When $X$ is compact and $\mathbb{F}$$\mathbb{R}$}, this is Lemma 15.4. Now, suppose that $X$ is locally compact. Let $X^{+}=X\cup\{+\}$ be the one-point compactification. As in Definition 11.3 we have
$KO^{0}(X)=\ker\left[KO^{0}(X^{+})\xrightarrow{\iota^{*}}KO^{0}(+)\right]\;$ (1)
where the map is induced by the inclusion map $\iota\colon\{+\}\hookrightarrow X^{+}$. On the other hand, if we set $A=C_{o}^{{\mathbb{R}}}(X)$, then $A$ is a non-unital real $C^{*}$-algebra and from Definition 6.5 we have
$KO_{0}(A)=\ker\left[KO_{0}(A^{+})\xrightarrow{\pi_{*}}KO_{0}({\mathbb{R}})\right]\;$ (2)
where the map is induced by the canonical map $\pi\colon A^{+}\rightarrow A^{+}/A\cong{\mathbb{R}}$ associated with a unitization.
In this situation, we have an isomorphism $A^{+}\cong\real(X^{+})$ and under this isomorphism the map $\pi$ corresponds to the evaluation map
$\rm{ev}\colon\real(X^{+})\rightarrow{\mathbb{R}}\;,$
specifically evaluation at the point ++. This evaluation map is dual to the inclusion map $\iota$. This shows that the maps shown in equations (1) and (2) are identical and hence there is an isomorphism $KO^{0}(X)\cong KO_{0}(A)$.
Finally, the proof in the complex case (for $X$ compact and locally compact) is exactly the same as the proof above, including the proofs of Lemmas 15.3 and 15.4, replacing projections in $M_{n}(\real(X))$ with projections in $M_{n}(C^{{\mathbb{C}}}(X))$, replacing real vector bundles with complex vector bundles, and replacing the semigroup $\rm{Vect}^{\mathbb{R}}(X)$ with $\rm{Vect}^{\mathbb{C}}(X)$. ∎
Corollary 15.5.
Suppose that $X$ is a connected compact space. If $i>0$, then
  1. (1)
    $KO^{-i}(X)\cong KO^{0}({\bf{R}}^{i}\times X)\;$
    understood as a subgroup of $K$-theory of the compactification
    $KO^{0}(({\bf{R}}^{i}\times X)^{+}).$
  2. (2)
    In particular, $KO^{0}({\bf{R}}^{i}\times X)$ is the subgroup of $KO^{0}(({\bf{R}}^{i}\times X)^{+})$ generated by formal differences of vector bundles [E][F][E]-[F] over $({\bf{R}}^{i}\times X)^{+}$ where $E$ and $F$ have the same dimension.
  3. (3)
    $KO_{0}(C_{o}^{{\mathbb{R}}}({\bf{R}}^{i}\times X))$ is the subgroup of $KO_{0}(\real(({\bf{R}}^{i}\times X)^{+}))$ generated by formal differences of projection classes [p][q][p]-[q] where $\mathrm{ev}_{*}[p]\cong\mathrm{ev}_{*}[q]$ where $\mathrm{ev}\colon\real(({\bf{R}}^{i}\times X)^{+})\rightarrow{\mathbb{R}}$ is point evaluation.
We can deal with negative degrees by writing (again for $i>0$),
  1. (1)
[…]
    where $\hat{\tau}$ is the involution on ${\bf{R}}^{i}\times X$ given by $(t_{1},\dots,t_{i},x)\mapsto(-t_{1},\dots,-t_{i},x)$.
  2. (2)
    In a similar way to the previous paragraph, this $K$-theory group can be expressed as a subgroup of $KR^{0}(({\bf{R}}^{i}\times X,\tau)^{+})$ consisting of differences $(E,\sigma_{1})]-[(F,\sigma_{2})]$ where $E$ and $F$ have the same dimension; or as the kernel of $\mathrm{ev}_{*}$ in $KO_{0}(A)$ where
    $A=\{f\in C^{{\mathbb{C}}}(({\bf{R}}^{i}\times X)^{+})\mid f(t_{1},\dots,f_{i},x)=\overline{f(-t_{1},\dots,-t_{i},x)}\}\;.$
Note that in (2) above, $KR^{0}(~)$” refers to Atiyah’s “Real $K$-theory” which is described in the next section.
Proof.
The first statement follows from Swan’s Theorem:
$KO^{-i}(X)=KO_{i}(\real(X))=KO_{0}(S^{i}\real(X))=KO_{0}(C_{o}^{{\mathbb{R}}}({\bf{R}}^{i}\times X))=KO^{0}({\bf{R}}^{i}\times X).\;$
In the negative degree setting we have, again from Swan’s Theorem,
[…]
and the other results follow. ∎
Remark 15.6.
We have discussed topological $K$-theory for real and complex vector bundles. There are many other families of compact Lie groups, such as $Spin(n)$, $Spin^{c}(n)$, $Sp(n)$, etc. with natural actions on ${\mathbb{R}}^{n}$, ${\mathbb{C}}^{n}$, or ${\mathbb{H}}^{n}$. Let $G$ be one of them, and assume it acts on ${\mathbb{R}}^{n}$. Then we may form $G$-vector-bundles by starting with principal $G$-bundles $P\to X$ (with $X$ compact for safety) and forming $V=P\times_{G}{\mathbb{R}}^{n}$. The set of isomorphism classes of such $G$-bundles is closed under direct sum, and so we may form a Grothendieck group to obtain $KG^{0}(X)$. Some of these groups turn out to be very useful indeed, but they take us too far away from our main object of interest. To get a taste of what is possible, check out Stong [82] Chapter XI, which has the great title SpinSpin, $Spin^{c}$, and Similar Nonsense.” Stong is writing about cobordism, but the two subjects are strongly related in some cases. [^73] Stong was an unusual person, to put it mildly. He would get to his office around 3 AM, work diligently until 8 or 9 AM, and then lounge in the coffee room for the rest of the day, so it looked like he was doing nothing. He is well-known, among other things, for the Hattori-Stong theorem. J.F. Adams once commented that it should have been called the Stong-Hattori theorem, since Stong proved it first, but people called it Hattori-Stong because they liked the emphasis implied by saying “Stong” second.
Remark 15.7.
Suppose that $G$ is a topological group (compact for simplicity). Then we may focus attention on a compact space $X$ for which there is a natural action $G\times X\to X$. Call $X$ a $G$-space. Suppose then given (real or complex) vector bundles $V\to X$ where $G$ acts on $X$ and the projection map commutes with the $G$-action. Then we get $G$-vector bundles. Taking Grothendieck groups again we obtain equivariant $K$-theory $K^{G}(X)$. [^74] We can do this more generally for operator algebras but this is abstract enough for a remark. These groups turn out to be computable (sometimes) and quite interesting (always). The cases $G$ finite and $G$ compact connected turn out to be quite different, and we refer the interested reader to N. C. Phillips [68] and to Rosenberg-Schochet [71] for examples of their use.
16. Atiyah’s $KR$- theory
Now we consider $K$-theory in the category of topological spaces with involution, a cohomology theory invented by M. Atiyah [6] and give a new proof that it can be realized as $KO$ of an appropriate real $C^{*}$-algebra.
Atiyah described the birth of $KR$-theory as follows:
In attempting to understand reality questions of elliptic operators Singer and I were, for a long time, held up by the fact that real $K$-theory behaves differently from complex $K$-theory. Eventually, Singer pointed out that a different notion of reality was needed. The essential point is that the Fourier transform of a real-valued function is not real but instead satisfies the relation $f(-x)={\overline{f(x)}}$. It was therefore necessary to develop a new version of real $K$-theory for spaces with involution. This was carried out in [6]. In fact $KR$-theory, as I called it, turned out to provide a better approach even for purely topological purposes. ([8], p. 4)
An object in this category is a pair $(X,\tau)$ where $X$ is a topological space (compact or locally compact) with involution $\tau$. (This is what Atiyah refers to as a “Real space.” We will not use that terminology.) Given such a pair $(X,\tau)$ when $X$ is compact, we will define a complex vector bundle with involution to be a pair $(E,\sigma)$ where $E$ is a complex vector bundle over $X$ and $\sigma$ is a conjugate-linear involution on $E$ that satisfies $\tau(\pi(e))=\pi(\sigma(e))$ for all $e\in E$.
Then, again following Atiyah, we define $\rm{Vect}(X,\tau)$ to be the abelian semigroup of isomorphism classes of complex vector bundles with involution over $(X,\tau)$ and
$KR^{0}(X,\tau)=\rm{Groth}(\rm{Vect}(X,\tau)).$
We assert that $KR^{*}$-theory for spaces with involution is determined by $KO_{*}$- theory for an associated real $C^{*}$-algebra. Here is the precise statement.
Theorem 16.1.
  1. (1)
    Let $(X,\tau)$ be a compact topological space with involution. Define a real $C^{*}$-algebra
    $A=A(X,\tau)=\{f\in C(X,{\mathbb{C}})\mid f(\tau(x))=\overline{f(x)}\}.\;$
    Then there is a natural isomorphism
    $\Theta:KO_{0}(A)\overset{\cong}{\longrightarrow}KR^{0}(X,\tau).$
  2. (2)
    Let $(X,\tau)$ be a locally compact topological space with involution. Define
    $A=\{f\in C_{0}(X,{\mathbb{C}})\mid f(\tau(x))=\overline{f(x)}\}\;.$
    Then there is a natural isomorphism
    $\Theta:KO_{0}(A)\overset{\cong}{\longrightarrow}KR^{0}(X,\tau).$
One should compare Karoubi [55] Theorem 4.1 for a related result.
Before beginning the proof, we establish a basic property of complex involutive bundles over spaces with involution.
Lemma 16.2.
Let $(X,\tau)$ be a compact space with involution and let $(E,\sigma)$ be a complex vector bundle over $X$ with a conjugate linear involution satisfying $\sigma(E_{x})=E_{\tau(x)}$. Then $(E,\sigma)$ is isomorphic (as a complex vector bundle with involution) to a sub-bundle $(E^{\prime},\sigma )$ of the trivial bundle $(X\times{\mathbb{C}}^{n},\tau\times c_{n})$ for some $n$ (where $c_{n}$ is complex conjugation on ${\mathbb{C}}^{n}$). Further, there is a subbundle $(F,\rho)$ of the trivial bundle such that
$(E^{\prime},\sigma^{\prime})\oplus(F,\rho)\cong(X\times{\mathbb{C}}^{n},\tau\times c_{n})$
as bundles with involution.
Proof.
First we work locally. We claim that for all $x_{0}\in X$ there exists a neighborhood $\mathcal{U}$ of $x_{0}$ in $X$ such that $\tau(\mathcal{U})=\mathcal{U}$ and $(E_{\mathcal{U}},\sigma)$ is isomorphic to $(\mathcal{U}\times{\mathbb{C}}^{n},\tau\times c_{n})$
Let $x_{0}\in X$ be given and let $\mathcal{U}$ be a neighborhood of $x_{0}$ on which $E$ is trivial. If $\tau(x_{0})\neq x_{0}$ we can arrange without loss of generality that $\mathcal{U}$ is a disjoint union of $\mathcal{V}$ and $\tau(\mathcal{V})$. Let $\phi\colon E_{\mathcal{V}}\rightarrow\mathcal{V}\times{\mathbb{C}}^{n}$ be an isomorphism; we use the notation $\phi_{x}\colon E_{x}\rightarrow{\mathbb{C}}^{n}$ for the map on the fibre. Extend $\phi$ to $E_{\mathcal{\tau(\mathcal{V})}}\rightarrow\tau(\mathcal{V})\times{\mathbb{C}}^{n}$ by
$\phi(\tau(x),v)=(\tau(x),c_{n}(\phi_{x}(\sigma_{\tau(x)}(v))))\;.$
Now $\phi\colon E_{\mathcal{U}}\rightarrow\mathcal{U}\times{\mathbb{C}}^{n}$ is defined and satisfies $(\tau\times c_{n})\circ\phi=\phi\circ\sigma$.
[…]
Since $\widetilde{\sigma}_{x}$ is conjugate-linear and depends continuously on $x$, we can write
$\widetilde{\sigma}_{x}=c_{n}\circ A_{x}$
where $A_{x}$ is an $n\times n$ matrix depending continuously on $x$. The condition that $\widetilde{\sigma}$ is an involution implies that ${A_{x}}\overline{A_{\tau(x)}}=I_{n}$ for all $x\in\mathcal{U}$. Indeed, $(\widetilde{\sigma})^{2}=\mathrm{id}$ can be rewritten as
$c_{n}A_{\tau(x)}c_{n}A_{x}v=v\qquad for\qquad v\in{\mathbb{C}}^{n}$
and then again as $\overline{A_{\tau(x)}}A_{x}v=v$, which is equivalent to
$\overline{A_{\tau(x)}}A_{x}=I_{n}\qquad or\qquad{A_{x}}\overline{A_{\tau(x)}}=I_{n}.$
Furthermore, we can assume without loss of generality that $A_{x_{0}}=I_{n}$, by use of the isomorphism
$\mathrm{id}\times A_{x_{0}}^{-1}\colon(\mathcal{U}\times{\mathbb{C}}^{n})\rightarrow(\mathcal{U}\times{\mathbb{C}}^{n})\;.$
The isomorphism $\psi$ that we are looking for is from $\mathcal{U}\times{\mathbb{C}}^{n}$ to itself, so it can be represented by a continuous matrix-valued function $B_{x}$ for $x\in\mathcal{U}$. We need $\psi$ to carry the involution $\widetilde{\sigma}$ to the involution $c_{n}$ on ${\mathbb{C}}^{n}$. In other words, we need the diagram
[…]
to commute, which can be expressed as
$(B_{\tau(x)}\circ c_{n}\circ A_{x})(v)=(c_{n}\circ B_{x})(v).$
This condition is equivalent to matrix equations $B_{\tau(x)}\overline{A_{x}}=\overline{B_{x}}$ for all $x\in\mathcal{U}$.
We know that there is a continuous square root defined on square matrices in a neighborhood of the identity. We replace $\mathcal{U}$ with a smaller neighborhood to insure that square roots are defined on $A_{x}$ for $x\in\mathcal{U}$, and let $B_{x}=(A_{x})^{1/2}$. Then using the formula $A_{\tau(x)}=(\overline{A_{x}})^{-1}$ we have $B_{\tau(x)}=(\overline{A_{x}})^{-1/2}$ so
${B_{\tau(x)}}\overline{A_{x}}=(\overline{A_{x}})^{-1/2}\overline{A_{x}}=(\overline{A_{x}})^{1/2}=\overline{B_{x}}$
as desired.
Now, using a partition of unity, these local maps can be patched to form an injective homomorphism of vector bundles-with-involution from
$(E,\sigma).\longrightarrow(X\times{\mathbb{C}}^{m},\tau\times c_{n})$
for some large enough integer $m$, as in the proof of Lemma 15.4.
Finally, take $(E,\sigma)=(E,\tau\times c^{n})$ to be a subbundle of the trivial bundle. Then take $F$ to be the point-wise orthogonal complement to $E$ again with the involution $\tau\times c^{n}$. Since vector conjugation on ${\mathbb{C}}^{n}$ respects the inner product we have the statement that
$(\tau\times c^{n})(E)=E\Longrightarrow(\tau\times c^{n})(F)=F.$
Clearly
$(E,\tau\times c^{n})\oplus(F,\tau\times c^{n})=(X\times{\mathbb{C}}^{n},\tau\times c^{n}).$
The rest of this section is devoted to a new proof of Theorem 16.1. Let $(X,\tau)$ be a compact space with involution and let $A=A(X,\tau)$. We shall show that there is a natural isomorphism of abelian semigroups
$\Theta\colon\mathcal{P}_{\infty}(A)\rightarrow\rm{Vect}(X,\tau)$
which implies that there is an induced natural isomorphism of groups
$\Theta\colon KO_{0}(A)\rightarrow KR^{0}(X,\tau).$
Proof.
Our proof follows the standard route.
  1. (1)
    Define the map $\Theta$.
  2. (2)
    Show that $\Theta$ is well-defined.
  3. (3)
    Show that $\Theta$ is 1-1.
  4. (4)
    Show that $\Theta$ is onto.
  5. (5)
    Check that $\Theta$ is natural with respect to maps of spaces with involution.
First let $p\in M_{n}(A)$ be a projection. Then $p\in M_{n}(C(X,{\mathbb{C}}))$ and $p(\tau(x))=\overline{p(x)}$. Let
$E={\rm ran~}p$ $=\{(x,v)\in X\times{\mathbb{C}}^{n}\mid v\in p(x)\cdot{\mathbb{C}}^{n}\}$
$=\{(x,v)\in X\times{\mathbb{C}}^{n}\mid p(x)v=v\}$
with the associated projection map $\pi\colon E\rightarrow X$ by $\pi(x,v)=x$. Now $E$ is a complex vector bundle and has a conjugate-linear involution $\sigma=\tau\times c_{n}$ given by $(x,v)\mapsto(\tau(x),\overline{v})$. This involution $\sigma=\tau\times c_{n}$ certainly commutes with the respective involutions on $X$ as required. To show that $(\tau(x),\overline{v})\in E$ we check that
$p(\tau(x))\overline{v}=\overline{p(x)}\,\overline{v}=\overline{p(x)v}=\overline{v}\;\;.$
Hence $(E,\sigma)\in\rm{Vect}(X,\tau)$. We define $\Theta(p)=(E,\sigma)$.
We must show that $\Theta$ is well-defined. Suppose that $p$ and $q$ are projections in $M_{n}(A)$ and
$p]=[q]\in\mathcal{P}_{\infty}(A).$
There is a unitary $u\in M_{n}(A)$ such that $upu^{*}=q$. Let
$E_{p}$ $=\{(x,v)\in X\times{\mathbb{C}}^{n}\mid v\in p(x)v=v\}$
$E_{q}$ $=\{(x,v)\in X\times{\mathbb{C}}^{n}\mid v\in q(x)v=v\}$
As in the proof of Lemma 15.3, there is a vector bundle homomorphism
$\phi:E_{p}\to E_{q}$
defined by
$\phi(x,v)=(x,u(x)v).$
The bundle homomorphism $\phi$ commutes with the involution $\sigma$:
$\phi(\sigma(x,v))=\phi(\tau(x),\overline{v})=(\tau(x),u(\tau(x))\overline{v})=(\tau(x),\overline{u}\;\overline{v})=\sigma(x,u(x)v)=\sigma(\phi(x,v))\;.$
It follows that $E_{p}$ and $E_{q}$ are isomorphic objects in the category of complex vector bundles with involution over $(X,\tau)$.
Thus the construction $p\mapsto(E,\sigma)$ defines a homomorphism of abelian semigroups
$\Theta\colon\mathcal{P}_{\infty}(A)\rightarrow\rm{Vect}(X,\tau).$
Next, we claim that $\Theta$ is 111-1. Assume that $p,q$ in $M_{n}(A)$ are projections that satisfy $E_{p}\cong E_{q}$. Similar to the proof of Lemma 15.4, there exists $a\in C(X,M_{n}({\mathbb{C}}))$ where $a_{x}$ is an isomorphism from $p(x){\mathbb{C}}^{n}$ to $q(x){\mathbb{C}}^{n}$ for each $x$. The hypothesis that this isomorphism commutes with the involutions on $E_{p}$ and $E_{q}$ is expressed by the formula $a_{\tau(x)}(\overline{v})=\overline{a_{x}(v)}$, which in turn implies that $a_{\tau(x)}(\overline{v})=\overline{a_{x}}\,\overline{v}$. Hence $a_{\tau(x)}=\overline{a_{x}}$ and so $a\in M_{n}(A)$.
As before, we let $s=f(a^{*}a)a^{*}$ where
$f(t)=\begin{cases}0&t=0\\ t^{-1/2}&t>0\end{cases}$
noting in particular that because $a\in M_{n}(A)$, we also have $s\in M_{n}(A)$. As before, $s$ is a partial isometry that satisfies $ss^{*}=p$ and $s^{*}s=q$.
Since
$u=\begin{bmatrix}s&1-q\\ 1-p&s^{*}\end{bmatrix}\;$
is still in $M_{2n}(A)=C(X,M_{2n}({\mathbb{R}}))$ and we still have
$u\begin{bmatrix}p&0\\ 0&0\end{bmatrix}u^{*}=\begin{bmatrix}q&0\\ 0&0\end{bmatrix}\;$,
we have again that $\mathrm{diag}(p,0)$ is unitarily equivalent to $\mathrm{diag}(q,0)$ in $M_{2n}(A)$, so $[p]=[q]$ in $\rm{Proj}(A)$. Thus $\Theta$ is 1-1.
Next, we show that $\Theta$ is onto. Let $(E,\sigma)$ be any complex vector bundle with involution over $(X,\tau)$. Using Lemma 16.2, identify $E$ with a sub-bundle of $X\times{\mathbb{C}}^{n}$ in such a way that the involution $\sigma$ on $E$ corresponds to conjugation on ${\mathbb{C}}^{n}$. In other words, $\sigma$ can be expressed by the formula
$\sigma(x,v)=(\tau(x),\overline{v})$
where $\overline{v}$ is component-wise conjugation on ${\mathbb{C}}^{n}$. Define $p(x)\in C(X,M_{n}({\mathbb{C}}))$ by letting $p(x)$ be the orthogonal projection of ${\mathbb{C}}^{n}$ to $E_{x}$. Then $p$ is continuous in $x$, and is a projection: $p^{2}=p^{*}=p$. Furthermore, we claim that our condition on $\sigma$ implies that
$p(\tau(x))=\overline{p(x)}\;.$
Indeed, suppose that $v\in E_{x}$ is in the range of $p(x)$. Then
$\overline{p(\tau(x))}\cdot v=p(x)\cdot v=v\;$
which shows that $v$ is in the range of $\overline{p(\tau(x))}$. As $\overline{p(\tau(x))}$ and p(x){p(x)} are both orthogonal projections with the same rank, they must be equal, implying that $\overline{p(\tau(x))}={p(x)}$ and hence $p(\tau(x))=\overline{p(x)}\;$. From this it follows that $p$ is a projection in $M_{n}(A)$. By construction we have $(E,\sigma)=(E_{p},\overline{~\cdot~})$, which shows that $\Theta$ is surjective.
Thus we have established that if $X$ is compact then there is an isomorphism of abelian semigroups
$\Theta\colon\mathcal{P}_{\infty}(A)\rightarrow\rm{Vect}(X,\tau)$
which implies that there is an induced group isomorphism
$\Theta\colon KO_{0}(A)\rightarrow KR^{0}(X,\tau).$
It remains to establish that $\Theta$ is natural with respect to maps of spaces with involution. Suppose that
$f\colon(X_{1},\tau_{1})\rightarrow(X_{2},\tau_{2})$
is a continuous homomorphism of compact spaces with involution. Then there is a $C^{*}$-algebra homomorphism
$f^{*}:A_{2}=A(X_{2},\tau_{2})\longrightarrow A_{1}=A(X_{1},\tau_{1})$
with associated map
$f^{*}:\mathcal{P}_{\infty}(A_{2})\longrightarrow\mathcal{P}_{\infty}(A_{1}).$
Similarly, the pullback construction induces a natural map
$f^{*}:\rm{Vect}(X_{2},\tau_{2})\longrightarrow\rm{Vect}(X_{1},\tau_{1})$
and it is easy to check that the diagram
[…]
commutes. This implies immediately that the diagram
[…]
commutes, establishing naturality. This completes the proof of Theorem 16.1 in the case where $X$ is compact.
The proof in the locally compact case follows the same strategy as that in the proof of Theorem 15.2. Note that if $X$ is locally compact then any involution $\tau$ on $X$ extends naturally to an involution $\tau^{+}$ on $X^{+}$, which satisfies $\tau^{+}(+)=+$. Thus $KR^{0}(X,\tau)$ is defined to be the kernel of
$\ker\left[KR^{0}(X^{+},\tau^{+})\xrightarrow{\iota_{*}}KR^{0}(+,\mathrm{id})\right]\;.$
We leave the details to the reader. ∎
Remark 16.3.
As far as we know, the product structure on $A=\real(X,\tau)$ with $\tau$ non-trivial has not been investigated. If we identify $KO_{*}(\real(X,\tau))$ with Atiyah’s Real $K$-theory $KR^{*}(X,\tau)$, then there is a natural coproduct
$(X,\tau)\longrightarrow(X,\tau)\times(X,\tau)$
that induces a product structure on $KO_{*}(\real(X,\tau))\cong KR^{*}(X)$ exactly analogous to the product structures on $KO^{*}(X)$ and $K^{*}(X)$. This structure needs more attention than we can give it at this time.
Part III Clifford Algebras and Symmetric Spaces
At first glance, there is no particular reason to suspect that a space such as $\Omega^{j}(\mathbf{O})$ has any special relationship to physics. It turns out that these spaces have a close relationship with certain symmetric spaces, and in fact the eight spaces in the real Bott Periodicity Theorem and the two in the Complex Bott Periodicity Theorem correspond to the renowned Ten-Fold Way in physics. We can see this connection to symmetric spaces best via Clifford algebras. So in Part 3 we will introduce Clifford algebras and show their connection to Bott Periodicity and to symmetric spaces. In Part 4 we will relate symmetric spaces to physics directly.
17. Clifford Algebras: Clifford and Lie algebra
The classic reference for Clifford algebras and their connection to $K$-theory is Atiyah, Bott, and Shapiro [9]. Husemoller [44] has a fine treatment. For connections with physics we rely on Baez [25], Stone et.al. [81] and G. Moore [64]. Notation for Clifford algebras varies a lot. When in doubt, please consult the chart below, which gives our notation.
Definition 17.1.
The real Clifford algebra ${\mathcal{C}}_{k}$ is defined to be the universal $\mathbb{R}$}-algebra generated by a unit $1$ and symbols $e_{1},\dots,e_{k}$ modulo the relations
$e_{j}^{2}=-1\qquad\qquad e_{i}e_{j}+e_{j}e_{i}=0\qquad i\neq j$
The complex Clifford algebra ${\mathcal{C}}_{k}^{c}$ is defined by
${\mathcal{C}}_{k}^{c}={\mathcal{C}}_{k}\otimes_{{\mathbb{R}}}{\mathbb{C}}.$
There are natural inclusions
${\mathcal{C}}_{k}\hookrightarrow{\mathcal{C}}_{k+1}\qquad\qquad and\qquad\qquad{\mathcal{C}}_{k}^{c}\hookrightarrow{\mathcal{C}}_{k+1}^{c}.$
Let’s look at the first few of these. For ${\mathcal{C}}_{0}$ we are handed $\mathbb{R}$} itself. The algebra ${\mathcal{C}}_{1}$ is given by additively by two copies of $\mathbb{R}$}, one for $1$ and one for $e_{1}$, with $e_{1}^{2}=-1$ and is easily seen to be the complex numbers $\mathbb{C}$}. For ${\mathcal{C}}_{2}$ we have additive generators $1$, $e_{1}$, $e_{2}$ and $e_{1}e_{2}$, and
$(e_{1}e_{2})^{2}=e_{1}e_{2}e_{1}e_{2}=-(e_{1}e_{1})(e_{2}e_{2})=-1$
and this is just the quaternions $\mathbb{H}$}.
Complexifying is easy as well.
${\mathcal{C}}_{0}^{c}\cong{\mathbb{R}}\otimes_{\mathbb{R}}{\mathbb{C}}\cong{\mathbb{C}}\qquad\qquad\qquad{\mathcal{C}}_{1}^{c}\cong{\mathbb{C}}\otimes_{\mathbb{R}}{\mathbb{C}}\cong{\mathbb{C}}\oplus{\mathbb{C}}.$
Here is a list of the first two complexified Clifford algebras, the first eight real Clifford algebras, and a statement of their periodicity.
Theorem 17.2.
[9]
  1. (1)
    In low dimensions, the Clifford algebras are given as follows:
    name Clifford algebra
    ${\mathcal{C}}_{0}^{c}$ $\mathbb{C}$}
    ${\mathcal{C}}_{1}^{c}$ $\mathbb{C}$$\oplus{\mathbb{C}}$
    ${\mathcal{C}}_{0}$ $\mathbb{R}$}
    ${\mathcal{C}}_{1}$ $\mathbb{C}$}
    ${\mathcal{C}}_{2}$ $\mathbb{H}$}
    ${\mathcal{C}}_{3}$ $\mathbb{H}$$\oplus{\mathbb{H}}$
    ${\mathcal{C}}_{4}$ $M_{2}({\mathbb{H}})$
    ${\mathcal{C}}_{5}$ $M_{4}({\mathbb{C}})$
    ${\mathcal{C}}_{6}$ $M_{8}({\mathbb{R}})$
    ${\mathcal{C}}_{7}$ $M_{8}({\mathbb{R}})\oplus M_{8}({\mathbb{R}})$
    where $M_{k}({\mathbb{F}})$ denotes $k\times k$-matrices with entries in $\mathbb{F}$}.
  2. (2)
    All other real and complex Clifford algebras are determined by the periodicity formulas:
    ${\mathcal{C}}_{k+8}\,\cong\,{\mathcal{C}}_{k}\otimes M_{16}({\mathbb{R}})\qquad\qquad{\mathcal{C}}^{c}_{k+2}\,\cong\,{\mathcal{C}}^{c}_{k}\otimes M_{2}({\mathbb{C}})$
  3. (3)
    $dim_{\mathbb{R}}({\mathcal{C}}_{k})=2^{k}\qquad\qquad dim_{\mathbb{C}}({\mathcal{C}}_{k}^{c})=2^{k}$
Each Clifford algebra $A$ has a Lie group $Lie(A)\subset A$ associated to it, namely the elements in it that are “unitary” in the sense that $aa^{*}=a^{*}a=1$. We write
${\mathcal{L}}_{j}=Lie({\mathcal{C}}_{j})\qquad\qquad{\mathcal{L}}_{j}^{c}=Lie({\mathcal{C}}_{j}^{c}).$
We can easily see what these are in low dimensions. We can also see what happens as $n\to\infty$ and that the result is periodicity of period 2 (in the complex case) and 8 (in the real case.)
name Clifford algebra Lie group      ${\mathcal{L}}_{j}$ stable Lie
${\mathcal{C}}_{0}^{c}$ $\mathbb{C}$} ${\mathcal{L}}_{0}^{c}=U_{1}$ $\mathbf{U}$
${\mathcal{C}}_{1}^{c}$ $\mathbb{C}$$\oplus{\mathbb{C}}$ ${\mathcal{L}}_{1}^{c}=U_{!}\times U_{1}$ $\mathbf{U}\times\mathbf{U}$
${\mathcal{C}}_{0}$ $\mathbb{R}$} ${\mathcal{L}}_{0}=O_{1}$ $\mathbf{O}$
${\mathcal{C}}_{1}$ $\mathbb{C}$} ${\mathcal{L}}_{1}=U_{1}$ $\mathbf{U}$
${\mathcal{C}}_{2}$ $\mathbb{H}$} ${\mathcal{L}}_{2}=Sp_{1}$ $\mathbf{Sp}$
${\mathcal{C}}_{3}$ $\mathbb{H}$$\oplus{\mathbb{H}}$ ${\mathcal{L}}_{3}=Sp_{1}\times Sp_{1}$ $\mathbf{Sp}\times\mathbf{Sp}$
${\mathcal{C}}_{4}$ $M_{2}({\mathbb{H}})$ ${\mathcal{L}}_{4}=Sp_{2}$ $\mathbf{Sp}$
${\mathcal{C}}_{5}$ $M_{4}({\mathbb{C}})$ ${\mathcal{L}}_{5}=U_{4}$ $\mathbf{U}$
${\mathcal{C}}_{6}$ $M_{8}({\mathbb{R}})$ ${\mathcal{L}}_{6}=O_{8}$ $\mathbf{O}$
${\mathcal{C}}_{7}$ $M_{8}({\mathbb{R}})\oplus M_{8}({\mathbb{R}})$ ${\mathcal{L}}_{7}=O_{8}\times O_{8}$ $\mathbf{O}\times\mathbf{O}$
${\mathcal{C}}_{8}$ $M_{16}({\mathbb{R}})$ ${\mathcal{L}}_{8}=O_{16}$ $\mathbf{O}$
Note that the order of the Clifford algebras
${\mathcal{C}}_{0}^{c}\hookrightarrow{\mathcal{C}}_{1}^{c}\hookrightarrow{\mathcal{C}}_{2}^{c}\hookrightarrow{\mathcal{C}}_{3}^{c}\hookrightarrow\dots$
and
${\mathcal{C}}_{0}\hookrightarrow{\mathcal{C}}_{1}\hookrightarrow{\mathcal{C}}_{2}\hookrightarrow{\mathcal{C}}_{3}\hookrightarrow\dots$
corresponds nicely to the order of the inclusions of the stable groups. The group $\mathbf{U}=\cup U_{n}({\mathbb{C}})$ repeats in two steps
$\mathbf{U}\,\hookrightarrow\,\mathbf{U}/(\mathbf{U}\times\mathbf{U})\,\hookrightarrow\,\mathbf{U}$
and similarly in the real and symplectic setting there are eight steps, namely
$\mathbf{O}\hookrightarrow\mathbf{U}\hookrightarrow\mathbf{Sp}\hookrightarrow\mathbf{Sp}\times\mathbf{Sp}\hookrightarrow\mathbf{Sp}\hookrightarrow\mathbf{U}\hookrightarrow\mathbf{O}\,\hookrightarrow\mathbf{O}\times\mathbf{O}\hookrightarrow\mathbf{O}$
Remark 17.3.
We are particularly interested in the relationship between the Lie groups, and so we reconstruct the chart as follows, showing the various natural inclusions. The complex situation is fairly simple:
[…]
The real case is analogous, but of course more complicated. Note that every vertical arrow below is an inclusion.
[…]
Note that the “limit space” column is exactly the list of the Big Ten spaces in the order previously presented in the Homotopy Road Map. Here it is again:
The Homotopy Road Map: Step I (Table 11.2 redux)
finite approx Big Ten Bott ${\mathcal{R}}_{k}$ $\pi_{0}({\mathcal{R}}_{k})$
$_{n} 𝐔\mathbf{U}$ $\Omega^{0}\mathbf{U}$ ${\mathcal{R}}_{1}^{c}$ 0
$U_{2n}/(U_{n}\times U_{n})$ $\mathbf{BU}\times{\mathbb{Z}}$ $\Omega^{1}\mathbf{U}$ ${\mathcal{R}}_{0}^{c}$ $\mathbb{Z}$}
$_{n} 𝐎\mathbf{O}$ $\Omega^{0}\mathbf{O}$ ${\mathcal{R}}_{1}$ ${\mathbb{Z}}_{2}$
$O_{2n}/U_{n}$ $\mathbf{O}/\mathbf{U}$ $\Omega^{1}\mathbf{O}$ ${\mathcal{R}}_{2}$ ${\mathbb{Z}}_{2}$
$U_{n}/Sp_{n}$ $\mathbf{U}/\mathbf{Sp}$ $\Omega^{2}\mathbf{O}$ ${\mathcal{R}}_{3}$ 0
$Sp_{2n}/(Sp_{n}\times Sp_{n})$ $\mathbf{BSp}\times{\mathbb{Z}}$ $\Omega^{3}\mathbf{O}$ ${\mathcal{R}}_{4}$ $\mathbb{Z}$}
$_{n} 𝐒𝐩\mathbf{Sp}$ $\Omega^{4}\mathbf{O}$ ${\mathcal{R}}_{5}$ 0
$Sp_{n}/U_{n}$ $\mathbf{Sp}/\mathbf{U}$ $\Omega^{5}\mathbf{O}$ ${\mathcal{R}}_{6}$ 0
$U_{n}/O_{n}$ $\mathbf{U}/\mathbf{O}$ $\Omega^{6}\mathbf{O}$ ${\mathcal{R}}_{7}$ 0
$O_{2n}/(O_{n}\times O_{n})$ $\mathbf{BO}\times{\mathbb{Z}}$ $\Omega^{7}\mathbf{O}$ ${\mathcal{R}}_{0}$ $\mathbb{Z}$}
This shows a striking connection between the structure of Clifford algebras (a clearly purely algebraic matter, at least at the start) to Bott periodicity. This very deep connection was first observed by Atiyah, Bott, Shapiro [9] (1964), who say in their introduction:
“It is this theorem which shows the significance of Clifford algebras in $K$-theory and strongly suggests that one should look for a proof of the periodicity theorem using Clifford algebras. Since this paper was written a proof on these lines has in fact been found by R. Wood. [89] (1966)”
Wood did more than this - he generalized $K$-theory from compact spaces to Banach algebras. It is his proof of periodicity in that context, together with Karoubi’s monumental work (we list many papers of his in the bibliography, some with Villamayor) which made possible the use of $K$-theory in functional analysis and the subsequent explosive developments due to Brown-Douglas-Fillmore, Kasparov, Atiyah, Karoubi, Connes, and many others in the 1970’s, 1980’s and to this day.
We may realize the the Lie groups ${\mathcal{L}}_{j}$ of the Clifford algebras in a different way. Fix some large index $t$ and integer $k$. Fix a faithful $C^{*}$-representation
$\phi:{\mathcal{C}}_{k}\longrightarrow M_{16t}({\mathbb{R}}).$
Following the custom in physics, we let $J_{i}=\phi(e_{i})\in O_{16t}$. Then $J_{r}J_{s}=-J_{s}J_{r}$ for $r\neq s$ and $J_{r}^{2}=-1$. Let
$\big<O_{16t},\,\,\,J_{1},\dots J_{n}\big$>
denote the subgroup of $O_{16t}$ consisting of all of the matrices that commute with $J_{1},\dots J_{n}$. Then we have the following result.
Theorem 17.4.
(cf. Stone et. al. [81], §2)
${\mathcal{L}}_{n}=\big<O_{16t},\,\,\,J_{1},\dots J_{n}\big$>
More precisely,
$U_{8t}=\big<O_{16t},\,\,\,J_{1}\big$>
$Sp_{4t}=\big<O_{16t},\,\,\,J_{1},J_{2}\big$>
$Sp_{2t}\times Sp_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3}\big$>
$Sp_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4}\big$>
$U_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5}\big$>
$O_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6}\big$>
$O_{t}\times O_{r}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6},J_{7}\big$>
$O_{t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6},J_{7},J_{8}\big$>
The sequence may be extended infinitely in each direction when $t$ is a sufficiently large power of $2$. The analogous result holds in the complex setting as well.
The proof of this result in [81] is quite concrete. It references Milnor [62] for the construction. We shall return to this construction when we introduce symmetric spaces.
Remark 17.5.
We want to highlight something that has confused us no end when looking casually at the situation. Remark 17.3 gives us the following list for the Lie groups, with inclusions left to right corresponding to the increasing sequence of Clifford algebras:
$O_{1}\hookrightarrow U_{1}\hookrightarrow Sp_{1}\hookrightarrow Sp_{1}\times Sp_{1}\hookrightarrow Sp_{2}\hookrightarrow U_{4}\hookrightarrow O_{8}\hookrightarrow O_{8}\times O_{8}\hookrightarrow O_{16}\qquad(**)$
whereas Theorem 17.4, following Stone et al [81] gives the following list, corresponding to the increasing number of $J_{i}$ being used and the corresponding decreasing sequence of Lie groups and Lie algebras, and with inclusions right to left:
$O_{16}\hookleftarrow U_{8}\hookleftarrow Sp_{4}\hookleftarrow Sp_{2}\times Sp_{2}\hookleftarrow Sp_{2}\hookleftarrow U_{2}\hookleftarrow O_{2}\hookleftarrow O_{1}\times O_{1}\hookleftarrow O_{1}\qquad(@@)$
If we take larger and larger groups, to reach stability, we wind up with two contrasting sequences, namely
$\mathbf{O}\,\overset{1}{\hookrightarrow}\,\mathbf{U}\,\overset{2}{\hookrightarrow}\,\mathbf{Sp}\,\overset{3}{\hookrightarrow}\,\mathbf{Sp}\times\mathbf{Sp}\,\overset{4}{\hookrightarrow}\,\mathbf{Sp}\,\overset{5}{\hookrightarrow}\,\mathbf{U}\,\overset{6}{\hookrightarrow}\,\mathbf{O}\,\overset{7}{\hookrightarrow}\,\mathbf{O}\times\mathbf{O}\,\overset{8}{\hookrightarrow}\,\mathbf{O}.$
from (**) and
$\mathbf{O}\,\overset{6}{\hookleftarrow}\,\mathbf{U}\,\overset{5}{\hookleftarrow}\,\mathbf{Sp}\,\overset{4}{\hookleftarrow}\,\mathbf{Sp}\times\mathbf{Sp}\,\overset{3}{\hookleftarrow}\,\mathbf{Sp}\,\overset{2}{\hookleftarrow}\,\mathbf{U}\,\overset{1}{\hookleftarrow}\,\mathbf{O}\,\overset{8}{\hookleftarrow}\,\mathbf{O}\times\mathbf{O}\,\overset{7}{\hookleftarrow}\,\mathbf{O}.$
from (@@). So the groups appear in opposite order. What is really neat is that the maps are the same - just re-ordered! We have inserted the numbering of the maps simply to facilitate comparison. The same thing happens in the complex setting, but there are only two groups involved, and so it passes without notice. We want to refer to these two different points of view.
Definition 17.6.
  1. (1)
    The Clifford perspective will entail regarding the ordering
    $\overset{8}{\hookrightarrow}\,\mathbf{O}\,\overset{1}{\hookrightarrow}\,\mathbf{U}\,\overset{2}{\hookrightarrow}\,\mathbf{Sp}\,\overset{3}{\hookrightarrow}\,\mathbf{Sp}\times\mathbf{Sp}\,\overset{4}{\hookrightarrow}\,\mathbf{Sp}\,\overset{5}{\hookrightarrow}\,\mathbf{U}\,\overset{6}{\hookrightarrow}\,\mathbf{O}\,\overset{7}{\hookrightarrow}\,\mathbf{O}\times\mathbf{O}\,\overset{8}{\hookrightarrow}\,\mathbf{O}\overset{1}{\hookrightarrow}\$
    with associated homogeneous spaces ordered correspondingly:
    $\mathbf{U}/\mathbf{O}\,\quad\,\mathbf{Sp}/\mathbf{U}\,\quad\,(\mathbf{Sp}\times\mathbf{Sp})/\mathbf{Sp}\,\quad\,\mathbf{Sp}/(\mathbf{Sp}\times\mathbf{Sp})\,\quad\,\mathbf{U}/\mathbf{Sp}\,\quad\,\mathbf{O}/\mathbf{U}\,\quad\,(\mathbf{O}\times\mathbf{O})/\mathbf{O}\,\quad\,\mathbf{O}/(\mathbf{O}\times\mathbf{O})$
    arising from the inclusions of Clifford algebras ${\mathcal{C}}_{0}\subset{\mathcal{C}}_{1}\dots$ as fundamental.
  2. (2)
    The Lie algebra perspective will entail regarding the ordering
    $\overset{7}{\hookleftarrow}\,\mathbf{O}\,\overset{6}{\hookleftarrow}\,\mathbf{U}\,\overset{5}{\hookleftarrow}\,\mathbf{Sp}\,\overset{4}{\hookleftarrow}\,\mathbf{Sp}\times\mathbf{Sp}\,\overset{3}{\hookleftarrow}\,\mathbf{Sp}\,\overset{2}{\hookleftarrow}\,\mathbf{U}\,\overset{1}{\hookleftarrow}\,\mathbf{O}\,\overset{8}{\hookleftarrow}\,\mathbf{O}\times\mathbf{O}\,\overset{7}{\hookleftarrow}\,\mathbf{O}\overset{6}{\hookleftarrow}\,$
    with associated homogeneous spaces
    $\quad\,\mathbf{O}/\mathbf{U}\,\quad\,\mathbf{U}/\mathbf{Sp}\,\quad\,\mathbf{Sp}/(\mathbf{Sp}\times\mathbf{Sp})\,\quad\,(\mathbf{Sp}\times\mathbf{Sp})/\mathbf{Sp}\,\quad\,\mathbf{Sp}/\mathbf{U}\,\quad\mathbf{U}/\mathbf{O}\,\quad\,\mathbf{O}/(\mathbf{O}\times\mathbf{O})\,\quad\,(\mathbf{O}\times\mathbf{O})/\mathbf{O}$
    arising from using the $J_{i}$ to shrink the Clifford algebras as fundamental. Note that the maps in the Lie algebra perspective match exactly to the maps in the Clifford perspective and hence exactly the same homogeneous spaces appear, albeit in reverse order!
18. Clifford Algebras and Bott Periodicity: the Real Case
In this section and the next we outline the Wood-Karoubi proofs of Bott periodicity via Clifford algebras. We try to balance the detail, and we cite but don’t present the one hard homotopy step involved in the proof. The real case is significantly longer in the sense that 8 is greater than 2, but the mathematical steps are essentially the same. In this section we present the real case. In the next section we present the complex case.
Definition 18.1.
Suppose that $A$ is a unital real Banach algebra. Let ${\mathcal{L}}A$ denote the Lie group associated to $A$. Let ${\mathcal{G}}A$ denote the group of invertible elements of $A$. Then ${\mathcal{L}}A\subset{\mathcal{G}}A$ and this map is a deformation retraction (essentially, by Gram-Schmidt orthogonalization) and hence a homotopy equivalence.
Using the larger groups has some technical conveniences in homotopy theory, and they are used by Wood, and so we shall use them as well.
Here’s some notation that we will use:
  1.
  2. (1)
    $A\otimes B=A\otimes_{\mathbb{R}}B$
  3. (2)
    ${\mathcal{G}}_{k}(A)={\mathcal{G}}(A\otimes M_{k}({\mathbb{R}}))$
  4. (3)
[…]
  5. (4)
    If $X$ is a based topological space, then $X_{o}$ denotes the path component of the base point of $X$. (If $X$ is a topological group then we take the identity of the group as the basepoint unless there is a good reason not to.)
There are natural inclusions of real Clifford algebras ${\mathcal{C}}_{r}\longrightarrow{\mathcal{C}}_{r+1}$ which induce a natural commutative diagram with Clifford perspective
[…]
We have seen instances of these spaces before. For example, taking $A={\mathbb{R}}$, we obtain (after a deformation retraction)
${\mathcal{G}}_{\infty}({\mathcal{C}}_{7})/{\mathcal{G}}_{\infty}({\mathcal{C}}_{6})\simeq(\mathbf{O}\times\mathbf{O})/\mathbf{O}=\mathbf{O}={\mathcal{R}}_{1}$
and
${\mathcal{G}}_{\infty}({\mathcal{C}}_{6})/{\mathcal{G}}_{\infty}({\mathcal{C}}_{5})\simeq\mathbf{O}/\mathbf{U}={\mathcal{R}}_{2}.$
Here is an example of the contrast/confusion generated between the Clifford perspective and the Lie algebra perspective in the numbering. We go “down” from $7\dots 6$ to $6\dots 5$ in the Clifford indexing of the quotients associated to the Clifford algebras and the number goes up from $1$ to $2$ in the Lie algebra indexing of the ${\mathcal{R}}_{j}$.
Next, we place ABS Clifford periodicity on record in the form that we need.
Proposition 18.2.
(Atiyah-Bott-Shapiro [9] ) Let $A$ be a unital real Banach algebra. Then there is a natural homotopy equivalence
$GL_{\infty}(A\otimes{{\mathcal{C}}}_{r+8})\simeq GL_{\infty}(A\otimes{\mathcal{C}}_{r}).$
Proof.
ABS prove that
${\mathcal{C}}_{r+8}\cong{\mathcal{C}}_{r}\otimes M_{16}({\mathbb{R}})$
which implies immediately that
$A\otimes{\mathcal{C}}_{r+8}\cong A\otimes{\mathcal{C}}_{r}\otimes M_{16}({\mathbb{R}}).$
Thus for each $k$,
${\mathcal{G}}_{k}(A\otimes{\mathcal{C}}_{r+8})\cong{\mathcal{G}}_{k}(A\otimes{\mathcal{C}}_{r}\otimes M_{16}({\mathbb{R}}))\cong{\mathcal{G}}_{k+16}(A\otimes{\mathcal{C}}_{r})$
and letting $k\to\infty$ implies the result.
Next, we invoke Wood’s result. It is essentially topological, using the internal structure of the topological groups involved and their symmetries. (Karoubi (cf. his excellent book [52] ) packages the matter differently but ultimately proves the analogous result.) For yet a different packaging, see Milnor [60], §24. For specifics, recall that the maps
${\mathcal{L}}_{r}\overset{\subseteq}{\longrightarrow}{\mathcal{G}}_{\infty}({\mathcal{C}}_{r})$
are homotopy equivalences, and then look at Remark 17.3 for a list of the various ${\mathcal{L}}_{r}$ and, most importantly, ${\mathcal{L}}_{r}/{\mathcal{L}}_{r-1}$.
Theorem 18.3.
Wood [89] ( 4.7, 4.9) Suppose that $A$ is a real unital Banach algebra. Then:
  1. (1)
    There is a weak homotopy equivalence
    $\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1})\big)_{o}\overset{\simeq}{\longrightarrow}\Omega^{1}\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r+1})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r})\big)_{o}.$
  2. (2)
    There is a weak homotopy equivalence
    $\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1})\big)_{o}\overset{\simeq}{\longrightarrow}\Omega^{8}\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r+8})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r+7})\big)_{o}.$
  3. (3)
    There is a weak homotopy equivalence
    $\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1})\big)_{o}\overset{\simeq}{\longrightarrow}\Omega^{8}\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1})\big)_{o}$
  4. (4)
    Taking $A={\mathbb{R}}$ and $r=7$ in the previous result, there is a weak homotopy equivalence
    $\mathbf{O}_{o}\overset{\simeq}{\longrightarrow}\Omega^{8}\mathbf{O}_{o}.$
  5. (5)
    Rephrasing in the Lie algebra perspective and recalling that ${\mathcal{R}}_{1}=\mathbf{O}$, there is a natural homotopy equivalence
    ${\mathcal{R}}_{s}\simeq\Omega^{s}{\mathcal{R}}_{0}$
    which implies that
    $\pi_{a}({\mathcal{R}}_{b})\cong\pi_{0}({\mathcal{R}}_{a+b})\cong\pi_{a+b}({\mathcal{R}}_{0})$
    and, in particular,
    $\pi_{k}(\mathbf{O})=\pi_{k}({\mathcal{R}}_{1})=\pi_{0}({\mathcal{R}}_{k+1}))$
    with the indexing understood to be mod $8$. So we know all of the homotopy groups of $\mathbf{O}$. [^75] This cannot be said for very many spaces. Topologists build Eilenberg-MacLane spaces, for instance, that have exactly one non-zero homotopy group. However, in situations (such as the 2-sphere $S^{2}$, to take a really simple example) where the higher homotopy groups do not all vanish, very little is actually known. Keep in mind that if somebody says that they know the first million homotopy groups of $S^{2}$ then they are talking about 0% of all of the homotopy groups of $S^{2}$.
Note that part (2) follows from part (1): you simply iterate (1) eight times. For part (3) you use the ABS result, Theorem 19.2. Part (4) is simply the recognition that
${\mathcal{G}}_{\infty}({\mathcal{C}}_{7})/{\mathcal{G}}_{\infty}({\mathcal{C}}_{6})_{o}\cong\mathbf{O}_{o}$
The hard part, then, is Part 1, and this is where Wood makes a major contribution.
Please note that this strongly ties Clifford algebras, or, more precisely, their Lie groups, to the Homotopy Road Map. Rather than regarding the first column as coming to us from Bott Periodicity, we may think of the column as simply presenting yet a different look at the Clifford and Lie algebra perspectives.
19. Clifford Algebras and Bott Periodicity: the Complex Case
Here we vaguely sketch Wood’s proof of Bott periodicity in the complex setting.
Definition 19.1.
Suppose that $A$ is a unital complex Banach algebra.
Here’s some notation that we will use in this section:
  1.
  2. (1)
    $A\otimes B=A\otimes_{\mathbb{C}}B$
  3. (2)
    ${\mathcal{G}}_{k}(A)={\mathcal{G}}(A\otimes M_{k}({\mathbb{C}}))$
  4. (3)
[…]
There are natural inclusions of complex Clifford algebras ${\mathcal{C}}^{c}_{r}\longrightarrow{\mathcal{C}}_{r+1}^{c}$ which induce a natural commutative diagram
[…]
Taking $A={\mathbb{C}}$, we obtain (after a deformation retraction)
${\mathcal{G}}_{\infty}({\mathcal{C}}_{1}^{c})/{\mathcal{G}}_{\infty}({\mathcal{C}}_{0}^{c})\simeq(\mathbf{U}\times\mathbf{U})/\mathbf{U}=\mathbf{U}.$
Next, we place ABS Clifford complex periodicity on record in the form that we need.
Proposition 19.2.
(Atiyah-Bott-Shapiro [9] ) Let $A$ be a unital complex Banach algebra. Then there is a natural homotopy equivalence
$GL_{\infty}(A\otimes{\mathcal{C}}_{r+2}^{c})\simeq GL_{\infty}(A\otimes{\mathcal{C}}_{r}^{c}).$
Proof.
ABS prove that
${\mathcal{C}}_{r+2}^{c}\cong{\mathcal{C}}_{r}^{c}\otimes M_{2}({\mathbb{C}})$
which implies immediately that
$A\otimes{\mathcal{C}}_{r+2}^{c}\cong A\otimes{\mathcal{C}}^{c}_{r}\otimes M_{2}({\mathbb{C}}).$
Thus for each $k$,
${\mathcal{G}}_{k}(A\otimes{\mathcal{C}}_{r+2}^{c})\cong{\mathcal{G}}_{k}(A\otimes{\mathcal{C}}^{c}_{r}\otimes M_{2}({\mathbb{C}}))\cong{\mathcal{G}}_{k+2}(A\otimes{\mathcal{C}}_{r}^{c})$
and taking limits, this implies the result.
Next, we invoke Wood’s result.
Theorem 19.3.
Wood [89] ( 4.7, 4.9) Suppose that $A$ is a complex unital Banach algebra. Then:
  1. (1)
    There is a weak homotopy equivalence
    $\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r}^{c})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1}^{c})\big)_{o}\overset{\simeq}{\longrightarrow}\Omega^{1}\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r+1}^{c})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r}^{c})\big)_{o}.$
  2. (2)
    There is a weak homotopy equivalence
    $\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r}^{c})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1}^{c})\big)_{o}\overset{\simeq}{\longrightarrow}\Omega^{2}\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r+2}^{c})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{1}^{c})\big)_{o}.$
  3. (3)
    There is a weak homotopy equivalence
    $\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r}^{c})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1}^{c})\big)_{o}\overset{\simeq}{\longrightarrow}\Omega^{2}\big({\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r}^{c})/{\mathcal{G}}_{\infty}(A\otimes{\mathcal{C}}_{r-1}^{c})\big)_{o}$
  4. (4)
    Taking $A={\mathbb{C}}$ there is a weak homotopy equivalence
    $\mathbf{U}_{o}\overset{\simeq}{\longrightarrow}\Omega^{2}\mathbf{U}_{o}.$
Note that part (2) follows from part (1): you simply iterate (1) twice. For part (3) you use the ABS result, Theorem 19.2. Part (4) is simply the recognition that
${\mathcal{G}}_{\infty}({\mathcal{C}}_{1}^{c})/{\mathcal{G}}_{\infty}({\mathcal{C}}_{0}^{c})_{o}\cong\mathbf{U}_{o}.$
Our comment at the end of the previous section regarding the Homotopy Road Map applies here as well.
20. The Basics: Symmetric Spaces and a Timely Example
We rely on Helgason [41] and Gorodski [37] for our basic references. This is actually a vast topic and we are simply skimming the surface.
Definition 20.1.
Suppose that $(M,g)$ and $(N,h)$ are Riemann manifolds [^76] A Riemann metric on a smooth manifold is a smooth choice of a positive definite inner product for each tangent space of the manifold. A Riemann manifold is a smooth manifold together with a Riemann metric. and $f:M\to N$ is a diffeomorphism (that is, $f$ is a homeomorphism which is smooth and has a smooth inverse.) Then $f$ is an isometry if $f$ preserves distance. In shorthand, we may say $g=f^{*}h$, and more explicitly,
$g_{p}(u,v)=f_{f(p)}(df_{p}(u),df_{p}(v))$
for all $p\in M$ and $u,v\in T_{p}M$.
Definition 20.2.
Suppose that $M$ is a connected Riemann manifold and let $p\in M$. There is an exponential map $exp_{p}:T_{p}M\to M.$ We may choose a small open ball $B_{p}$ about the origin of $T_{p}M$ and label its image as $V_{p}\equiv exp_{p}(B_{p})\subset M.$ If $B_{p}$ is small enough then the map
$exp_{p}:B_{p}\longrightarrow V_{p}$
is a diffeomorphism. Define the geodesic symmetry at $p$, written $s_{p}:V_{p}\to V_{p}$ to be the composite
$V_{p}\overset{exp_{p}^{-1}}{\longrightarrow}B_{p}\overset{u\to-u}{\longrightarrow}B_{p}\overset{exp_{p}}{\longrightarrow}V_{p}.$
The key point is that $s_{p}$ reverses the direction of geodesics.
Definition 20.3.
A Riemann manifold $M$ is a symmetric space if for each $p\in M$ the map $s_{p}$ is a globally defined isometry of $M$. (Sometimes these are referred to as Riemannian symmetric spaces but we will leave off the first word.)
Here are some examples:
  1. (1)
    Euclidean space ${\mathbb{R}}^{n}$, with $s_{p}(y)=2p-y$
  2. (2)
    spheres $S^{n}$
  3. (3)
    projective spaces
  4. (4)
    compact Lie groups
In order to simplify matters and to concentrate immediately upon the cases that concern us, we will stick to compact symmetric spaces. We will also not discuss the exceptional Lie groups. These seem to not play a role in physics, as yet. [^77] For example, the lowest dimensional irreducible representation of the group $E_{8}$ is in dimension 248.
Proposition 20.4.
(cf. Helgason [41], §3-4. ) Every symmetric space arises as a homogeneous space of Lie groups.
For example, the symmetric space $S^{k}$ has isometry group $O_{k+1}$, its isometry group at the point $(1,0,0,\dots,0)$ is $O_{k}$ and hence we may write
$S^{k}\cong O_{k+1}/O_{k}.$
Note that had we used the connected component of the identity, then we would get the same outcome, since $SO_{k+1}/SO_{k}\cong S^{k}$.
We follow, more or less, the 1926 approach of Élie Cartan.
First, any irreducible symmetric space $M$ fits into one of three categories:
  1. (1)
    $M$ is isometric to some Euclidean space.
  2. (2)
    $M$ is compact. (It will then have nonnegative (but not identically zero) sectional curvature.)
  3. (3)
    $M$ is not compact and has nonpositive (but not identically zero) sectional curvature. This class is dual to the compact class.
As mentioned, we are restricting to the second category. The classification of É. Cartan uses heavily the Lie algebra associated to the Lie group. Here’s a quick introduction. We will focus completely on Lie algebras over either $\mathbb{R}$} or $\mathbb{C}$}, and we denote the field by $\mathbb{F}$}.
Definition 20.5.
A Lie algebra $\mathfrak{g}$ over a field $\mathbb{F}$} is an $\mathbb{F}$}-vector space together with a map
$\mathfrak{g}\times\mathfrak{g}\longrightarrow\mathfrak{g}$
such that
  1. (1)
    The map is bilinear.
  2. (2)
    $[x,x]=0$ for all $x\in\mathfrak{g}$.
  3. (3)
    The pairing satisfies the Jacobi identity. That is, for all $x,y,z\in\mathfrak{g}$,
    $[x,[y,z]]+[y,[z,x]]+[z,[x,y]]=0.$
For example, take $\mathfrak{g}=M_{k}({\mathbb{F}})$ with [A,B]=ABBA[A,B]=AB-BA.
The most important examples for us are the Lie algebras associated to particular Lie groups. These ideas date back to Sophus Lie (in the period 1869-1873), to his collaborators, particularly to Felix Klein, and to his students, particularly Élie Cartan.
Definition 20.6.
Fix a Lie group $G$. Then it is a smooth manifold, and so at each point $h\in G$ there is a tangent space $T_{h}G$. Suppose that $L_{g}:G\to G$ is a vector field on $G$ induced by $L_{g}(x)=gx$. We say that $L$ is a left invariant vector field if for all $g,h$,
$(dL_{g})_{h}:T_{h}G\longrightarrow T_{gh}G.$
We denote by $Lie(G)$ the real vector space of all left-invariant vector fields on $G$. It is a Lie algebra with Lie bracket defined by
$L_{g},L_{h}]=[L_{g}L_{h}-L_{h}L_{g}]$
There is special notation for the classical Lie groups. The Lie algebra of the group $SO_{n}$ is ${\mathfrak{so}}_{n}$ and similarly for $O_{n}$, $U_{n}$, $SU_{n}$ and $Sp_{n}$.
Now, what is the relationship between Lie groups and Lie algebras. Here are the facts:
Theorem 20.7.
(S. Lie)
  1. (1)
    If $G$ is a Lie group then $Lie(G)$ is a Lie algebra, with $dim(G)$ (as a compact smooth manifold) the same as $dim(Lie(G))$ (as a real vector space.)
  2. (2)
    If $\mathfrak{g}$ is a finite-dimensional Lie algebra, then there exists a connected Lie group $G$ with $Lie(G)\cong\mathfrak{g}$. Thus the map
    $\{Lie\,\,Groups\}\overset{Lie(-)}{\longrightarrow}\{Finite-Dimensional\,\,Lie\,\,Algebras\}$
    is onto.
  3. (3)
    Any two Lie groups with the same Lie algebra are locally isomorphic and, in particular, have the same universal cover.
  4. (4)
    If we restrict attention to simply connected797979i.e. with $\pi_{1}(G)=0$. Lie groups, then there is a 1-1 correspondence
    $\{Simply\,\,Connected\,\,Lie\,\,Groups\}\overset{Lie(-)}{\longrightarrow}\{Finite-Dimensional\,\,Lie\,\,Algebras\}.$
  5. (5)
    If $K$ is a closed Lie subgroup of the Lie group $G$ with associated Lie algebras $\mathfrak{k}$ and $\mathfrak{g}$, then $\mathfrak{g}$ may be written as a direct sum
    $\mathfrak{g}\cong\mathfrak{k}\oplus\mathfrak{m}$
    as vector spaces, with the commutator Lie bracket attending to this decomposition as follows: if $k,k_{1},k_{2}\in\mathfrak{k}$ and $m,m_{1},m_{2}\in\mathfrak{m}$, then
    $k_{1},k_{2}]\in\mathfrak{k}\qquad[k,m]\in\mathfrak{m}\qquad[m_{1},m_{2}]\in\mathfrak{k}.$
Theorem 20.8.
(E. Cartan) Ignoring the exceptional Lie groups, there are ten different families of compact symmetric spaces. They fall into three groups:
  1. (1)
    There are three compact connected Lie groups that appear in the classification as is: they are $SO_{n}$, $SU_{n}$, and $Sp_{n}$.
  2. (2)
    There are four symmetric spaces built in the $G/K$ format that are built out of two different groups. They are
    $SU_{n}/SO_{n}\qquad SU_{2n}/Sp_{n}\qquad SO_{2n}/U_{n}\qquad\qquad Sp_{n}/U_{n}.$
  3. (3)
    There are three symmetric spaces, also in the $G/K$ format, that have the form $G_{p+q}/(G_{p}\times G_{q})$
    $SO_{p+q}/(SO_{p}\times SO_{q}).\qquad SU_{p+q}/SU_{p}\times SU_{q})\qquad Sp_{p+q}/(Sp_{p}\times Sp_{q})$
These ten different categories of spaces correspond exactly to the ten categories of spaces we have seen, namely the eight spaces of the type $GL_{\infty}({\mathcal{C}}_{r})/GL_{\infty}({\mathcal{C}}_{r-1})$ associated to the real numbers and the two space types associated with the complex numbers!
Here’s an extension of the Homotopy Road Map demonstrating that.
Table 20.9.
The Homotopy Road Map: Step 2
label name finite approx Big Ten Bott ${\mathcal{R}}_{k}$ $\pi_{0}({\mathcal{R}}_{k})$
A unitary $_{n} 𝐔\mathbf{U}$ $\Omega^{0}\mathbf{U}$ ${\mathcal{R}}_{1}^{c}$ 0
AIII chiral unitary $U_{2n}/(U_{n}\times U_{n})$ $\mathbf{BU}\times{\mathbb{Z}}$ $\Omega^{1}\mathbf{U}$ ${\mathcal{R}}_{0}^{c}$ $\mathbb{Z}$}
D BdG $_{n} 𝐎\mathbf{O}$ $\Omega^{0}\mathbf{O}$ ${\mathcal{R}}_{1}$ ${\mathbb{Z}}_{2}$
DIII BdG $O_{2n}/U_{n}$ $\mathbf{O}/\mathbf{U}$ $\Omega^{1}\mathbf{O}$ ${\mathcal{R}}_{2}$ ${\mathbb{Z}}_{2}$
AII symplectic $U_{n}/Sp_{n}$ $\mathbf{U}/\mathbf{Sp}$ $\Omega^{2}\mathbf{O}$ ${\mathcal{R}}_{3}$ 0
CII chiral symplectic $Sp_{2n}/(Sp_{n}\times Sp_{n})$ $\mathbf{BSp}\times{\mathbb{Z}}$ $\Omega^{3}\mathbf{O}$ ${\mathcal{R}}_{4}$ $\mathbb{Z}$}
C BdG $_{n} 𝐒𝐩\mathbf{Sp}$ $\Omega^{4}\mathbf{O}$ ${\mathcal{R}}_{5}$ 0
CI BdG $Sp_{n}/U_{n}$ $\mathbf{Sp}/\mathbf{U}$ $\Omega^{5}\mathbf{O}$ ${\mathcal{R}}_{6}$ 0
AI orthogonal $U_{n}/O_{n}$ $\mathbf{U}/\mathbf{O}$ $\Omega^{6}\mathbf{O}$ ${\mathcal{R}}_{7}$ 0
BDI chiral orthogonal $O_{2n}/(O_{n}\times O_{n})$ $\mathbf{BO}\times{\mathbb{Z}}$ $\Omega^{7}\mathbf{O}$ ${\mathcal{R}}_{0}$ $\mathbb{Z}$}
A Timely Example
For the rest of this section we are going to consider a timely example, a case that was used very recently by Orion and Akkermans [66] to understand the entanglement of two qubits. They consider the case where the Lie group is $U_{4}$ and the closed subgroup is $O_{4}$ and then restrict to $SU_{4}$. So let’s look at the chart, featuring as a special case $U_{2t}/O_{2t}$. Here are some basic topological facts:
Proposition 20.10.
  1. (1)
    $U_{n}$ and $SU_{n}$ are compact smooth manifolds of dimension $n^{2}$ and $(n^{2}-1)$ respectively. Hence their Lie algebras $\mathfrak{u}_{n}$ and $\mathfrak{su}_{n}$ have dimension $n^{2}$ and $(n^{2}-1)$ respectively as real vector spaces.
  2. (2)
    $O_{n}$ and $SO_{n}$ are compact smooth manifolds, both of dimension $\frac{n(n-1)}{2}$. Their Lie algebras are isomorphic:
    $\mathfrak{o}_{n}\cong\mathfrak{so}_{n}$
    and have dimension $\frac{n(n-1)}{2}$ as real vector spaces.
  3. (3)
    In particular,
    $dim\,\mathfrak{su}_{4}=15\qquad\text{and}\qquad dim\,\mathfrak{o}_{4}=6.$
  4. (4)
    Using the Cartan decomposition, we may decompose
    $\mathfrak{su}_{4}\cong\mathfrak{o}_{4}\oplus\mathfrak{m}$
    as vector spaces, where $\dim\,\mathfrak{m}=9$. Under this decomposition, the Lie bracket has the following property:
    $h_{1},h_{2}]\in\mathfrak{o}_{4}\qquad\forall h_{j}\in\mathfrak{o}_{4}$
    $m_{1},m_{2}]\in\mathfrak{o}_{4}\qquad\forall m_{j}\in\mathfrak{m}$
    $h,m]\in\mathfrak{m}\qquad\forall h\in\mathfrak{o}_{4}\quad m\in\mathfrak{m}$
In order to move forward we need a some definitions known to every physicist but not, perhaps, to every mathematician. We will use physics notation for the remainder of this section and for the next. Beware!
Definition 20.11.
  1. (1)
    We denote the complex conjugate of a complex number $c$ by $c^{*}$ rather than $\bar{c}$.
  2. (2)
    If $A$ is a complex matrix then the Hermitian adjoint $A^{\dagger}$ of $A$ is defined by
    $a_{ij}]^{\dagger}=[a_{ji}^{*}]$
  3. (3)
    AA is said to be Hermitian (or self-adjoint) if $A^{\dagger}=A$ and $A$ is skew-Hermitian if $A^{\dagger}=-A$.
  4. (4)
    If $f:V\to W$ where $V$ and $W$ are complex vector spaces, then $f$ is said to be anti-linear or conjugate linear if for all $x,y\in V$ and $z\in{\mathbb{C}}$, we have
    $f(x+y)=f(x)+f(y)\qquad{\text{and}}\qquad f(zx)=\bar{z}f(x).$
  5. (5)
    If $U:\mathcal{H}\to\mathcal{H}$ where $\mathcal{H}$ is a complex Hilbert space, then $U$ is said to be anti-unitary if it is anti-linear and, in addition, if for all $x,y\in\mathcal{H}$,
    $Ux,Uy>=<x,y>^{*}.$
  6. (6)
    If $V$ is a complex vector space with basis elements $v_{j}$ then we define the complex conjugation operator
    $K:V\longrightarrow V$
    by $Kv_{j}=v_{j}$ and $K(iv_{j})=-iv_{j}$. Then one checks that
    $K^{\dagger}K=K^{2}=I.$
We need a basis for the vector space $\mathfrak{su}_{4}$ that respects the Cartan decomposition
$\mathfrak{su}_{4}\cong\mathfrak{o}_{4}\oplus\mathfrak{m}$
It is constructed as follows. Let
$\sigma_{1}=\left(\begin{matrix}0&1\\ 1&0\end{matrix}\right)\qquad\sigma_{2}=\left(\begin{matrix}0&-i\\ i&0\end{matrix}\right)\qquad\sigma_{3}=\left(\begin{matrix}1&0\\ 0&-1\end{matrix}\right)\qquad\sigma_{0}=\mathbb{I}=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right)$
denote the $2\times 2$ Pauli matrices in $M_{2}({\mathbb{C}})$.
Proposition 20.12.
The Pauli matrices satisfy the following properties:
  1. (1)
    Each matrix is Hermitian. [^80] The matrices $\sigma_{0},\sigma_{1},\sigma_{2},\sigma_{3}$ } form a basis for the Hermitian elements of $M_{2}({\mathbb{C}})$.
  2. (2)
    The Cayley table, their multiplication table, shows the values of the row times the column:
    $\times$ $\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$
    $\sigma_{1}$ $\mathbb{I}$ $i\sigma_{3}$ $i\sigma_{2}$
    $\sigma_{2}$ $i\sigma_{3}$ $\mathbb{I}$ $i\sigma_{1}$
    $\sigma_{3}$ $i\sigma_{2}$ $i\sigma_{1}$ $\mathbb{I}$
    (To read the table correctly, note that $\sigma_{1}\sigma_{2}=i\sigma_{3}$.)
  3. (3)
    In particular, for $i,j=1,2,3$ and $i\neq j$, then
    $\sigma_{i}\sigma_{j}=-\sigma_{j}\sigma_{i}.$
  4. (4)
    Complex conjugation is given by
    $\sigma_{1}^{*}=\sigma_{1}\qquad\sigma_{2}^{*}=-\sigma_{2}\qquad\sigma_{3}^{*}=\sigma_{3}$
    and
    $(i\sigma_{1})^{*}=-i\sigma_{1}\qquad(i\sigma_{2})^{*}=i\sigma_{2}\qquad(i\sigma_{3})^{*}=-i\sigma_{3}$
  5. (5)
    The matrices $i\sigma_{1},i\sigma_{2},i\sigma_{3}$ form a basis for the real Lie algebra $\mathfrak{su}_{2}$.
  6. (6)
    $det\,\sigma_{j}=-1\qquad\qquad tr\,\sigma_{j}=0\qquad\forall j$
    and hence each matrix $\sigma_{j}$ has eigenvalues +1+1 and 1-1.
Here is the basis for $\mathfrak{su}_{4}$ that respects the Cartan decomposition, as promised.
Proposition 20.13.
Regard
$M_{4}({\mathbb{C}})\cong M_{2}({\mathbb{C}})\otimes M_{2}({\mathbb{C}}).$
Then the Lie algebra $\mathfrak{o}_{4}$ is the span of the six basis elements
$i\sigma_{j}\otimes\mathbb{I},\quad i\mathbb{I}\otimes\sigma_{k}\qquad\qquad j,k\in\{1,2,3\}.$
The remaining space $\mathfrak{m}$ is the span of the nine basis elements
$i\sigma_{j}\otimes\sigma_{k}\qquad\qquad j,k\in\{1,2,3\}.$
We now can define the operators that we need.
Definition 20.14.
  1. (1)
    Let $K$ denote complex conjugation with respect to the basis we have constructed for $\mathfrak{su}_{4}$ as defined above. We have noted that $K^{2}=I$.
  2. (2)
    Define the antiunitary operator
    $\Theta:{\mathbb{C}}^{4}\longrightarrow{\mathbb{C}}^{4}$
    by
    $\Theta=(\sigma_{2}\otimes\sigma_{2})K.$
    Check that $\Theta^{2}=I$.
  3. (3)
    Define
    $Q=K(\sigma_{1}\otimes i\sigma_{2}).$
    Check once again that $Q^{2}=-I$.
  4. (4)
    . Finally, define $T$ by
    $T(A)=\Theta A\Theta^{\dagger}.$
    Then $T$ is a map
    $T:\mathfrak{su}_{4}\longrightarrow\mathfrak{su}_{4}$
    and $T^{2}=I$.
Theorem 20.15.
Let $T:\mathfrak{su}_{4}\longrightarrow\mathfrak{su}_{4}$ as described above. Then the decomposition $\mathfrak{su}_{4}\cong\mathfrak{o_{4}}\oplus\mathfrak{m}$ coincides with the eigenspace decomposition of $T$. That is:
  1. (1)
    $T(V)=V\quad\quad{\text{if}}\quad V\in\mathfrak{o_{4}}$.
  2. (2)
    $T(V)=-V\quad{\text{if}}\quad V\in\mathfrak{m}.$
We omit the proof and refer the reader to [66]
20.1. Topology and Physics, by Nadav Orion
In quantum mechanics, the physical description of a system consists of two notions:
  1. (1)
    The state of the system. It contains the information about the system as it is in a certain moment in time. It is described by a vector in a complex inner product space (specifically a Hilbert space), the state space, which in turn contains all possible system states. These vector spaces may be of any complex dimension, from two (such a system is called “qubit”) to infinity.
  2. (2)
    The Hamiltonian $H$. It is a Hermitian operator on that space of physical states that determines how a system evolves in time. $H$ generates the unitary evolution operator
    $e^{-iHt/\hbar}$
    where $t$ is the time and $\hbar$ is the reduced Planck constant. If a system was described at $t=0$ by the vector $v$, then at time $t$ the state of the system is
    $e^{-iHt/\hbar}v.$
    For a finite interval $t\in\left[0,T\right]$, the equation above describes a path in the state space.
In a lab setting, physicists control certain aspects of the Hamiltonian (e.g. by applying electromagnetic fields), while other aspects depend on the specific fabrication and noise and thus change from one experiment to another. For these reasons a set of Hamiltonians, rather than a single one, is usually considered. In many cases this set corresponds to a Lie algebra, Clifford algebra, or a well-defined subset of one of these.
Different elements in the set of Hamiltonians may produce different paths in the state space. These paths are at the basis of many physical phenomena such as
  1. (1)
    Aharonov-Bohm effect,
  2. (2)
    Berry phase,
  3. (3)
    quantum Hall effect, and
  4. (4)
    entanglement
to name a few. These are captured as topological properties of Hamiltonians as maps (or a set of maps) from the space of physical parameters, both controlled and uncontrolled, to the set of Hamiltonians. The space of Hamiltonians is defined by applying additional equivalence relations, depending upon the situation. The topological properties of the evolution operator space (defined using the set of Hamiltonians) can also be examined, and they are related to those of the space of Hamiltonians. This relation is exactly Bott periodicity when the state space dimension $\rightarrow\infty$, but is very similar for large enough dimensions.
In “A Timely Example” two qubits are considered. For each qubit the state space is $\mathbb{C}^{2}$, giving a state space of
$\mathbb{C}^{2}\otimes\mathbb{C}^{2}\cong\mathbb{C}^{4}$
for the joint system (both spaces with the usual inner product). A typical Hamiltonian for such a system contains both terms that affect a single qubit ($\sigma_{j}\otimes I$ and $I\otimes\sigma_{k}$) and terms that describe interactions between the two ($\sigma_{j}\otimes\sigma_{k}$), such that the set of Hamiltonians is all of the Hermitian $4\times 4$ matrices, which is the Lie algebra $\mathfrak{u}(4)$. The evolution operator space is thus $U_{4}$, a well-known compact complex symmetric space.
In some physical cases not all terms are allowed, for example when the system possess time reversal symmetry, i.e.
$T\left(H\right)\equiv\Theta H\Theta^{\dagger}=H.$
For two qubit systems the natural time reversal symmetry is
$\Theta=\left(\sigma_{2}\otimes\sigma_{2}\right)K,$
thus
$H\in\text{span}\left\{\sigma_{j}\otimes\sigma_{k}\right\}_{j,k\in\left\{1,2,3\right\}},$
which is (up to a factor of $i$) the set $\mathfrak{m}$ defined in proposition 20.10. The disallowed terms are the Lie algebra $\mathfrak{o}_{4}$. To get the space of evolution operators one must then identify all operators in the original $U_{4}$ space that differ by the disallowed terms $O_{4}$, giving the homogeneous space
$U_{4}/O_{4},$
a known real symmetric space.
Entanglement is a very valuable resource in physics, and two qubits is the simplest case in which it can be observed. The Cartan decomposition distinguishes between Hamiltonians that may create and destroy entanglement and those that cannot interact with it, and as a consequence also distinguishes between paths where entanglement changes and paths where it does not. The evolution operators in $U_{4}/O_{4}$ can change the entanglement and those in $O_{4}$ cannot. Furthermore, the different Hamiltonians in $\mathfrak{m}$ generate paths in $U_{4}/O_{4}$ that differ by the amount of entanglement they can create.
One of the equivalence relations used to define the space of Hamiltonians is homotopic equivalence. In this case the topological properties of the Hamiltonian do not depend on the size of its eigenvalues, so one may then ”flatten” $H$ by deforming its eigenvalues to be +1+1 or 1-1 (or 0 in extremely special cases). This leads to the associated evolution operator paths mentioned being loops. [^81] The loop starts at $t=0$. To define the number for the Hamiltonian, we stop the first time the loop returns to its starting place. These loops then are maps
$S^{1}\longrightarrow U_{4}/O_{4}$
and their homotopy classes lie in the group $\pi_{1}(U_{4}/O_{4})$, which is isomorphic to $\mathbb{Z}$ as we have shown in Section 12. Thus to each time evolution is associated an integer, which is a topological invariant of that path. By homotopy theory, if a single loop of a given Hamiltonian is given the invariant 1, performing the loop again will give 2. We can define the topological number of the Hamiltonian by the number given by a single loop.
21. Quantum Symmetry
Suppose that $G$ is a compact Lie group and $K$ is a closed sub-Lie group in $G$. Then the Lie algebra of $G$ decomposes as real vector spaces
$Lie(G)\cong Lie(K)\oplus\mathfrak{m}$
as noted previously, and the decomposition preserves the Lie algebra structure in $Lie(K)$. However, $\mathfrak{m}$ is not a Lie algebra as we saw in the Timely Example.
Recall that $G$ is in fact a compact smooth manifold with a closed submanifold $K$ and homogeneous space $G/K$. Topologically, $Lie(G)$ corresponds to the tangent space $T_{o}G$ of $G$ at the origin, $Lie(K)$ corresponds to the subspace $T_{o}K$ that represents the tangent space of $K$ at the origin. Taking quotients, we see that $\mathfrak{m}$ corresponds to the tangent space of $G/K$ at the point on the smooth manifold $G/K$ that is the image of the origin of $G$. Thus symmetries of $\mathfrak{m}$ correspond to symmetries of the tangent space of $G/K$ at the origin.
Next we provide a very simplified introduction to the T,C,ST,\,C,\,S symmetries so beloved by followers of the Ten-Fold Way. The assignment of these symmetries to the ten Cartan classes dates to 1996 and is attributed to Altland and Zirnbauer (cf. [3], [90]). [^82] We note that the whole TCS assignment is considered somewhat tentative by some physicists- Emil Prodan alerted us to this matter and has an alternate approach which is too involved to be introduced in this work.
We start with a quantum system with some real Hamiltonian $H$ and seek topological invariants. To say they are topological means, roughly, that they pay no attention to gentle continuous perturbations of the Hamiltonian. For the Hamilton to have a topological invariant it must have energy gaps, since if not then every Hamiltonian can be deformed to any other.
We seek the simplest examples. If the Hamiltonian is invariant under a unitary operator then it may be block-diagonalized with respect to the operator into smaller blocks. We do this until there are no further unitary operators, so that it is irreducible.
There are, however, other possibilities, the first of which being time-reversal symmetry $T$. Possessing this symmetry means that the system would look the same, even at the atomic level, if time were reversed. So in general we can write time-reversal symmetry as $T$ with $T^{2}=I$ or $T^{2}=-I$ and the custom is to write 0 if it is not present. The Timely Example was indeed an example of time-reversal symmetry with $T^{2}=I$.
The second sort of symmetry that might be present is charge conjugation symmetry, also known as particle-hole symmetry, where we, at least symbolically, switch particles and holes. This is modeled by an antiunitary operator $C$ with $C^{2}=1$ or $C^{2}=-1$ and we write 0 if it is not present.
Finally, it is possible to have a system that has neither time reversal nor charge conjugation symmetry, but it’s symmetrical under the combination of time reversal and charge conjugation. We also may have $S=CT$ with $S^{2}=1$ .
Finally, we complete the Homotopy Road Map.
Table 21.1.
The Complete Homotopy Road Map
label Name T C S BigTenBig\,\,Ten $\mathcal{R}$$_{?}$ $\pi_{0}({\mathcal{R}}_{j})$
A unitary 0 0 0 $\mathbf{U}\cong\Omega^{0}\mathbf{U}$ ${\mathcal{R}}_{1}^{c}$ 0
AIII ch. unitary 0 0 +1 $\mathbf{BU}\times{\mathbb{Z}}\,\,\,\cong\,\,\,\Omega^{1}\mathbf{U}$ ${\mathcal{R}}_{0}^{c}$ $\mathbb{Z}$}
D BdG 0 +1 0 $\mathbf{O}\,\,\,\cong\,\,\,\Omega^{0}\mathbf{O}$ ${\mathcal{R}}_{1}$ ${\mathbb{Z}}_{2}$
DIII BdG -1 +1 +1 $\mathbf{O}/\mathbf{U}\,\,\,\cong\,\,\,\Omega^{1}\mathbf{O}$ ${\mathcal{R}}_{2}$ ${\mathbb{Z}}_{2}$
AII symplectic -1 0 0 $\mathbf{U}/\mathbf{Sp}\,\,\,\cong\,\,\,\Omega^{2}\mathbf{O}$ ${\mathcal{R}}_{3}$ 0
CII ch. symplectic -1 -1 +1 $B\mathbf{Sp}\times{\mathbb{Z}}\,\,\,\cong\,\,\,\Omega^{3}\mathbf{O}$ ${\mathcal{R}}_{4}$ $\mathbb{Z}$}
C BdG 0 -1 0 $\mathbf{Sp}\,\,\,\cong\,\,\,\Omega^{4}\mathbf{O}$ ${\mathcal{R}}_{5}$ 0
CI BdG +1 -1 +1 $\mathbf{Sp}/\mathbf{U}\,\,\,\cong\,\,\,\Omega^{5}\mathbf{O}$ ${\mathcal{R}}_{6}$ 0
AI orthogonal +1 0 0 $\mathbf{U}/\mathbf{O}\,\,\,\cong\,\,\,\Omega^{6}\mathbf{O}$ ${\mathcal{R}}_{7}$ 0
BD1 chiral orth +1 +1 +1 $\mathbf{BO}\times{\mathbb{Z}}\,\,\,\cong\,\,\,\Omega^{7}\mathbf{O}$ ${\mathcal{R}}_{0}$ $\mathbb{Z}$}
Each of the ten rows of this chart corresponds to one of the ten types of compact symmetric spaces and hence to the Ten-Fold way. We have discussed each column previously; here is a reminder.
  1. (1)
    label gives the É. Cartan indicator for this class of symmetric spaces.
  2. (2)
    Name gives the Cartan name for this class of symmetric spaces.
  3. (3)
    $\bf{T,C,S}$ indicate the types of symmetry present on this class of symmetric spaces. For example, the ”orthogonal” row, second row from the bottom, includes the Timely Example, in which time-reversal symmetry $T$ was present with $T^{2}=+1$ and $C$ and $S$ were not present.
  4. (4)
    Big Ten This shows the stable version of this class of spaces. For example, look at the ”orthogonal” row, second from the bottom. The space $\mathbf{U}/\mathbf{O}$ is the union of symmetric spaces $U_{n}/O_{n}$. The space $\mathbf{U}/\mathbf{O}$ came to our attention first via the isomorphism
    $\mathbf{U}/\mathbf{O}\cong\Omega^{6}\mathbf{O}$
    as part of our exposition of Bott Periodicity. We named this space ${\mathcal{R}}_{7}$ following physics custom. It arose a second time via the Lie algebra perspective in the context of Clifford algebras. It comes up a third time via $U_{4}/O_{4}$ in the Timely Example of the previous section.
  5. (5)
    $\bf{\pi_{0}({\mathcal{R}}_{j})}$} shows the lowest homotopy group of one of the Big Ten spaces. In the case of $\mathbf{U}/\mathbf{O}$ the group is 0, indicating that the space is path-connected.
This is laid out in Stone et al [81], §3 in very concrete detail. They demonstrate there that each of the symmetric spaces ${\mathcal{R}}_{j}$ listed above has certain symmetries associated to it, or, more precisely, associated to $\mathfrak{m}_{j}$. Their Table 4 lists which symmetry occurs in each of the eight real cases. These symmetries are vital for the applications in physics.
Since
$\widetilde{KO}^{-n}(X)\cong[X,{\mathcal{R}}_{n}]$
and
$\widetilde{K}^{-n}(X)\cong[X,{\mathcal{R}}^{c}_{n}],$
we see at once that $KO$-theory and $K$-theory are intimately related to this story. The homotopy properties of the ${\mathcal{R}}_{n}$ are detected by $KO$ and hence connected to real vector bundles. This allows topological properties of the various symmetries to be detected and utilized. Much current physics research in topological insulators and related matters uses $KO^{*}$, and this was one of the main motivations for writing this work .
It is important to remember that the ${\mathcal{R}}_{j}$ are unions of compact symmetric spaces, as this structure helps us very much in understand the homotopy side of the story. For example,
${\mathcal{R}}_{2}=\mathbf{O}/\mathbf{U}=\lim O_{16t}/U_{8t}$
and the map
$\pi_{j}(O_{16t}/U_{8t})\to\pi_{j}(\mathbf{O}/\mathbf{U})=\pi_{j}({\mathcal{R}}_{2})$
is an isomorphism for $j<<t$ by cellular approximation. Thus, for example, if $X$ is some compact space then
$\widetilde{KO}^{-2}(X)\cong[X,{\mathcal{R}}_{2}]=[X,O_{16t}/U_{8t}]$
for $dim(X)<<t$.
For many of the applications we need $\pi_{j}$ of the finite approximations (e.g. $\pi_{j}(O_{4}/U_{4})$ for $j=0,1,2$,) and these may not coincide with $\pi_{j}(\mathbf{O}/\mathbf{U})$ so one has to be careful. See Chapter 12 for a deep discussion of this matter.
Part IV A Deeper look at the Structure of $KO$ and $K$
22. Real and Complex $K$-theory
This section is concerned with four inter-related matters:
  1. (1)
    The structural maps (aka natural transformations) that connect $KO_{*}(A)$ and $K_{*}(A_{\mathbb{C}})$ ;
  2. (2)
    General KOKO-$K$ relationships;
  3. (3)
    The product structure on the ring $KO_{*}({\mathbb{R}})$.
  4. (4)
    A spectral sequence for filtered algebras that arises both in the real and complex cases, with a natural transformation connecting them.
For most of this information one should refer to Karoubi [52], Bousfield [26], Boersema [15].
For a real $C^{*}$-algebra $A$, there are five important natural transformations
$c_{n}$ $\colon KO_{n}(A)\rightarrow K_{n}(A_{\mathbb{C}})$
$r_{n}$ $\colon K_{n}(A_{\mathbb{C}})\rightarrow KO_{n}(A)$
$\psi_{n}$ $\colon K_{n}(A_{\mathbb{C}})\rightarrow K_{n}(A_{\mathbb{C}})$
$\beta$ $\colon K_{n}(A_{\mathbb{C}})\rightarrow K_{n+2}(A_{\mathbb{C}})$
$\beta_{0}$ $\colon KO_{n}(A)\rightarrow KO_{n+8}(A)$
which are defined as follows:
  1. (1)
    The map $c\colon KO_{*}(A)\rightarrow K_{*}(A_{\mathbb{C}})$ is defined as in Section 6.
  2. (2)
    There is also a standard inclusion ${\mathbb{C}}\rightarrow M_{2}({\mathbb{R}})$ given by
    $x+iy\mapsto\left(\begin{matrix}x&-y\\ y&x\end{matrix}\right).$
    Tensoring by $A$ we obtain the homomorphism $A_{\mathbb{C}}\rightarrow M_{2}(A)$, which induces
    ${\mathcal{P}}_{\infty}({\mathbb{C}})\to{\mathcal{P}}_{\infty}(M_{2}({\mathbb{R}}))\cong{\mathcal{P}}_{\infty}({\mathbb{R}})$
    and hence, for each $n$,
    $r_{n}\colon K_{n}(A_{\mathbb{C}})\rightarrow KO_{n}(M_{2}(A))\cong KO_{n}(A).$
  3. (3)
    The conjugate-linear map $A_{\mathbb{C}}\rightarrow A_{\mathbb{C}}$ defined by $a+ib\mapsto a-ib$ induces the map $\psi$.
  4. (4)
    The map $\beta$ is the complex Bott periodicity isomorphism.
  5. (5)
    The map $\beta_{0}$ is the real Bott periodicity isomorphism.
For example, if $A={\mathbb{R}}$ then we have $KO_{0}({\mathbb{R}})={\mathbb{Z}}$ and $K_{0}({\mathbb{R}}_{\mathbb{C}})=K_{0}({\mathbb{C}})={\mathbb{Z}}$. As both groups are generated by the class of the projection $1\in{\mathbb{R}}\subset{\mathbb{C}}$, it follows that $c_{0}\colon KO_{0}({\mathbb{R}})\rightarrow K_{0}({\mathbb{R}}_{\mathbb{C}})$ is an isomorphism. As the identity $1\in{\mathbb{C}}$ corresponds to the identity $1_{2}$ under the inclusion ${\mathbb{C}}\subset M_{2}({\mathbb{R}})$, the map $r_{0}\colon K_{0}({\mathbb{R}}_{\mathbb{C}})\rightarrow KO_{0}({\mathbb{R}})$ is multiplication by 2.
If $A={\mathbb{H}}$, we also have $KO_{0}({\mathbb{H}})=K_{0}({\mathbb{H}}_{\mathbb{C}})={\mathbb{Z}}$. However
$c_{0}\colon KO_{0}({\mathbb{H}})\rightarrow K_{0}({\mathbb{H}}_{\mathbb{C}})$
is multiplication by 2. Indeed, $KO_{0}({\mathbb{H}})$ is generated by the class of the identity projection $1_{\mathbb{H}}\in{\mathbb{H}}$. But under the standard inclusion ${\mathbb{H}}\subset{\mathbb{H}}_{\mathbb{C}}=M_{2}({\mathbb{C}})$, the identity of $\mathbb{H}$} corresponds to the projection $1_{2}\in M_{2}({\mathbb{C}})$ and $1_{2}]$ is twice the generator of $K_{0}({\mathbb{C}})$. Then it follows from the following proposition that
$r_{0}\colon K_{0}({\mathbb{H}}_{\mathbb{C}})\rightarrow KO_{0}({\mathbb{H}})$
is an isomorphism.
Proposition 22.1.
The natural transformations defined above satisfy the following relations:
$rc=2,\qquad cr=1+\psi,\qquad\psi^{2}=1,\qquad\psi c=c,\qquad r\psi=r.$
Proof.
  1. (1)
    The composition $A\rightarrow A_{\mathbb{C}}\rightarrow M_{2}(A)$ satisfies
    $p\mapsto\left(\begin{smallmatrix}p&0\\ 0&p\end{smallmatrix}\right),$
    and from this it follows that $rc([p])=2[p]$ for any projection in $\mathcal{P}_{\infty}(A)$.
  2. (2)
    The composition $A_{\mathbb{C}}\rightarrow M_{2}(A)\rightarrow M_{2}(A_{\mathbb{C}})$ satisfies
    $a+ib\mapsto\left(\begin{matrix}a&-b\\ b&a\end{matrix}\right)$
    for any projection $p=a+ib$. Conjugating the latter by the unitary
    $u=\tfrac{1}{\sqrt{2}}\left(\begin{matrix}1&i\\ i&1\end{matrix}\right)$
    yields
    $\left(\begin{matrix}a+ib&0\\ 0&a-ib\end{matrix}\right).$
    Therefore
    $cr([p])=(1+\psi)([p])$
    for any projection.
  3. (3)
    The relation $\psi^{2}=1$ follows directly from the fact that $a+ib\mapsto a-ib$ is an involution on $A_{\mathbb{C}}$.
  4. (4)
    It is easy to check the relation $\psi c=c$ holds exactly on the level of $C^{*}$-algebra maps.
  5. (5)
    Finally to verify the relation $r\psi=r$, note that if $p=a+ib$ is a projection in $A_{\mathbb{C}}$ (or in $M_{n}(A_{\mathbb{C}})$ for some $n$) then
    $r\psi(p)=\left(\begin{smallmatrix}a&b\\ -b&a\end{smallmatrix}\right).$
    But this projection is unitarily equivalent to $\left(\begin{smallmatrix}a&-b\\ b&a\end{smallmatrix}\right)$, by conjugating by $u=\left(\begin{smallmatrix}-1&0\\ 0&-1\end{smallmatrix}\right)$. It follows that $r\psi[p]=r[p]$.
We now discuss the ring structure of $KO_{*}({\mathbb{R}})$. Recall from Section 8 that $KO_{*}({\mathbb{R}})$ has period 8 and that in degrees 0 through $7$,
n= 0 11 22 33 44 55 66 77
$KO_{n}({\mathbb{R}})$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} 0 0 0
We specify generators for these groups as follows:
  1. (1)
    The generator of $KO_{0}({\mathbb{R}})={\mathbb{Z}}$ is the class of the unit $[1]$, and it is easy to check that
    $1]\cdot[p]=\mu([1]\otimes[p])=[p]{\text{\qquad for any\quad}}p\in\mathcal{P}_{k}(A).$
    In particular, $[1]$ is a unit in the ring $KO_{*}({\mathbb{R}})$ and serves as the identity under multiplication.
  2. (2)
    Let $\eta\in KO_{1}({\mathbb{R}})$ be the non-zero element. Then it turns out that $\eta^{2}$ is also the non-zero element of $KO_{2}({\mathbb{R}})$.
  3. (3)
    We take $\xi$ to be a generator of $KO_{4}({\mathbb{R}})={\mathbb{Z}}$.
  4. (4)
    The Bott element $\beta_{O}$ is a generator of $KO_{8}({\mathbb{R}})={\mathbb{Z}}$. It turns out that $\xi^{2}=4\beta_{O}$.
Theorem 22.2.
(Karoubi [52], III.5.19 )
$KO_{*}({\mathbb{R}})$ is the graded commutative ring with generators
  • $\eta$ in degree 1,
  • $\xi$ in degree 3,
  • $\beta_{O}$ in degree 8,
  • $\beta_{O}^{-1}$ in degree -8,
subject to the relations:
  • $2\eta=0$,
  • $\eta^{3}=0,$
  • $\xi^{2}=4\beta_{O},$
  • $\beta_{O}\beta_{O}^{-1}=1$
That is, $KO_{*}({\mathbb{R}})$ is the quotient of the graded polynomial ring over the integers generated by $\eta,\xi,\beta_{0},$ and $\beta_{0}^{-1}$ modulo the ideal generated by the relations specified above.
Since $KO_{*}(A)$ is a module over $KO_{*}({\mathbb{R}})$, the generators listed above induce natural transformations on $KO_{*}(A)$ induced by multiplication by these elements of $KO_{*}({\mathbb{R}})$. Thus we have
$\eta_{n}$ $\colon KO_{n}(A)\rightarrow KO_{n+1}(A)$
$\xi_{n}$ $\colon KO_{n}(A)\rightarrow KO_{n+4}(A)$
$\beta_{O}$ $\colon KO_{n}(A)\rightarrow KO_{n+8}(A)\;,$
and these natural transformations are subject to the same relations listed in Theorem 22.2.
Proposition 22.3.
For any real $C^{*}$-algebra $A$, there is a long exact sequence
$\cdots\rightarrow KO_{n}(A)\xrightarrow{~\eta~}KO_{n+1}(A)\xrightarrow{~c~}K_{n+1}(A_{\mathbb{C}})\xrightarrow{~r\beta^{-1}}KO_{n-1}(A)\rightarrow\cdots$ (3)
Proof.
We start with the inclusion homomorphism $c\colon{\mathbb{R}}\rightarrow{\mathbb{C}}$. Then the mapping cone construction from Section 4 gives a fibration
$0\rightarrow S{\mathbb{C}}\hookrightarrow Cc\rightarrow{\mathbb{R}}\rightarrow 0$
which gives rise to a long exact sequence
$\cdots\longrightarrow KO_{n+1}(A\otimes{\mathbb{C}})\rightarrow KO_{n}(A\otimes Cc)\rightarrow KO_{n}(A)\xrightarrow{\partial}KO_{n}(A\otimes{\mathbb{C}})\rightarrow\cdots\,.$
We know that $\partial=c*$ and it still remains to identify the other maps. For that we refer to either Section 1.11 of Schröder [76] or Theorem 1.18 of Boersema [15] for details. ∎
Corollary 22.4.
Suppose that $f:A\to B$ is a homomorphism of real $C^{*}$-algebras. Then the following are equivalent:
  1. (1)
    The map
    $f_{*}:KO_{n}(A)\longrightarrow KO_{n}(B)$
    is an isomorphism for all $n$.
  2. (2)
    The map
    $f_{*}:K_{n}(A_{\mathbb{C}})\longrightarrow K_{n}(B_{\mathbb{C}})$
    is an isomorphism for all $n$.
Proof.
Consider the map $f_{*}$ applied to the short exact sequence of Equation (3).
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${KO_{n-1}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$$\scriptstyle{f_{*}}$${KO_{n}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c}$$\scriptstyle{f_{*}}$${K_{n}(A_{\mathbb{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r\beta^{-1}}$$\scriptstyle{f_{*}}$${KO_{n-2}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$$\scriptstyle{f_{*}}$${KO_{n-1}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}\textstyle{\cdots}\textstyle$$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${KO_{n-1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$${KO_{n}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c}$${K_{n}(B_{\mathbb{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r\beta^{-1}}$${KO_{n-2}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$${KO_{n-1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots}$
If $f_{*}\colon KO_{n}(A)\rightarrow KO_{n}(B)$ is an isomorphism for all $n$, then the five-lemma implies that $f_{*}\colon K_{n}(A_{\mathbb{C}})\rightarrow K_{n}(B_{\mathbb{C}})$ is an isomorphism for all $n$. For the converse, we refer to Section 2.3 of Bousfield [26]. ∎
Corollary 22.5.
Suppose that $A$ is a real $C^{*}$-algebra. Then the following are equivalent:
  1. (1)
    $KO_{n}(A)=0$ for all $n$.
  2. (2)
    $K_{n}(A_{\mathbb{C}})=0$ for all $n$.
Proof.
If $KO_{n}(A)=0$ for all $n$, then the long exact sequence (3) immediately implies that $K_{n}(A_{\mathbb{C}})=0$ for all $n$. Conversely, suppose that $K_{n}(A_{\mathbb{C}})=0$ for all $n$, then the long exact sequence (3) implies that $\eta_{n}$ is an isomorphism for all $n$. However, the relation $\eta^{3}=0$ then implies that $KO_{n}(A)=0$ for all $n$. ∎
Finally, there is a deep result in spectral sequences.
Definition 22.6.
Suppose that $A$ is a real or a complex $C^{*}$-algebra. We say that $A$ is filtered if there is an increasing sequence of closed ideals
$A_{0}\subset A_{1}\subset A_{2}\subset\dots$
with $\overline{\cup_{j}A_{j}}=A$.
Filtered $C^{*}$-algebras occur naturally in several situations, typically in the structure theory of certain types of $C^{*}$-algebras. The following result was established in [72] Theorem 1.2 for complex $C^{*}$-algebras and the proof there is equally valid in the real context of real $C^{*}$-algebras.
Theorem 22.7.
([72], Thm. 1.2 )
  1. (1)
    Let $A$ be a real filtered $C^{*}$-algebra. Then there is a spectral sequence $\{E^{r},d^{r}\}$ with $d^{r}:E_{p,q}^{r}\to E_{p-r,q+r-1}^{r}$ which converges to the graded object associated to the natural filtration on
[…]
    with
    $E_{p,q}^{1}(A)=KO_{p+q}(A_{p}/A_{p-1}),$
    where $KO_{p+q}$ is graded mod 8.
  2. (2)
    If $A$ is a complex filtered $C^{*}$-algebra then there is an analogous $K_{*}$-spectral sequence. with $d^{r}:E_{p,q}^{r}\to E_{p-r,q+r-1}^{r}$ which converges to the graded object associated to the natural filtration on
[…]
    with
    $E_{p,q}^{1}(A)=K_{p+q}(A_{p}/A_{p-1}),$
    where $K_{p+q}$ is graded mod 2.
  3. (3)
    If $A$ is a real $C^{*}$-algebra with complexification $A\overset{c}{\longrightarrow}A_{\mathbb{C}}$, then the map $c$ induces a natural transformation of spectral sequences
    $E_{p,q}^{r}(A)\overset{c}{\longrightarrow}E_{p,q}^{r}(A_{\mathbb{C}}).$
A special case of the above result, established much earlier by Atiyah and Hirzebruch [10], the very first paper on topological $K$-theory, is the Atiyah-Hirzebruch spectral sequence. We state it for finite complexes $X$, though it holds for compact spaces and, with a little modification, for locally compact spaces.
Corollary 22.8.
Suppose that $X$ is a finite simplicial complex838383imagine a CW-complex with finitely many cells attached with skeleta [^84] For CW complexes, let $X_{0}$ be the finite set of points you start with, $X_{1}$ is what you have after adding all of the one-dimensional cells, etc. $X_{j}$. Let $\{x_{o}\}$ denote the space with a single point. Then there is a spectral sequence $E_{r}^{p,q}$ ($r\geq 1$, $\infty<p,q<+\infty$), with
$E_{1}^{p,q}\cong C^{p}(X;K^{q}(\{x_{o}\})$
($p$- cochains taking values in $K^{q}(\{x_{o}\}$ ), with $d^{1}$ usual coboundary operator, and
$E_{2}^{p,q}\cong H^{p}(X;K^{q}(\{x_{o}\})$
$E_{\infty}^{p,q}\cong G_{p}K^{p+q}(X).$
where $G_{p}K^{p+q}(X)$ is the associated graded group built from the skeletal filtration of $K^{*}(X)$ built from the filtration of $X$. The differential
$d_{r}:E_{r}^{p,q}\longrightarrow E_{r}^{p+r,q-r+1}$
vanishes for $r$ even.
23. Examples and Calculations
In this section, we present a number of different examples of the results of calculations of $K$-theory for real $C^{*}$-algebras and topological spaces with involution. Because of the structure discussed in the previous section, it is convenient to consider both the “real” and “complex” $K$-theory together. That is, for a real $C^{*}$-algebra $A$ we will often consider both $KO_{*}(A)$ and $K_{*}(A_{\mathbb{C}})$; and for a compact space $X$ we will consider both
$KO^{-*}(X)\cong KO_{*}(\real(X))\qquad{\text{and}}\qquad K^{-*}(X)\cong K_{*}(C^{{\mathbb{C}}}(X)).$
These results are found in Schroeder [76] and in Boersema [15].
Example 23.1.
As discussed in the previous section, we have (with $\{x\}$ designating the one-point space),
n= 0 11 22 33 44 55 66 77
$KO^{-n}(\{x\})=KO_{n}({\mathbb{R}})$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} 0 0 0
$K^{-n}(\{x\})=K_{n}({\mathbb{R}}_{\mathbb{C}})$ $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0
$c$ 1 0 0 0 2 0 0 0
$r$ 2 0 1 0 1 0 0 0
$\eta$ 1 1 0 0 0 0 0 0
Example 23.2.
We can consider $\mathbb{C}$} as a real $C^{*}$-algebra. When we do that we find there is an isomorphism
$\mathbb{C}$$_{{\mathbb{C}}}\cong{\mathbb{C}}\otimes_{{\mathbb{R}}}$$\mathbb{C}$$\cong{\mathbb{C}}\oplus{\mathbb{C}}$.
Thus $K_{*}({\mathbb{C}}_{\mathbb{C}})=K_{*}({\mathbb{C}}\oplus{\mathbb{C}})$, so we have the following:
n= 0 11 22 33 44 55 66 77
$KO_{n}({\mathbb{C}})$ $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0
$K_{n}({\mathbb{C}}_{\mathbb{C}})$ ${\mathbb{Z}}^{2}$ 0 ${\mathbb{Z}}^{2}$ 0 ${\mathbb{Z}}^{2}$ 0 ${\mathbb{Z}}^{2}$ 0
$c$ $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0
$r$ $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0
$\eta$ 0 0 0 0 0 0 0 0
Example 23.3.
The suspension algebra is $S{\mathbb{R}}=\real_{0}({\bf{R}})$. Since
$KO_{n}(S{\mathbb{R}})=KO_{n+1}({\mathbb{R}})\,\,\,\,\text{and}\,\,\,\,K_{n}(S{\mathbb{C}})=K_{n+1}({\mathbb{C}})$
we have
n= 0 11 22 33 44 55 66 77
$KO^{-n}({\bf{R}})\cong KO_{n}(\real_{0}({\bf{R}}))$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} 0 0 0 $\mathbb{Z}$}
$K^{-n}({\bf{R}})\cong K_{n}(C^{{\mathbb{C}}}_{0}({\bf{R}}))$ 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$}
Example 23.4.
The desuspension algebra is $V{\mathbb{R}}=\real_{0}(i{\bf{R}})$. Since
$KO_{n}(V{\mathbb{R}})=KO_{n-1}({\mathbb{R}})\,\,\,\,\text{but}\,\,\,\,V{\mathbb{R}}_{\mathbb{C}}\simeq S{\mathbb{R}}_{\mathbb{C}},$
we have
n= 0 11 22 33 44 55 66 77
$KO_{n}(C_{o}^{{\mathbb{R}}}(i{\bf{R}}))$ 0 $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} 0 0
$K_{n}(C_{o}^{{\mathbb{C}}}(i{\bf{R}}))$ 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$} 0 $\mathbb{Z}$}
Compare the first row with the first row of the previous (suspension) example to see the difference in how the dimension shifts.
Definition 23.5.
If $M$ is a graded group (indexed by $\mathbb{Z}$}), we let $\Sigma^{k}M$ be the $k$th suspension of $M$, defined by
$(\Sigma^{k}M)_{i}=M_{i+k}\;.$
With this notation, Examples 17.3 and 17.4 are generalized by the following:
Example 23.6.
$KO_{*}(\real_{0}({\bf{R}}^{n}))=KO_{*}(S^{n}{\mathbb{R}})=\Sigma^{n}KO_{*}({\mathbb{R}})$
and
$KO_{*}(\real_{0}(i{\bf{R}}^{n}))=KO_{*}(V^{n}{\mathbb{R}})=\Sigma^{-n}KO_{*}({\mathbb{R}}).$
Furthermore, let $\tau^{p,q}$ be the involution on ${\bf{R}}^{n}$ (where $n=p+q$) given by the identity on $p$ coordinates and by multiplication by 1-1 on the remaining $q$ coordinates. Then
$KO_{*}(\real_{0}({\bf{R}}^{n},\tau^{p,q}))=\Sigma^{p-q}KO_{*}({\mathbb{R}}).$
Example 23.7.
n= 0 11 22 33 44 55 66 77
$KO^{-n}(S^{1})\cong KO_{n}(\real(S^{1}))$ ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}^{2}$ ${\mathbb{Z}}_{2}$ $\mathbb{Z}$} $\mathbb{Z}$} 0 0 $\mathbb{Z}$}
$K^{-n}(S^{1})\cong K_{n}(C^{{\mathbb{C}}}(S^{1}))$ $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$}
Proof.
We regard $S^{1}$ as the usual subset of the complex plane and let
$\mathrm{ev}_{1}\colon\real(S^{1})\rightarrow{\mathbb{R}}$
[…]
$0\rightarrow\real_{0}({\bf{R}})\rightarrow\real(S^{1})\xrightarrow{\mathrm{ev}_{1}}{\mathbb{R}}\rightarrow 0\;.$
Furthermore, there is a splitting $s\colon{\mathbb{R}}\rightarrow\real(S^{1})$ which is given by constant functions. This implies that the long exact sequence
$\cdots\rightarrow KO_{*}(\real_{0}({\bf{R}}))\rightarrow KO_{*}(\real(S^{1}))\xrightarrow{(\mathrm{ev}_{1})_{*}}KO_{*}({\mathbb{R}})\rightarrow\cdots$
splits and we have an isomorphism
$KO_{*}(\real(S^{1}))\cong KO_{*}({\mathbb{R}})\oplus KO_{*}(\real_{0}({\bf{R}}))\cong KO_{*}({\mathbb{R}})\oplus\Sigma KO_{*}({\mathbb{R}})\;.$
The complex part $K_{*}(C^{{\mathbb{C}}}(S^{1}))$ is computed using the same split sequence. ∎
Example 23.8.
Recall that $iS^{1}$ denotes the circle with the involution $z\mapsto\overline{z}$ and
$\real(iS^{1})=\{f\in C^{{\mathbb{C}}}(S^{1})\mid\overline{f(z)}=f(\overline{z})\}$
is the corresponding real $C^{*}$-algebra.
n= 0 11 22 33 44 55 66 77
$KO_{n}(\real(iS^{1}))$ $\mathbb{Z}$} ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}^{2}$ ${\mathbb{Z}}_{2}$ $\mathbb{Z}$} $\mathbb{Z}$} 0 0
$K_{n}(C^{{\mathbb{C}}}(S^{1}))$ $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$}
Proof.
Since $1\in S^{1}$ is a fixed point of the involution $z\mapsto\overline{z}$, the condition that $\overline{f(z)}=f(\overline{z})$ implies that $f(1)\in{\mathbb{R}}$. So as in the previous calculation, we have a map $\mathrm{ev}_{1}\colon\real(iS^{1})\rightarrow{\mathbb{R}}$; this time the kernel is isomorphic to $V{\mathbb{R}}=\real_{0}(i{\mathbb{R}})$. Again, as in the previous calculation, the map splits yielding
$KO_{*}(\real(iS^{1}))\cong KO_{*}({\mathbb{R}})\oplus KO_{*}(V{\mathbb{R}})\cong KO_{*}({\mathbb{R}})\oplus\Sigma^{-1}KO_{*}({\mathbb{R}})\;.$
Example 23.9.
More generally, we have
$KO_{*}(\real(S^{n}))$ $=(1+\Sigma^{n})KO_{*}({\mathbb{R}})\;$
$\text{and~}KO_{*}(\real(iS^{n}))$ $=(1+\Sigma^{-n})KO_{*}({\mathbb{R}})\;.$
This follows from the Künneth Formula for real $K$-theory.
Example 23.10.
For completeness, we discuss the other possible involution on $S^{1}$. Recall that $\sigma$ is the involution on $S^{1}$ given by $z\mapsto-z$. For this example, we refer to Corollary 1.6 of [15].
n= 0 11 22 33 44 55 66 77
$KO_{n}(\real(S^{1},\sigma))$ $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$} $\mathbb{Z}$} ${\mathbb{Z}}_{2}$ 0 $\mathbb{Z}$}
$K_{n}(\real(S^{1},\sigma))$ $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$} $\mathbb{Z}$}
Example 23.11.
Let $S^{1}\times S^{1}$ be the torus. Then
$KO^{*}(S^{1}\times S^{1})=KO_{*}({\mathbb{R}})\oplus\Sigma KO_{*}({\mathbb{R}})^{2}\oplus\Sigma^{2}KO_{*}({\mathbb{R}})$
That is,
n= 0 11 22 33 44 55 66 77
$KO^{-n}(S^{1}\times S^{1})$ ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}^{3}$ ${\mathbb{Z}}_{2}^{3}$ ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ ${\mathbb{Z}}^{2}$ $\mathbb{Z}$} 0 $\mathbb{Z}$} ${\mathbb{Z}}^{2}\oplus{\mathbb{Z}}_{2}$
$K^{-n}(S^{1}\times S^{1})$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$ ${\mathbb{Z}}^{2}$
Proof.
Since $KO_{*}(\real(S^{1}))=KO_{*}({\mathbb{R}})\oplus\Sigma KO_{*}({\mathbb{R}})$, we can use the version of the Künneth formula expressed in Theorem 7.3. We have
$KO_{*}(\real(X))$ $=KO_{*}(\real(S^{1})\otimes_{{\mathbb{R}}}\real(S^{1}))$
$\cong KO_{*}(\real(S^{1}))\otimes_{KO_{*}({\mathbb{R}})}KO_{*}(\real(S^{1}))$
$=(1+\Sigma)KO_{*}({\mathbb{R}})\otimes_{KO_{*}({\mathbb{R}})}(1+\Sigma)KO_{*}({\mathbb{R}})$
$=(1+\Sigma)^{2}KO_{*}({\mathbb{R}})\;.$
Example 23.12.
More generally, consider the $n$-torus $S^{1}\times\overset{n}{\cdots}\times S^{1}$. Then
$KO^{-*}(S^{1}\times\overset{n}{\cdots}\times S^{1})\cong KO_{*}(\real(X))=\left(\prod_{i=1}^{n}(1+\Sigma)\right)KO_{*}({\mathbb{R}})\;.$
Even more generally, consider a generalized torus where each factor is a sphere of arbitrary dimension: $X=\prod_{i=1}^{n}S^{k_{i}}$. Then
$KO_{*}(\real(X))=\left(\prod_{i=1}^{n}(1+\Sigma^{k_{i}})\right)KO_{*}({\mathbb{R}})\;.$
24. $KO$-theory with Coefficients
Let $G$ be a countable abelian group and let $N_{G}$ be any complex $C^{*}$-algebra in the bootstrap class satisfying
$K_{*}(N_{G})=(G,0)\;.$
That such a $C^{*}$-algebra exists is a standard result (see for example the proof of Proposition 7.4 of [71] where such an $N_{G}$ is constructed and is also commutative). Furthermore, any two such $C^{*}$-algebras are $KK$-equivalent by Proposition 7.3 of [71] (this is a consequence of the Universal Coefficient Theorem for complex $K$-theory). Furthermore, by the Künneth Theorem in [73], the following definition of $K$-theory with coefficients is well-defined (independent of the choice of $N_{G}$).
Definition 24.1.
In the context of complex $C^{*}$-algebras, $K$-theory with $G$-coefficients is defined as
$K_{*}(A;G)=K_{*}(A\otimes N_{G})\;.$
In [75], the second author introduced mod $p$ coefficients into complex $K$-theory. In the present section we do something similar but considerably more complicated with real $K$-theory. It is much more complicated because of the group $G$ being used and especially by the fact that $KO$-theory itself is more complicated and has both torsion and torsion-free elements in its coefficient group.
Let $G$ be a countable abelian group and let $N^{\mathbb{R}}_{G}$ denote any real $C^{*}$-algebra in the real bootstrap category such that
  • $K_{*}((N^{\mathbb{R}}_{G})_{\mathbb{C}})=(G,0)$
  • $KO_{0}(N^{\mathbb{R}}_{G})=G$
  • $c_{0}\colon KO_{0}(N^{\mathbb{R}}_{G})\rightarrow K_{0}((N^{\mathbb{R}}_{G})_{\mathbb{C}})$ is an isomorphism.
Proposition 24.2.
For any countable abelian group $G$, there exists a real $C^{*}$-algebra $N_{G}^{\mathbb{R}}$ satisfying the given properties. Furthermore, the algebra $N_{G}^{\mathbb{R}}$ is unique in the sense that if $N_{G}^{\mathbb{R}}$ and $\widetilde{N}_{G}^{\mathbb{R}}$ both satisfy the given properties, then there is an isomorphism $KO_{*}(N_{G}^{\mathbb{R}})\cong KO_{*}(\widetilde{N}_{G}^{\mathbb{R}})$.
The stronger and more precise statement is that $N_{G}^{\mathbb{R}}$ and $\widetilde{N}_{G}^{\mathbb{R}}$ are $KK$-equivalent (via real $KK$-theory). That is, there is an element
$\alpha\in KK^{{\mathbb{R}}}(N_{G}^{\mathbb{R}},\widetilde{N}_{G}^{\mathbb{R}})$
that implements the isomorphism stated in the Proposition 24.2. We will not prove this proposition since it requires the use of united $K$-theory which we have not developed. It relies on results from Section 8.4 of [42] and the Main Theorem of [17].
Definition 24.3.
For a real $C^{*}$-algebra $A$, define
$KO_{*}(A;G)=K_{*}(A\otimes N_{G}^{\mathbb{R}})\;.$
The next two Theorems will also be stated without proof. The proofs also require the development of united $K$-theory and the full Künneth Formula and the Universal Coefficient Theorem for real $C^{*}$-algebras in [15] and [16].
Theorem 24.4.
$KO_{*}(A;G)$ is well-defined, independent of the choice of $N_{G}^{\mathbb{R}}$. The functor
$A\Longrightarrow KO_{*}(A;G)$
is a homology theory in sense of Definition 5.1.
Theorem 24.5.
Let $\alpha\colon G\rightarrow H$ be a homomorphism of countable abelian groups. Then there exists a natural transformation
$\theta_{\alpha}\colon KO_{*}(A;G)\rightarrow KO_{*}(A;H)$
such that $(\theta_{\alpha})_{0}=\alpha$.
Theorem 24.6.
For any real $C^{*}$-algebra $A$, there is an isomorphism
$KO_{i}(A;{\mathbb{Q}})=KO_{i}(A)\otimes{\mathbb{Q}}\;.$
Proof.
Let $\{\alpha\}_{n\in{\mathbb{N}}}$ be a sequence of homomorphisms $\alpha_{n}\colon{\mathbb{Z}}\rightarrow{\mathbb{Z}}$ such that
$\lim_{n\to\infty}({\mathbb{Z}},\alpha_{n})={\mathbb{Q}}\;.$
Each homomorphism $\alpha_{n}$ is multiplication by some integer $k_{n}$. Let
$\phi_{n}\colon N_{\mathbb{Z}}^{\mathbb{R}}\rightarrow N_{\mathbb{Z}}^{\mathbb{R}}$
be a homomorphism of real $C^{*}$-algebras such that $\phi_{n}$ induces the homomorphism $\alpha_{n}$ on real $K$-theory in degree 0. Then it is easy to verify that
$\lim_{n\to\infty}(N_{\mathbb{Z}}^{\mathbb{R}},\phi_{n})=N_{\mathbb{Q}}^{\mathbb{R}}\;.$
Then we also have
$\lim_{n\to\infty}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}},\mathrm{id}\otimes\phi_{n})=A\otimes N_{\mathbb{Q}}^{\mathbb{R}}\;.$
Note that each map
$(\mathrm{id}\otimes\phi_{n})_{*}\colon KO_{*}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}})\rightarrow KO_{*}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}})$
is multiplication by $k_{n}$ on $KO_{*}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}})=KO_{*}(A)$.
Thus
$KO_{*}(A;{\mathbb{Q}})$ $=KO_{*}(A\otimes N_{\mathbb{Q}}^{\mathbb{R}})$
$=\lim_{n\to\infty}(KO_{*}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}}),(\mathrm{id}\otimes\phi_{n})_{*})$
$=\lim_{n\to\infty}(KO_{*}(A),k_{n})$
$=KO_{*}(A)\otimes{\mathbb{Q}}$
In the remainder of this section, we investigate in more detail the properties of real $K$-theory with ${\mathbb{Z}}_{n}$-coefficients. Following the calculations of Section 5.1 of [15], we have the following structure for the $K$-theory of the real $C^{*}$-algebra $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$:
For $n$ odd:
$i$ 0 11 22 33 44 55 66 77
$KO_{i}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})$ ${\mathbb{Z}}_{n}$ 0 0 0 ${\mathbb{Z}}_{n}$ 0 0 0
$K_{i}((N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})_{\mathbb{C}})$ ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0
For $n\equiv 2\pmod{4}$:
$i$ 0 11 22 33 44 55 66 77
$KO_{i}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}))$ ${\mathbb{Z}}_{n}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{4}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{n}$ 0 0 0
$K_{i}((N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})_{\mathbb{C}})$ ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0
For $n\equiv 0\pmod{4}$:
$i$ 0 11 22 33 44 55 66 77
$KO_{i}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})$ ${\mathbb{Z}}_{n}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{2}^{2}$ ${\mathbb{Z}}_{2}$ ${\mathbb{Z}}_{n}$ 0 0 0
$K_{i}((N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})_{\mathbb{C}})$ ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0 ${\mathbb{Z}}_{n}$ 0
For the record, there are are three natural ways to obtain models for $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$.
  1. (1)
    Commutative Models.
    Let $D^{2}$ be the unit disk and let $S^{1}$ be the unit circle in the complex plane. For $n\geq 2$, let $\zeta_{n}=\exp(2\pi j/n)\in S^{1}$ be the principal root of unity and define an equivalence relation on $S^{1}=\partial(D^{2})\subset{\mathbb{C}}$ generated by $z\sim\zeta_{n}\cdot z$. Then define the reduced lens space by
    $L_{n}=(D^{2}-\{1\})/\sim\;.$
    There is a short exact sequence of complex $C^{*}$-algebras
    $0\rightarrow C_{0}(\rm{int}(D^{2}),{\mathbb{C}})\xrightarrow{\iota}C_{0}(L_{n},{\mathbb{C}})\xrightarrow{\rm{~ev~}}C_{0}(S^{1}-\{1\},{\mathbb{C}})\rightarrow 0$
    where $\rm{ev}$ is the evaluation map at the boundary. In the resulting 6-term exact sequence on $K$-theory, the differential map
    $\partial_{1}\colon K_{1}(C(S^{1}-\{1\},{\mathbb{C}}))\rightarrow K_{0}(C_{0}(\rm{int}(D^{2}),{\mathbb{C}}))$
    is multiplication by $n$ (from $\mathbb{Z}$} to $\mathbb{Z}$}) and it follows that
    $K_{*}(C_{0}(L_{n}),{\mathbb{C}})=(0,{\mathbb{Z}}_{n}).$
    This is the complex ${\mathbb{Z}}_{n}$-model used by the second author in [75].
    The same short exact sequence for real $C^{*}$-algebras
    $0\rightarrow C_{0}(\rm{int}(D^{2}),{\mathbb{R}})\xrightarrow{\iota}C_{0}(L_{n},{\mathbb{R}})\xrightarrow{\rm{~ev~}}C_{0}(S^{1}-\{1\},{\mathbb{R}})\rightarrow 0$ (4)
    yields a 24-term exact sequence on $KO$-theory. Note that
    $KO_{*}(C_{0}(S^{1}-\{1\},{\mathbb{R}}))=KO_{*}(S{\mathbb{R}})=\Sigma KO_{*}({\mathbb{R}})$
    and
    $KO_{*}(C_{0}(\rm{int}(D^{2})),{\mathbb{R}})=KO_{*}(S^{2}{\mathbb{R}})=\Sigma^{2}KO_{*}({\mathbb{R}})\ ;.$
    The differential map
    $\partial_{1}\colon KO_{-1}(C(S^{1}-\{1\},{\mathbb{R}}))\rightarrow KO_{-2}(C_{0}(\rm{int}(D^{2})),{\mathbb{R}})$
    is again multiplication by $n$ and from this the 24-term long exact sequence yields
    $KO_{-2}(C_{0}(L_{n},{\mathbb{R}}))={\mathbb{Z}}_{n}\quad\text{and}\quad KO_{-3}((C_{0}(L_{n},{\mathbb{R}}))=0\;.$
    This is a double shift of what is specified in the definition, so to correct for this we take $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=V^{2}C_{0}(L_{n},{\mathbb{R}})$.
  2. (2)
    Matrix Models.
    The real “dimension-drop algebra” $\mathbb{I}_{n}^{\mathbb{R}}$ for any $n\geq 2$ is defined by
    $\mathbb{I}_{n}^{\mathbb{R}}=\{f\in C_{0}((0,1],M_{n}({\mathbb{R}}))\mid f(1)\in{\mathbb{R}}\cdot I_{n}\}$
    as in Section 6 of [19]. The $K$-theory of these algebras is calculated in Theorem 6.1 of [19] and we find that $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=V\mathbb{I}_{n}^{\mathbb{R}}$. This is the real analog of the complex $C^{*}$-algebra used in [34] for complex $K$-theory with coefficients.
  3. (3)
    Unital, Simple, Purely Infinite Models.
    The real Cuntz-algebra $\mathcal{O}^{\mathbb{R}}_{n+1}$ is defined to be the unitary real $C^{*}$-algebra generated by partial isometries $s_{1},\dots,s_{n+1}$ that satisfy
    $s_{n}^{*}s_{n}=1\text{~for all $k$}\quad\text{and}\quad\sum_{k=1}^{n+1}s_{k}s_{k}^{*}=1\;.$
    Then by the real $K$-theory calculations for $\mathcal{O}^{\mathbb{R}}_{n+1}$ in Theorem 1.6.8 of [76] and Section 5.1 of [15] we have $KO_{-1}(\mathcal{O}_{n+1}^{\mathbb{R}})=0$ and $KO_{0}(\mathcal{O}_{n+1}^{\mathbb{R}})={\mathbb{Z}}_{n}$. Therefore $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=\mathcal{O}^{\mathbb{R}}_{n+1}$. Note that by Theorem 10.2 of [20], any unital simple purely infinite model for $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$ is isomorphic to $\mathcal{O}_{n+1}^{\mathbb{R}}$ (this is stronger than what is stated in Proposition 24.2).
Theorem 24.7.
For an integer $n\geq 2$, there exist natural transformations
$\rho_{n}$ $\colon KO_{*}(A)\rightarrow KO_{*}(A;{\mathbb{Z}}_{n})$ (of degree 0)
$\beta_{n}$ $\colon KO_{*}(A;{\mathbb{Z}}_{n})\rightarrow KO_{*}(A)$ (of degree 1-1)
such that the following Bockstein sequence is exact:
$\cdots\rightarrow KO_{k}(A)\xrightarrow{\rho_{n}}KO_{k}(A;{\mathbb{Z}}_{n})\xrightarrow{\beta_{n}}KO_{k-1}(A)\xrightarrow{\times n}KO_{k-1}(A)\rightarrow\cdots$
Proof.
The existence of $\rho_{n}$ follows from Theorem 24.5 based on the surjective homomorphism ${\mathbb{Z}}\rightarrow{\mathbb{Z}}_{n}$, but to obtain the exact sequence we take a different approach using the lens space model of $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. Start with the short exact sequence shown in displayed Equation (4) and tensor with a real $C^{*}$-algebra $A$. Since commutative $C^{*}$-algebras are exact, the resulting sequence
$0\rightarrow C_{0}(\rm{int}(D^{2}),A)\xrightarrow{\iota}C_{0}(L_{n},A)\xrightarrow{\rm{~ev~}}C_{0}(S^{1}-\{1\},A)\rightarrow 0$
is still exact. Using the isomorphisms
$C_{0}(\rm{int}(D^{2}),A)\cong S^{2}A\quad\text{and}\quad C_{0}(S^{1}-\{1\},A)\cong SA\;,$
the long exact sequence on $K$-theory can be formulated as
$\cdots\rightarrow KO_{k}(S^{2}A)\xrightarrow{\iota_{*}}KO_{k}(C_{0}(L_{n},A))\xrightarrow{\mathrm{ev}_{*}}KO_{k}(SA)\xrightarrow{\partial}KO_{k-1}(S^{2}A)\rightarrow\cdots\;.$
Then by using $C_{0}(L_{n},{\mathbb{R}})=S^{2}N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$ this becomes
$\cdots\rightarrow KO_{k+2}(A)\xrightarrow{\iota_{*}}KO_{k+2}(A;{\mathbb{Z}}_{n}))\xrightarrow{\mathrm{ev}_{*}}KO_{k+1}(A)\xrightarrow{\partial}KO_{k+1}(A)\rightarrow\cdots$
This gives the natural transformations $\rho_{n}=i_{*}$ and $\beta_{n}=\mathrm{ev}_{*}$. The differential $\partial\colon KO_{*}(A)\rightarrow KO_{*}(A)$ is known from [75] to be multiplication by $n$ in the case of $K_{0}({\mathbb{C}})={\mathbb{Z}}$ for complex $K$-theory. Then it must be multiplication by $n$ in the real case $KO_{0}({\mathbb{R}})$ since there is an isomorphism
$c\colon KO_{0}({\mathbb{R}})\rightarrow K_{0}({\mathbb{C}})\;.$
It follows that $\partial$ is multiplication by $n$ for any $A$ in the bootstrap class, using the Universal Coefficient Theorem for real $C^{*}$-algebras in [16]. ∎
Proposition 24.8.
For integers $m,n\geq 1$, there are natural transformations
$\kappa_{n,m}$ $\colon KO_{i}(A;{\mathbb{Z}}_{m})\rightarrow KO_{i}(A;{\mathbb{Z}}_{n})$
$\beta_{n,m}$ $\colon KO_{i}(A;{\mathbb{Z}}_{m})\rightarrow KO_{i-1}(A;{\mathbb{Z}}_{n})$
such that the following sequence is exact:
$\cdots\rightarrow KO_{k}(A;{\mathbb{Z}}_{m})\xrightarrow{\kappa_{mn,m}}KO_{k}(A;{\mathbb{Z}}_{mn})\xrightarrow{\kappa_{m,mn}}KO_{k}(A;{\mathbb{Z}}_{n})\xrightarrow{\beta_{m,n}}KO_{k-1}(A;{\mathbb{Z}}_{m})\rightarrow\cdots$
Proof.
The existence of $\kappa_{n,m}$ follows from Theorem 24.5 based on the homomorphism
${\mathbb{Z}}_{m}\rightarrow{\mathbb{Z}}_{n}\;$
which is multiplication by $n/\gcd(m,n)$. Also, $\beta_{n,m}$ is defined to be the composition $\beta_{n,m}=\rho_{m}\circ\beta_{n}$.
For the long exact sequence, we use the lens space model of $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. There is an obvious injection $L_{nm}\rightarrow L_{n}$ which induces a homomorphism
$\alpha\colon N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}\rightarrow N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}}\;.$
Now we apply the mapping cone construction (Definition 4.9) to get the short exact sequence
$0\rightarrow SN_{{\mathbb{Z}}_{m}n}^{\mathbb{R}}\rightarrow C\alpha\rightarrow N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}\rightarrow 0\;$
and the corresponding long exact sequence is
$\cdots\rightarrow KO_{i-1}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})\xrightarrow{\alpha_{*}}KO_{i-1}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{i}(C\alpha)\rightarrow KO_{i}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})\rightarrow\cdots\;.$
Note that the partial map arising from this long exact sequence coincides with the map $\alpha$ by Theorem 3.5 of [74], which applies in the case of real $K$-theory just as for complex $K$-theory. Using the fact that $\alpha_{*}$ is multiplication by $m$ on
$KO_{0}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})\rightarrow KO_{0}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\quad\text{(which is ${\mathbb{Z}}_{n}\rightarrow{\mathbb{Z}}_{mn}$)}$
we can calculate that $KO_{-2}(C\alpha)=0$ and $KO_{-1}(C\alpha)=K_{-1}((C\alpha)_{\mathbb{C}})={\mathbb{Z}}_{m}$. It follows that $C\alpha=SN_{{\mathbb{Z}}_{m}}^{\mathbb{R}}$ and that the map
$KO_{i}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{i+1}(C\alpha)$
is equivalent to the map
$KO_{i}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{i}(N_{{\mathbb{Z}}_{m}}^{\mathbb{R}})$
induced by multiplication by the quotient map ${\mathbb{Z}}_{mn}\rightarrow{\mathbb{Z}}_{m}$. ∎
We finish with a statement of the existence of a short exact sequence which can be used to compute $K$-theory with general coefficients. In the case of complex $K$-theory and a complex $C^{*}$-algebra $A$, there is a short exact sequence
$0\rightarrow K_{i}(A)\otimes G\rightarrow K_{i}(A;G)\rightarrow\rm{Tor}(K_{i+1}(A),G)\rightarrow 0\;$
for $i=0,1$, which is a special case of the Künneth Formula for complex $C^{*}$-algebras in [73]. In the real case, the Künneth Formula for real $C^{*}$-algebras must be expressed in terms of united $K$-theory, giving the result below. We state it here for completeness, but a discussion of united $K$-theory is beyond the scope of this paper. The middle term is the united $K$-theory of $A$ and it includes $KO_{*}(A;G)$, which is likely the group of interest, but also includes $K_{*}(A_{\mathbb{C}};G)$ and another group $KO_{*}(A\otimes T;G)$. It is necessary to move to this larger invariant to articulate the result. See [15] for the details.
Theorem 24.9.
For a countable abelian group $G$ and a real $C^{*}$-algebra, there is a short exact sequence
$0\rightarrow K^{\mathcal{CRT}}(A)\otimes_{\mathcal{CRT}}\mathcal{M}_{G}\xrightarrow{\mu}K^{\mathcal{CRT}}(A;G)\xrightarrow{\partial}\rm{Tor}_{\mathcal{CRT}}(K^{\mathcal{CRT}}(A),\mathcal{M}_{G})\rightarrow 0\;$
where $\partial$ has degree -1.
25. $K$-theory and Unitaries
In this section, we will present a description of $K$-theory entirely in terms of unitaries. The picture of $K$-theory that we describe here is well-known in the complex case, as well as in the real case for $KO_{0}$ and $KO_{1}$. In [18] a unitary picture of $K$-theory was expanded to incorporate all cases of the 10-fold way. This approach can be especially useful to understand the elements of $KO_{i}(A)$ in a concrete way for all values of $i$.
The real $K$-theory groups $KO_{i}(A)$ of a real $C^{*}$-algebra $A$ were defined in Definitions 6.10 and 6.11 in terms of homotopy groups, or in terms of suspensions. We have also seen that it can be defined in terms of Clifford algebras as in Section 18. The picture of $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ here is equivalent to those definitions but is a distinct picture of $K$-theory. It has the advantage of giving us a way to represent all elements of $KO_{i}(A)$ in terms of a specific (homotopy class of a) unitary in a matrix algebra over $A_{\mathbb{C}}$. This concrete representation allows physicists to develop formulas for topological markers.
In this alternative definition of real $K$-theory, we represent elements of $KO_{i}(A)$ using unitaries in a matrix algebra over $A_{\mathbb{C}}$ that satisfy a particular symmetry relation (depending on $i$). Given a real $C^{*}$-algebra $A$, let $\tau$ be the anti-multiplicative involution on $A_{\mathbb{C}}$ given by
$\tau(a+ib)\equiv(a+ib)^{\tau}=a^{*}+ib^{*}.$
Also, let $\tau$ be the extended anti-multiplicative involution defined by $\tau=\rm{Tr}\otimes\tau$ on $M_{n}(A_{\mathbb{C}})=M_{n}({\mathbb{R}})\otimes A_{\mathbb{C}}$. Note that we are using the condensed notation $a^{\tau}=\tau(a)$ for $a\in A$ or $a\in M_{n}(A_{\mathbb{C}})$ and recall that an element $a$ in $M_{n}(A_{\mathbb{C}})$ is in the underlying real $C^{*}$-algebra $M_{n}(A)$ if and only if it satisfies $a^{\tau}=a^{*}$.
Table 1 is reproduced from [18], and it summarizes the unitary picture of $K$-theory that will be described in this section. Each line gives the way of viewing $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ in terms of unitaries. For example, Propositions 25.1 and 25.2 below are summarized in the lines of Table 1 corresponding to $K_{0}(A_{\mathbb{C}})$ and $KO_{0}(A)$, respectively. The other lines in Table 1 similarly show how to represent $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ for $i\neq 0$ in this unitary picture, and they correspond to the propositions in this section which state the result explicitly.
Table 1. Unitary Picture of $K$-theory
K-group $\ell$ $\mathcal{S}$ II
complex $K_{0}(A_{\mathbb{C}})$ 2 $u=u^{*}$ $\left(\begin{smallmatrix}{1}&{0}\\ {0}&{-1}\end{smallmatrix}\right)$
$K_{1}(A_{\mathbb{C}})$ 1 11
real $KO_{-1}(A)$ 1 $u^{\tau}=u$ 11
$KO_{0}(A)$ 2 $u=u^{*}$, $u^{\tau}=u^{*}$ $\left(\begin{smallmatrix}{1}&{0}\\ {0}&{-1}\end{smallmatrix}\right)$
$KO_{1}(A)$ 1 $u^{\tau}=u^{*}$ 11
$KO_{2}(A)$ 2 $u=u^{*}$, $u^{\tau}=-u$ $\left(\begin{smallmatrix}{0}&{i\cdot 1}\\ {-i\cdot 1}&{0}\end{smallmatrix}\right)$
$KO_{3}(A)$ 2 $u^{\sharp\otimes\tau}=u$ $1_{2}$
$KO_{4}(A)$ 4 $u=u^{*}$, $u^{\sharp\otimes\tau}=u^{*}$ ${\mathrm{diag}}(1_{2},-1_{2})$
$KO_{5}(A)$ 2 $u^{\sharp\otimes\tau}=u^{*}$ $1_{2}$
$KO_{6}(A)$ 2 $u=u^{*}$, $u^{\sharp\otimes\tau}=-u$ $\left(\begin{smallmatrix}{0}&{i\cdot 1}\\ {-i\cdot 1}&{0}\end{smallmatrix}\right)$
Explanation of Table 1. For a real $C^{*}$-algebra $A_{\mathbb{C}}$, define
$M_{1,\infty}(A_{\mathbb{C}})$ $=\vee_{n=1}^{\infty}M_{n}(A_{\mathbb{C}})$
$M_{2,\infty}(A_{\mathbb{C}})$ $=\vee_{n=1}^{\infty}M_{2n}(A_{\mathbb{C}})$
$M_{4,\infty}(A_{\mathbb{C}})$ $=\vee_{n=1}^{\infty}M_{4n}(A_{\mathbb{C}})\;.$
Similarly, define $\mathcal{U}_{n}(A_{\mathbb{C}})$ to be the set of unitary elements in $M_{n}(A_{\mathbb{C}})$ and let
$\mathcal{U}_{1,\infty}(A_{\mathbb{C}})$ $=\vee_{n=1}^{\infty}\mathcal{U}_{n}(A_{\mathbb{C}})$
$\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$ $=\vee_{n=1}^{\infty}\mathcal{U}_{2n}(A_{\mathbb{C}})$
$\mathcal{U}_{4,\infty}(A_{\mathbb{C}})$ $=\vee_{n=1}^{\infty}\mathcal{U}_{4n}(A_{\mathbb{C}})\;.$
where $\vee$ denotes disjoint unions. We will discuss the identification of matrices of different sizes later. The way that we do this is not the same for all $i$.
For each row of the table, there is a given value of $\ell$ and the corresponding group $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ is isomorphic to a group [^85] In each of the ten cases we are taking equivalence classes of certain spaces of matrices. We must show in each case that this has indeed the structure of a group. consisting of equivalence classes of elements in $\mathcal{U}_{\ell,\infty}(A_{\mathbb{C}})$ that satisfy the relation $\mathcal{S}$. More precisely, if $a\in\mathcal{U}_{n}(A_{\mathbb{C}})$ (where $n$ is a multiple of $\ell$) and $a$ satisfies the relation $\mathcal{S}$, then $[a]$ represents an element of $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$. If two elements $a$ and $b$ in $\mathcal{U}_{n}(A_{\mathbb{C}})$ that satisfy $\mathcal{S}$ can be connected by a path $a_{t}$ in $\mathcal{U}_{n}(A_{\mathbb{C}})$ with $a_{0}=a$, $a_{1}=b$ and $a_{t}$ satisfying $S$ for each $t$, then we have $[a]=[b]$ in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ (depending on which row of the table we are working in). In each case the group operation for $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ is given by
$a]\oplus[b]=\left[\left(\begin{smallmatrix}a&0\\ 0&b\end{smallmatrix}\right)\right]\;.$
Finally the last entry of each row of Table 1 shows the unitary $I\in\mathcal{U}_{\ell,\infty}(A_{\mathbb{C}})$ that represents the trivial group element in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$. In particular this trivial element gives us the formulas
$I]=0\quad\text{and}\quad[a]=\left[\left(\begin{smallmatrix}a&0\\ 0&I\end{smallmatrix}\right)\right]\quad\text{in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$}.$
In what remains of this section, we explain this picture more fully and we state the propositions that form the basis of this picture, and we present some examples. For proofs of these propositions, including how one passes from disjoint unions to unions, see Boersema and Loring [20] Section 5 for the even case and Section 6 for the odd case.
We first discuss the cases for $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ when $i$ is even. As we know, $KO_{0}(A)$ is, generally speaking, built from projections in $A$ and $K_{0}(A_{\mathbb{C}})$ is built from projections in $A_{\mathbb{C}}$. Note that if $p$ is a projection in a real or complex unital $C^{*}$-algebra $A$, then u=2(p1)u=2(p-1) is a self-adjoint unitary element. Conversely, if $u$ is a self-adjoint unitary, then $p=\tfrac{1}{2}(u+1)$ is a projection. Using this relationship, we can express $K_{0}(A_{\mathbb{C}})$ and $KO_{0}(A)$ in terms of unitaries instead of projections, as in the following propositions.
Proposition 25.1.
Let $A$ be a real unital $C^{*}$-algebra. Then $K_{0}(A_{\mathbb{C}})$ is isomorphic to the group of homotopy classes of self-adjoint elements in $\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$.
Proposition 25.2.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{0}(A)$ is isomorphic to the group of homotopy classes of self-adjoint elements in $\mathcal{U}_{2,\infty}(A)$. Equivalently, $KO_{0}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$ of self-adjoint elements $u$ that satisfy $u=u^{\rm{Tr}\otimes\tau}$.
The identity of $KO_{0}(A)$ or $K_{0}(A_{\mathbb{C}})$ is the equivalence class $[I]$ where $I=\left(\begin{smallmatrix}1&0\\ 0&-1\end{smallmatrix}\right)$. Accordingly, an element $a\in M_{2n}(A_{\mathbb{C}})$ is identified with $\left(\begin{smallmatrix}a&0\\ 0&I\end{smallmatrix}\right)$ in $\mathcal{U}_{2n+2}(A_{\mathbb{C}})$. In other words,
$a]=\left[\left(\begin{smallmatrix}a&0\\ 0&I\end{smallmatrix}\right)\right]\quad\text{in $KO_{0}(A)$}.$
For $u],[v]\in K_{0}(A_{\mathbb{C}})$, the group operation is given by $u]+[v]=\left[\left(\begin{smallmatrix}u&0\\ 0&v\end{smallmatrix}\right)\right]$, and the inverse rule is given by $u]^{-1}=[-u]$.
Remarkably, by making a small sign change to the relation $\mathcal{S}$ characterizing $KO_{0}(A)$, we obtain the unitary picture of $KO_{2}(A)$, as in the next Proposition and as shown in the line of Table 1 for $KO_{2}(A)$. We will not give proofs for this proposition as indicated before.
Proposition 25.3.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{2}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$ of self-adjoint elements $u$ that satisfy $u=-u^{\tau}$.
In the group of unitaries for $KO_{2}(A)$ described above, the identity element is given by $[I]$ where $I=\left(\begin{smallmatrix}0&i\cdot 1\\ -i\cdot 1&0\end{smallmatrix}\right)$.
For $n=4$ and $n=6$ we let $\sharp$ be the involution
$\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\mapsto\left(\begin{smallmatrix}d&-b\\ -c&a\end{smallmatrix}\right)$
on $M_{2}(A)$, and let $\sharp\otimes\tau$ be the corresponding involution on $M_{2n}(A)=M_{2}({\mathbb{R}})\otimes M_{n}(A)$. Then we have the following propositions. For these propositions, the unitary $I$ such that $[I]$ represents the identity element of $KO_{i}(A)$ is listed in Table 1.
Proposition 25.4.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{4}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{4,\infty}(A_{\mathbb{C}})$ of self-adjoint unitaries $u$ that satisfy $u=u^{\sharp\otimes\tau}$.
Proposition 25.5.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{6}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$ of self-adjoint unitaries $u$ that satisfy $u=-u^{\sharp\otimes\tau}$.
This completes the discussion of the even cases of Table 1. In the odd cases, we first discuss the well-known exponential map in the complex case. The following is Theorem 10.1.3 [69].
Proposition 25.6.
Let $A$ be a real unital $C^{*}$-algebra. Then $K_{1}(A_{\mathbb{C}})$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{\infty}(A_{\mathbb{C}})$.
Sketch of Proof.
We have $K_{1}(A_{\mathbb{C}})=K_{0}(SA_{\mathbb{C}})$ (as presented in Theorem 8.5) and for the moment we let $K_{1}^{u}(A_{\mathbb{C}})$ denote the unitary group referred to in this proposition. It consists of homotopy classes of elements in $\mathcal{U}_{\infty}(A)_{\mathbb{C}})$, subject to the equivalence relation in which a unitary $u\in M_{n}(A_{\mathbb{C}})$ is identified with $\left(\begin{smallmatrix}u&0\\ 0&1\end{smallmatrix}\right)$ in $M_{n+1}(A_{\mathbb{C}})$. The identity is the class of the unit $[1]$ and the group inverse of an element $[u]$ in $K_{1}^{u}(A_{\mathbb{C}})$ is given by $u^{*}]$, since the unitary $\left(\begin{smallmatrix}u&0\\ 0&u^{*}\end{smallmatrix}\right)$ can be connected by a path of unitaries to $\left(\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right)$.
We will sketch the map $\Theta$ that gives an isomorphism between $K_{1}^{u}(A_{\mathbb{C}})$ and $K_{0}(SA_{\mathbb{C}})$, following that in [69].
Let $u$ be a unitary in $M_{n}(A_{\mathbb{C}})$. Let $x_{t}$ be a path of unitaries in $\mathcal{U}_{2n}({\mathbb{R}})$ from $\left(\begin{smallmatrix}0&1_{n}\\ 1_{n}&0\end{smallmatrix}\right)$ to $1_{2n}$, and let $v_{t}$ be the path
$v_{t}=\left(\begin{smallmatrix}u&0\\ 0&1_{n}\end{smallmatrix}\right)x_{t}\left(\begin{smallmatrix}1_{n}&0\\ 0&u^{*}\end{smallmatrix}\right)x_{t}^{*}$
from $1_{2n}$ to $\left(\begin{smallmatrix}u&0\\ 0&u^{*}\end{smallmatrix}\right)$.
Then $p_{t}=v_{t}\left(\begin{smallmatrix}1_{n}&0\\ 0&1_{n}\end{smallmatrix}\right)v_{t}^{*}$ is a projection in $C([0,1],M_{2n}(A_{\mathbb{C}}))$ with $p_{0}=\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)$ and $p_{1}=\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)$. The element
$p_{t}]-\left[\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)\right]\in K_{0}(C([0,1],M_{2n}(A_{\mathbb{C}}))$
is in the kernel of $(\mathrm{ev}_{0})_{*}\colon K_{0}(C([0,1],A_{\mathbb{C}}))\rightarrow K_{0}(A_{\mathbb{C}})$, which therefore defines an element $\chi_{u}\in K_{0}(SA_{\mathbb{C}})$. Then we define $\Theta([u])=\chi_{u}$.
We refer to the proof of Theorem 10.1.3 for the details that $\Theta$ is well-defined and gives an isomorphism, proving the Proposition. ∎
Proposition 25.7.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{1}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{\infty}(A_{\mathbb{C}})$ of unitaries $u$ that satisfy $u^{*}=u^{\tau}$.
Note here that the condition $u^{*}=u^{\tau}$ is equivalent to just requiring that $u$ is a unitary in $A\subset A_{\mathbb{C}}$.
Sketch of Proof.
This proof is the same as in the complex case, except we note that $\left(\begin{smallmatrix}0&1_{n}\\ 1_{n}&0\end{smallmatrix}\right)$ is not in the same homotopy class of $1_{2n}$ in $M_{2n}({\mathbb{R}})$ when $n$ is odd. However, by moving to a larger matrix algebra they become equivalent.
If $n$ is odd, then
$\left(\begin{smallmatrix}0&1_{n}&0\\ 1_{n}&0&0\\ 0&0&-1\end{smallmatrix}\right)$
is in the same homotopy class of unitaries as $1_{2n+1}$. Using this, we can produce a path of from $1_{2n+1}$ to $\mathrm{diag}(u,u^{*},1)$ in $M_{2n+1}({\mathbb{R}})$. The rest of the proof is the same as for Proposition 25.6. ∎
Proposition 25.8.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{-1}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{\infty}(A_{\mathbb{C}})$ of elements $u$ that satisfy $u=u^{\tau}$.
Proposition 25.9.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{3}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$ of elements $u$ that satisfy $u=u^{\sharp\otimes\tau}$.
Proposition 25.10.
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{5}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{2,\infty}(A_{\mathbb{C}})$ of elements $u$ that satisfy $u^{*}=u^{\sharp\otimes\tau}$.
To understand better how this picture of real $K$-theory works, we will discuss some examples of real $C^{*}$-algebras and see how some of the specific non-trivial $KO$-classes can be represented by unitaries satisfying their appropriate relations. These examples and more can be found with more detailed exposition in [18] and [19].
Example 25.11.
$KO_{i}({\mathbb{R}})$.
The table below shows the unitaries representing generators of all the non-trivial $KO_{i}({\mathbb{R}})$ groups.
[…]
As indicated by the table, the self-adjoint untiary $x_{0}=1_{2}$ represents a generator of $KO_{0}({\mathbb{R}})\cong{\mathbb{Z}}$ (and recall that the self-adjoint unitary $I=\mathrm{diag}(1,-1)$ represents the trivial class of $KO_{0}({\mathbb{R}})$). More generally, if $x$ is a self-adjoint unitary in $M_{2n}({\mathbb{R}})$, then $\tfrac{1}{2}\rm{tr}(x)$ is an integer, which determines the $KO_{0}({\mathbb{R}})$ class represnted by $x$. That is, $x]\mapsto\rm{tr}(x)$ is well-defined and gives an isomorphism from the unitary picture of $KO_{0}({\mathbb{R}})$ to $\mathbb{Z}$}. The same is true for $K_{*}({\mathbb{R}}_{\mathbb{C}})={\mathbb{Z}}$.
The unitary $x_{1}=-1$ represents the non-trivial element of $KO_{1}({\mathbb{R}})\cong{\mathbb{Z}}_{2}$. More generally, if $x$ is a unitary in $M_{2n}({\mathbb{R}})$, then $\det(x)$ is either $1$ or 1-1, and this map characterizes the class represented by $x$ in $KO_{1}({\mathbb{R}})$.
In the same way, the $KO_{4}({\mathbb{R}})$ class of a self-adjoint, skew-symmetric matrix is determined by its pfaffian.
Example 25.12.
$KO_{i}(S{\mathbb{R}})$
[…]
Since $S{\mathbb{R}}=C_{o}^{{\mathbb{R}}}({\mathbb{R}})$ is not unital, we must consider the unitization
$(S{\mathbb{R}})^{+}\cong\real(S^{1})$
and represent $KO_{*}(S{\mathbb{R}})$ as the kernel of the evaluation map
$\pi_{*}\colon KO_{*}(\real(S^{1}))\rightarrow KO_{*}({\mathbb{R}})$
where $\pi$ is evaluation at $z=1$. The unitaries in the fourth column of the table shows the generators of the various non-trivial $KO_{i}$ groups in this picture.
(Alternatively, we can represent the unitization by $(S{\mathbb{R}})^{+}\cong\real([0,1])$, and then we identify $KO_{*}(S{\mathbb{R}})$ with the kernel of
$\pi_{*}\colon KO_{*}(\real([0,1]))\rightarrow KO_{*}({\mathbb{R}})$
where $\pi$ is evaluation at $t=0$. The unitaries in the third column of the table show the generators in this picture.)
Furthermore, to make sense of this table, we need to think about the complex $C^{*}$-algebra $C^{{\mathbb{C}}}(S^{1})$ and the corresponding involution $\tau$ on $C^{{\mathbb{C}}}(S^{1})$. In this case we have $\tau=\mathrm{id}$.
For $KO_{0}(S{\mathbb{R}})$, we observe that $x_{0}=\left(\begin{smallmatrix}{x}&{y}\\ {y}&{-x}\end{smallmatrix}\right)$ is a self-adjoint unitary matrix in $M_{2}(\real(S^{1}))$. Since $\pi(x_{0})=\mathrm{diag}(1,-1)$ represents the trivial element of $KO_{0}({\mathbb{R}})$, we find that $x_{0}$ is a well-defined element of $KO_{0}(S{\mathbb{R}})$. According to the table, $x_{0}]$ represents the non-trivial element of $KO_{0}({\mathbb{R}})\cong{\mathbb{Z}}_{2}$ (whereas the trivial element of $KO_{0}(S{\mathbb{R}})$ would be represented by the constant matrix $\mathrm{diag}(1,-1)$). We also point out that one can confirm the fact that $KO_{0}(S{\mathbb{R}})\cong{\mathbb{Z}}_{2}$ by verifying that $\mathrm{diag}(x_{0},x_{0})$ can be connected to the identity by a homotopy through self-adjoint unitaries.
Now $KO_{-1}(S{\mathbb{R}})={\mathbb{Z}}$ is represented by unitaries $u$ that satisfy the relation $u=u^{\tau}$, which in this case collapses to $u=u^{\rm{Tr}}$. This relation is clearly satisfied by the unitary $x_{-1}=x+iy\in M_{1}(C^{{\mathbb{C}}}(S^{1}))$ and we also have that $\pi[x_{-1}]=[1]$ which is trivial in $KO_{-1}({\mathbb{R}})=0$. Hence, we have $x_{-1}]\in KO_{-1}(S{\mathbb{R}})$, and in fact $x_{-1}$ represents one of the generators of $KO_{-1}(S{\mathbb{R}})$.
Finally, the reader can check that the unitary $x_{3}$ satisfies $x_{3}^{\sharp\otimes\tau}=x_{3}$. This element turns out to be a non-trivial element of $KO_{3}(S{\mathbb{R}})\cong{\mathbb{Z}}$ and represents a free generator of that group.
Example 25.13.
$KO_{i}(V{\mathbb{R}})$.
[…]
Recall that
$V{\mathbb{R}}\cong\{f\in C_{0}({\mathbb{R}},{\mathbb{C}})\mid\text{$f(-x)=\overline{f(x)}$ for all $x\in{\mathbb{R}}$}\}$
and that
$KO_{*}(V{\mathbb{R}})=\Sigma^{-1}KO_{*}({\mathbb{R}})\;.$
The algebra $V{\mathbb{R}}$ and the previous one $S{\mathbb{R}}$ both have the same complexification (and the same unitzation of the complexification $C^{{\mathbb{C}}}(S^{1})$). The difference is the associated involution on the complex $C^{*}$-algebra $C^{{\mathbb{C}}}(S^{1})$ is now $f^{\tau}(z)=f(\overline{z})$ for $f\in C^{{\mathbb{C}}}(S^{1})$ and $z\in S^{1}\subset{\mathbb{C}}$. The element $x_{1}=x+iy$ satisfies $x_{1}^{\tau}=x_{1}^{*}$, so it represents an element in $KO_{1}(V{\mathbb{R}})$ (whereas in the previous example the same unitary represented an element in $KO_{-1}(S{\mathbb{R}})$).
We leave it to the reader to check that given elements $x_{2}$, $x_{3}$, and $x_{5}$ are also unitaries and that they satisfy the appropriate relations to represent the appropriate group $KO_{i}(V{\mathbb{R}})$. See [19] for further details.
We find it elucidating to note that $KO_{0}(V{\mathbb{R}})$ is represented by unitaries in $M_{2n}(C^{{\mathbb{C}}}(S^{1}))$ satisfying $u=u^{*}=u^{\tau}$. Furthermore, the condition that $\pi(1)$ represents the trivial element of $KO_{0}({\mathbb{R}})$ implies that the trace of such a unitary must be constantly 0. Using the functional calculus, then one can easily show that $u$ is equivalent to a constant unitary with trace 0. Thus we see in a very concrete way that $KO_{0}(V{\mathbb{R}})=0$.